### v2

# TRiSM for Agentic AI: A Review of Trust, Risk, and Security Management in LLM-based Agentic Multi-Agent Systems

Abstract
Agentic AI systems, built upon large language models (LLMs) and deployed in multi-agent configurations, are redefining intelligence, autonomy, collaboration, and decision-making across enterprise and societal domains. This review presents a structured analysis of Trust, Risk, and Security Management (TRiSM) in the context of LLM-based Agentic Multi-Agent Systems (AMAS). We begin by examining the conceptual foundations of Agentic AI and highlight its architectural distinctions from traditional AI agents. We then adapt and extend the AI TRiSM framework for Agentic AI, structured around four key pillars: Governance, Explainability, ModelOps, and Privacy/Security , each contextualized to the challenges of multi-agent LLM systems. A novel risk taxonomy is proposed to capture the unique threats and vulnerabilities of Agentic AI, ranging from coordination failures to prompt-based adversarial manipulation. To support practical assessment in Agentic AI works, we introduce two novel metrics: the Component Synergy Score (CSS), which quantifies the quality of inter-agent collaboration, and the Tool Utilization Efficacy (TUE), which evaluates the efficiency of tool use within agent workflows. We further discuss strategies for improving explainability in Agentic AI , as well as approaches to enhancing security and privacy through encryption, adversarial robustness, and regulatory compliance. The review concludes with a research roadmap for the responsible development and deployment of Agentic AI, outlining critical directions to align emerging systems with TRiSM principles for safe, transparent, and accountable operation.
keywords:
Agentic AI, LLM-based Multi-Agent Systems, TRiSM, AI Governance, Explainability, ModelOps, Application Security, Model Privacy, AI Agents, Trustworthy AI, Risk Management, AI Safety, Privacy-Preserving AI, Adversarial Robustness, Human-in-the-Loop
1 Introduction
1.1 Background and Motivation
The global market for Artificial Intelligence (AI) agents is projected to grow from \$5.4 billion in 2024 to \$7.6 billion in 2025  [1]. By mid-2025, over 70% of enterprise AI deployments are expected to involve multi-agent or action-based systems, reflecting a dramatic shift from traditional single-agent or rule-based conversational models [2]. An AI agent, typically defined as a computational entity that perceives its environment and takes actions to achieve goals  [3], has undergone a fundamental transformation. While early agents were task-specific and deterministic, modern systems have evolved significantly ,increasing operational autonomy by over 70% on benchmark tasks [4]. These Agentic AI systems now integrate large language models (LLMs), external tools, and persistent memory to support complex planning and coordinated decision-making [5, 6].
This rapid evolution presents new opportunities but also surfaces significant challenges in coordination, control, transparency, and risk assessment across heterogeneous agents. In particular, the shift toward Agentic AI architectures amplifies and evolves longstanding concerns around safety, fairness, interpretability, and systemic harms. While such issues have been central to machine learning (ML) systems broadly, they become particularly critical in Agentic AI, where autonomous agents operate in high-stakes domains like healthcare, law, and public services, areas where their decisions can carry profound and lasting societal consequences.
From isolated agents to Agentic AI The autonomous software agents date back to the 1990s [7], however, the advent of LLM-driven multi-agent ecosystems [8] since 2023 has made “ Agentic AI ” a distinctly new and rapidly expanding research focus [5]. Traditional agents primarily automated narrow tasks, such as information retrieval [9, 10], data summarization [11], and dialogue response [12]. Typically driven by single step logic [13] or scripted rules, these agents lacked the deep reasoning, adaptability, and persistence required for complex, multi-step problem solving [14, 15].
In contrast, Agentic AI consists of collaborative agents with specialized roles (e.g., planner, coder, analyst), enabled by LLMs and tool use [16]. These systems dynamically decompose tasks, share context, and pursue high-level goals across long timelines [17, 18, 19]. This shift reflects not only a change in technology but also a fundamental increase in complexity and autonomy, marking the emergence of machine collectives capable of exhibiting emergent and decentralized behavior.
1.2 Trust, Risk, and Security Challenges in Agentic AI
The rise of Agentic AI introduces complex challenges that extend beyond those of traditional ML systems. Unlike deterministic agents [20], LLM-based multi-agent systems exhibit emergent and often opaque decision-making, increasing the risk of cascading errors, biases, and unintended behaviors. These characteristics necessitate a robust approach to have trust and ensuring trustworthiness, in high-stakes domains such as healthcare, finance, and law.
Trust and Trustworthiness
In the context of Agentic AI, trust refers to a user willingness to rely on an AI system, while trustworthiness denotes whether the system consistently behaves in a safe, fair, and predictable manner, thereby deserving that trust. This distinction is critical in high-stakes settings, where undue reliance on an untrustworthy system can lead to severe consequences, such as misdiagnoses in healthcare or biased rulings in legal applications. The stochastic nature of LLM reasoning [21] further complicates trustworthiness, as randomness in output generation, while promoting creativity and ensemble diversity, introduces inconsistency and non-repeatability. These factors hinder traceability, verification, and compliance auditing, all of which are essential for trustworthy systems.
Security Risks in Multi-Agent Systems
Security risks escalate as autonomous agents gain access to tools, external APIs, and persistent memory, amplifying the potential for privacy breaches, adversarial misuse, or regulatory violations in the absence of robust human oversight [22]. For example, a multi-agent supply chain optimizer may autonomously coordinate between procurement and logistics agents but inadvertently leak sensitive data or violate compliance protocols (e.g., GDPR or HIPAA) if safeguards are inadequate. The distributed and cooperative nature of multi-agent systems introduces unique vulnerabilities that traditional security measures often fail to address.
Specific Threats: Prompt Injection, Spoofing, and Impersonation
LLM-based agents are susceptible to novel security threats exploiting their language-based interfaces and cooperative behaviors. One prominent threat is prompt injection [23], where attackers craft inputs with hidden or malicious instructions to manipulate agent behavior. Recent studies [24] describe “prompt infection,” a phenomenon where malicious prompts propagate from one agent to another, akin to a computer virus spreading across networks [25]. Such corrupted outputs can lead to cascading effects, including data leaks, fraudulent transactions, misinformation, or coordinated misbehavior within an agent society.
Another critical vulnerability involves spoofing and impersonation [26]. Spoofing entails faking an entity’s identity or credentials (e.g., using a false name or address) to deceive a system, while impersonation involves actively pretending to be a specific agent or user to gain trust, access, or privileges. In multi-agent systems, where agents rely on credentials or tokens for authentication during coordination, an adversary could steal an API key or mimic a trusted peer to issue unauthorized commands or access sensitive information.
Privacy Risks in Multi-Agent LLM Systems
LLM-based agents often need to handle user data, proprietary business information, or other sensitive inputs to fulfill their tasks [27]. In a multi-agent context, this challenge is amplified by the fact that agents may share information with each other (e.g. via a shared memory store or message passing) to collaborate.
Limitations of Traditional Frameworks
Traditional evaluation and safety frameworks, designed for deterministic agents or conventional AI systems, are increasingly inadequate for addressing the autonomy and complexity of Agentic AI. These frameworks often fail to account for emergent behaviors, inter-agent dependencies, or the dynamic nature of LLM-driven systems, leaving gaps in security, trust, and risk management.
Towards a TRiSM Framework for Agentic AI
To address these gaps, recent frameworks like AI TRiSM (Trust, Risk, and Security Management) [28, 29] propose lifecycle-level controls, including explainability, secure model orchestration, and privacy management. These components are vital for deploying Agentic AI in high-stakes domains such as finance, healthcare, and defense, where trust, security, and accountability are non-negotiable.
Figure 1: Taxonomy of TriSM for Agentic AI systems presented in this review. It includes five major sections: Fundamentals of Agentic AI Systems ( Agentic AI vs. Traditional AI Agents, LLM-Based Architecture), The TRISM Framework (Governance, Explainability, ModelOps, Application Security, Privacy), Threats and Risks , Trust and Explainability, and Evaluation Metrics (including Human-Centered and Composite Metrics).
1.3 Problem Statement
Agentic AI systems increasingly rely on the coordination of multiple LLM-powered agents to perform complex, long-horizon tasks. However, no unified framework exists for managing their Trust, Risk, and Security Management (AI TRiSM) across dynamic, multi-agent workflows. Recent failures, such as, an autonomous research agent citing unverifiable medical sources or collaborative swarms generating inconsistent forecasts due to memory drift, highlight that traditional safeguards are inadequate [30]. These traditional systems are vulnerable to orchestration breakdowns, agent collusion, data leakage, and opaque decision-making pathways. To ensure safety and reliability in Agentic AI systems, a structured TRiSM perspective is urgently needed. This study summarizes emerging findings and offers actionable guidelines to build secure, transparent, and trustworthy Agentic AI ecosystems.
1.4 Scope and Objectives
Objectives This study is a review article that maps the TRiSM (Trust, Risk, and Security Management) [28] landscape as it applies to LLM-powered, multi-agent “Agentic AI” systems. We structure the discussion around four foundational pillars of TRiSM [31]: explainability, security, lifecycle governance, and privacy. We argue that conventional safeguards developed for bounded, rule-based agents are inadequate for managing the complexities of open-ended, self-directed agent collectives. The objective of this review is to offer researchers and practitioners a coherent and actionable reference for the design, deployment, and governance of trustworthy Agentic AI ecosystems.
For the sake of this study, we define Agentic AI as Multi-Agent Systems (MAS) [32] powered by LLMs that exhibit autonomous planning, tool use, and memory retention, with or without human oversight. In contrast to classical MAS, these systems exhibit emergent reasoning capabilities and dynamically adapt their workflows based on interactions with tools, users, and other agents.
Scope This paper is intended as a conceptual survey and synthesis. It does not propose a new Agentic AI architecture or introduce synthetic benchmark evaluations. Rather, it integrates emerging perspectives on trust, risk, and security to support future research and responsible development in this evolving area.
Necessity of this Survey Agentic AI research has largely centered on agent modeling, planning, and task coordination, while comparatively overlooking critical dimensions such as adversarial robustness, lifecycle governance, and system-level explainability. As these systems are increasingly deployed in high-stakes domains - such as healthcare, scientific research, and finance—the absence of an integrated TRiSM framework exposes stakeholders to opaque decision-making and unmanaged risks. This survey addresses that gap by synthesizing TRiSM principles in the context of LLM-driven multi-agent systems and providing actionable guidance for researchers, engineers, and policymakers.
To the best of our knowledge, few existing surveys systematically explore LLM-based multi-agent systems through the lens of trust, risk, and security. Table 1 underscores this gap by comparing prior reviews with respect to their coverage of TRiSM dimensions.
1.5 Contributions
Our key contributions in this review are:
  1. 1.
    We propose the first TRiSM framework that is explicitly tailored to systems in which several LLM-powered agents collaborate and cooperate. The framework is organized around four pillars: Explainability, Lifecycle Governance, Application Security, and Model Privacy. It targets the specific governance challenges created by autonomous decision making, inter-agent coordination, and persistent memory (see Fig. 1 and Sec. 5).
  2. 2.
    We offer a systematic taxonomy of risks that emerge when multiple agents interact (Sec. 6). This taxonomy discusses prompt injection, memory poisoning, collusive failure, emergent misbehavior, and related threats, and it shows why defenses designed for single-model LLM applications are not sufficient.
  3. 3.
    We introduce a five-dimension evaluation framework that covers Trustworthiness, Explainability, User-centered Performance, Coordination, and Composite Assessment (Sec. 7.3). To support this template, we define two new metrics: the Component Synergy Score (CSS), which measures the quality of inter-agent collaboration, and the Tool Utilization Efficacy (TUE), which measures how correctly and efficiently agents invoke external tools.
  4. 4.
    We review state-of-the-art techniques for each TRiSM pillar (Secs. 78) and examine methods such as Local Interpretable Model-Agnostic Explanations (LIME), SHapley Additive exPlanations (SHAP), decision provenance, prompt hygiene, sandboxing, and ModelOps (model operations) pipelines. These techniques must be critically re-evaluated and adapted to suit environments involving multiple interacting LLM agents.
  5. 5.
    Building on the identified research gaps, we outline a research agenda that prioritizes scalable, verifiable, and regulation-compliant Agentic AI (Sec. 9).
1.6 Paper Organization
The remainder of this paper is structured as follows. Section 2 discuss the related work. Section 3 presents the literature search methodology. Section 4 reviews the core characteristics and architectures of Agentic AI systems, providing a foundation for the TRiSM analysis. Section 5 introduces the TRiSM framework and its four pillars: Explainability, ModelOps, Application Security, and Model Privacy. Section 6 outlines key risks, vulnerabilities, and attack surfaces in multi-agent AI ecosystems. Section 7 discusses trust-building and explainability techniques, while Section 8 covers security and privacy-enhancing methods. Section 9 synthesizes insights and proposes directions for research and policy. Section 10 concludes the paper with final reflections on the implications of TRiSM for the responsible development of Agentic AI. A list of key terms used throughout this paper is provided in Table A.1.
2 Related Work
Table 1: Comparison of Related Surveys on LLM-Based Multi-Agent Systems and TRiSM Aspects. Existing work offers limited coverage of trust, risk, and security in Agentic AI.
[…]
Guo et al. (2024) [33] ✓ (simulated env’ts) ~ (research challenges)
Chen et al. (2025) [34] ~ (task/simulation focus) ~ (future research areas)
Yan et al. (2025) [35] ✓ (mentions diverse) ~ (future work directions)
Tran et al. (2025) [36] ✓ (networks, QA, etc.) ~ (open challenges)
Lin et al. (2025) [37] ~ (creative tasks) ~ (roadmap for research)
Fang et al. (2025) [38] ~ ~ ✓ (health, finance cited) ~ (technical “next steps”)
This Survey (2025) ✓ (high-stakes domains) ✓ (tailored guidance)
Legend: ✓= explicitly addressed; ~= partially/indirectly addressed; ✗= not addressed.
[…]
We present the related work in this section.
2.1 LLM-Based Multi-Agent Surveys (Technical Focus)
Prior surveys on LLM-driven multi-agent systems predominantly focus on technical dimensions such as system architectures, agent capabilities, and domain-specific applications of Agentic AI, yet they rarely address critical TRiSM (Trust, Risk, and Security Management) considerations. For example, reviews like [33] and [34] delve into simulated environments, inter-agent communication protocols, and performance benchmarks. These works excel at detailing how agents coordinate or scale but fall short in examining adversarial threats, such as prompt injection attacks or agent collusion, lifecycle governance (e.g., continuous monitoring or regulatory compliance), or system-wide explainability, especially for emergent behaviors in complex systems.
More specialized surveys, such as [35] on natural language communication and [36] on collaboration mechanisms, explore interaction protocols and practical use cases in domains like question-answering systems, 5G networks, and Industry 5.0. While [35] briefly notes security risks tied to communication channels, and [36] mentions deployment contexts, both remain narrowly focused on coordination strategies. They lack deeper analysis of robustness against adversarial manipulation (e.g., Byzantine failures), governance frameworks for multi-agent deployments, or explainability challenges at scale.
2.2 Trustworthy/Responsible AI Surveys (Broad Trust Focus)
Surveys addressing TRiSM in AI, such as [38], tackle ethical and risk-related issues including alignment with human values, fairness in decision-making, and defenses against privacy attacks (e.g., data leakage). However, these works typically target general ML systems [39] and overlook the unique dynamics of LLM-based multi-agent systems. For instance, while they may advocate for privacy-preserving techniques like differential privacy, they do not address privacy risks in inter-agent data sharing or the distributed nature of LLM-driven systems. Moreover, critical TRiSM aspects [40] like governance (e.g., auditing autonomous agents) and explainability (e.g., tracing decisions across multiple agents) are often absent. As a result, their recommendations, while valuable for single-agent or centralized AI, do not fully translate to the complexities of multi-agent coordination in high-stakes domains.
2.3 Current Study (LLM Multi-Agent)
This review distinguishes itself by weaving TRiSM principles: adversarial robustness, end-to-end lifecycle governance, and system-wide explainability, into the study of LLM-based multi-agent systems, with a specific lens on high-stakes domains like healthcare, science, and finance. Unlike prior studies (see Table 1), which offer limited insight into governance and security, this work positions these elements as foundational design priorities. For example, it addresses targeted threats like backdoor attacks in distributed LLMs and governance challenges such as ensuring compliance with regulations (e.g., HIPAA in healthcare or GDPR in finance) across autonomous agents - areas largely ignored by technical surveys. This work provides actionable guidance for researchers, engineers, and policymakers, and fills critical gaps in both the technical multi-agent literature and broader responsible AI landscape.
3 Literature Review Methodology
To provide a comprehensive review of the literature on trust, risk, and security in Agentic AI systems, we adopted a structured methodology informed by best practices in systematic reviews [41, 42]. While this is not a formal systematic review, we followed a transparent and reproducible process to ensure coverage and rigor. This section outlines our research objectives, data sources, inclusion criteria, and classification strategy.
(a) Agentic AI papers per year (2022–2025).
(b) TRiSM pillars papers by topic.
Figure 2: (a) Annual trend in Agentic AI research, showing a sharp rise post-2023. (b) Distribution of TRiSM-related publications across governance and safety topics.
3.1 Research Objectives
The review was guided by the following four specific questions:
  • RQ1: What are the key trust, risk, and security challenges posed by LLM-enabled Agentic AI systems? How do current approaches establish and measure the trustworthiness of Agentic AI outputs and decisions?
  • RQ2: What technical and governance strategies have been proposed to address these challenges?
  • RQ3: How do existing approaches map onto the TRiSM pillars: explainability (including trustworthiness), model operations, security, and privacy?
  • RQ4: What gaps remain in current research, and what directions are promising for future work?
3.2 Data Sources and Search Strategy
To conduct a thorough literature review on Agentic AI and TRiSM, we systematically searched major digital libraries, including IEEE Xplore, ACM Digital Library, SpringerLink, arXiv, ScienceDirect, and Google Scholar. The search encompassed publications from January 2022 to May 2025, with select seminal works prior to 2022 included to provide foundational context. We utilized Boolean search queries with an expanded set of keywords to capture the interdisciplinary intersection of LLM-based multi-agent systems and TRiSM considerations. The search terms were structured as follows:
(“ Agentic AI” OR “multi-agent systems” OR “MAS” OR “multi-agent LLMs” OR “AI agents” OR “autonomous agents” OR “intelligent agents” OR “collaborative AI” OR “distributed AI” OR “LLM-based agents” OR “agent coordination” OR “agent-based modeling” OR “swarm intelligence”)
AND (“trust” OR “trustworthiness” OR “risk” OR “security” OR “safety” OR “governance” OR “oversight” OR “compliance” OR “explainability” OR “interpretability” OR “transparency” OR “privacy” OR “data protection” OR “robustness” OR “accountability” OR “ethical AI” OR “fairness” OR “adversarial attacks” OR “regulatory compliance” OR “bias mitigation” OR “system reliability”)
This comprehensive search strategy ensured the identification of relevant studies addressing the integration of TRiSM principles with LLM-based multi-agent systems, particularly in high-stakes domains such as healthcare, science, and finance.
3.3 Inclusion and Exclusion Criteria
We applied the following criteria during the screening process:
  • Inclusion:
    • Studies published between 2022 and 2025, to capture developments following the introduction of LLMs (mostly post-ChatGPT).
    • Studies that explicitly discuss Agentic AI, Multi-Agents, or LLM-powered AI systems in the context of at least one TRiSM dimension: trust, risk, security, governance, explainability, or privacy.
    • Peer-reviewed articles, preprints, or reports from credible sources (e.g., arXiv, NeurIPS, IEEE, ACM, Nature, or governmental standards bodies).
  • Exclusion:
    • Papers focused solely on traditional, rule-based, or symbolic agents without any integration of LLMs or emergent coordination.
    • Studies addressing only low-level ML components (e.g., training optimization, architecture design) without reference to agentic behavior or TRiSM concerns.
    • Non-English papers or those lacking sufficient metadata (e.g., missing abstracts or publication details).
After screening titles and abstracts, we shortlisted 250 papers. A subsequent full-text review narrowed the selection to 180 primary studies that directly addressed one or more TRiSM pillars. In addition to peer-reviewed literature, we also incorporated relevant technical reports, blogs, and website links to contextualize and map the emerging landscape of Agentic AI within the TRiSM framework. An analysis on the literature is shown in Figure 2.
3.4 Quality Assurance
To ensure the reliability of included studies, we conducted a quality assessment adapted from established systematic review guidelines [43, 42]. Each paper was evaluated against the following criteria: (1) Does the study clearly define its objectives in the context of agentic or LLM-based systems? (2) Are the methods, experiments, or architectural designs well-documented and reproducible? (3) Does the paper substantively engage with at least one of the trust, risk, security, explainability, or privacy aspects? (4) Does the study offer empirical, theoretical, or normative contributions relevant to Agentic AI governance? Each criterion was rated as low, medium, or high. Studies scoring “low” in more than two dimensions were excluded or flagged for contextual relevance only. Quality ratings were assigned independently by four reviewers, with discrepancies resolved through discussion.
4 Fundamentals of Agentic AI Systems
In this section, we present the fundamentals of Agentic AI systems.
4.1 Agentic AI vs. Traditional AI Agents
Agentic AI systems fundamentally transform the design of autonomous systems, and are different from traditional AI agents. Traditional agents operate through predefined rules [44], heuristic workflows [45, 46], or deterministic logic [47], and perform well in narrow, well-defined environments. Cognitive Agents [48], a subset of traditional agents, model human-like reasoning by integrating perception, memory, and decision-making, often using architectures like SOAR [49] or ACT-R [50]. However, they remain constrained by static knowledge bases and limited adaptability [48].
In contrast to traditional AI agents, Agentic AI systems leverage foundation models (mainly LLMs) to achieve adaptive goal-oriented behavior through multi-agent coordination and emergent reasoning capabilities. Agentic AI systems fundamentally redefine autonomy through three core innovations: (1) Multi-agent coordination: Specialized agents, such as planners and verifiers, collaborate using structured protocols [17, 19]. (2) Persistent context: Memory architectures maintain task state across workflows, ensuring continuity [51]. (3) Dynamic meta-orchestration: Systems delegate tasks and resolve conflicts dynamically [52].
This Agentic AI architecture supports longitudinal task execution and cross-domain generalization, overcoming the context fragmentation inherent in traditional approaches [53]. For example, systems like AutoGen [18] enable LLM-backed agents to decompose complex tasks, such as software development, through multi-step reasoning [54]. This capability of Agentic AI represents a significant shift from reactive to proactive problem-solving systems [19]. A comparison of traditional AI agents and Agentic AI systems is presented in Table 2.
Table 2: Traditional AI Agents vs. Agentic AI Systems
Dimension Traditional AI Agents Agentic AI Systems
Autonomy Model × Reactive or deliberative (e.g., cognitive architectures) Goal-driven planning & adaptation
$\bullet$ Fixed action sequences or symbolic reasoning $\bullet$ Recursive self-improvement
Cognitive Foundation × Symbolic logic, FSMs, or cognitive models Foundation models (LLMs/LIMs)
$\bullet$ Hand-coded or structured KBs $\bullet$ Emergent reasoning
Intelligence Scope × Narrow & task-specific Broad & compositional
$\bullet$ Single-domain focus $\bullet$ Cross-domain transfer
Reasoning Approach × Deterministic or multi-step symbolic Multi-step & contextual
$\bullet$ Rule-based or production rules $\bullet$ Chain-of-Thought (CoT)
Collaboration × Isolated or limited coordination Role-specialized coordination
$\bullet$ Manual decomposition $\bullet$ Automated hierarchy
Temporal Context × Episodic, often stateless Persistent memory
$\bullet$ Session-bound or episodic memory $\bullet$ VectorDB/LTM
Orchestration × Hard-coded or fixed workflows Dynamic meta-agents
$\bullet$ Sequential or predefined goals $\bullet$ Conflict resolution
Tool Utilization × Static or domain-specific tools Planned API invocation
$\bullet$ Handcrafted interfaces $\bullet$ Function calling
Context Awareness × Bounded or heuristic context Memory-augmented
$\bullet$ Limited or cognitive context $\bullet$ RAG architectures
Learning Mechanism × Limited or rule-based learning Emergent learning
$\bullet$ Fixed rules or chunking $\bullet$ Fine-tuning/RL
Human-AI Interaction × Scripted or cognitive modeling Natural language interfaces
$\bullet$ Limited or human-like reasoning $\bullet$ Conversational adaptability
Exemplars $\triangleright$ MYCIN  $\triangleright$ ELIZA  $\triangleright$ SOAR  $\triangleright$ ACT-R $\triangleright$ AutoGen  $\triangleright$ ChatDev  $\triangleright$ MetaGPT  $\triangleright$ AgentVerse
= Core strength  × = Fundamental limitation  $\bullet$ = Implementation characteristic  $\triangleright$ = Modern framework  $\triangleright$ = Classical system
4.2 LLM-Based Agentic AI System Architecture
Agentic Multi-Agent System (AMAS) represent an emerging paradigm in AI where multiple LLM-powered agents operate semi-autonomously, interact with external tools, and collaborate to achieve complex tasks. As illustrated in Figure 3, a typical AMAS architecture comprises several key components that together form a flexible yet highly dynamic ecosystem. At the core are multiple LLM-Based Agents, each capable of reasoning, planning, and tool invocation [55]. These agents access a Shared Toolchain Interface [56] to execute code, perform searches, or interact with domain-specific APIs.
Figure 3: Architecture of LLM-based Agentic Multi-Agent System (AMAS), illustrating key functional layers: Perception/Input Layer (text, image, audio processing), Cognitive/Reasoning Layer (goal-setting, planning, decision-making), Action/Execution Layer (digital and physical task execution), Learning Module (supervised and reinforcement learning), Communication and Collaboration (agent messaging and coordination), Data Storage & Retrieval (centralized/distributed databases), and Monitoring & Governance (ethical oversight, observability, and compliance mechanisms). The modular design highlights adaptive intelligence and inter-agent synergy.
Communication and coordination are facilitated through a Communication Middleware [57] that allows agents to share goals, observations, or intermediate results. A Task Manager or Orchestrator [58] governs high-level planning and delegates subtasks to agents based on their roles or specializations. Agents can read from and write to a World Model or Shared Memory [33], which stores contextual knowledge, system state, or evolving task data. Human oversight is supported through a Human-in-the-Loop Interface [46], enabling users to prompt, correct, or halt agent behavior.
To ensure accountability, a Trust and Audit module monitors agent actions, logs tool usage, and generates behavioral traces [59]. However, this modular and distributed structure introduces significant TRiSM [31] challenges. With multiple autonomous agents accessing external resources, the Security Gateway becomes critical for enforcing access controls, authentication, and sandboxing [60].
Likewise, a dedicated Privacy Management Layer is essential to prevent leakage of sensitive or personally identifiable information [61], especially when data traverses multiple agents or tools. Finally, an Explainability Interface must provide interpretable rationales for multi-agent decisions, supporting transparency and trust calibration [62]. Together, these architectural elements make AMAS powerful yet complex, raising unique and urgent questions about how to ensure their trustworthiness, mitigate systemic risks, and secure them against adversarial behaviors. Below, we discuss the architecture of AMAS.
Language Model Core (Agent Brain). At the center of an Agentic AI system lies a LLM serving as the primary decision-making controller or “brain” [63]. The core LLM is initialized with a user goal and a structured agent prompt (defining its role, capabilities, and tool access). It then generates step-by-step decisions or actions, interpreting instructions, producing reasoning traces, and selecting next steps in either natural language or structured action formats. In many agent frameworks, such as AutoGPT [64], Baby AGI [65] and GPT Engineer [66], the LLM governs the full control loop, orchestrating the overall system behavior.
Planning and Reasoning Module. To handle complex goals, an explicit planning mechanism decomposes tasks into manageable sub-goals. This can be done internally via chain-of-thought (CoT) or tree-of-thoughts prompting (ToT) [67], where the model performs intermediate reasoning before arriving at a final decision. Reasoning is also one aspect of cognitive functionalities [68]. Some implementations employ external planning systems by translating goals into structured planning languages and using classical planners for long-term decision-making. Planning is often interleaved with execution and feedback: the agent refines its plan based on outcomes, alternating between reasoning, acting, and integrating observations. Techniques like Reasoning and Acting (ReAct) [69] exemplify this loop by guiding the agent through repeated reasoning–action–observation cycles, improving performance on complex tasks.
Memory Module. Agentic AI systems integrate memory to maintain context across iterations. This includes short-term memory (recent interactions held within the prompt context) and long-term memory (accumulated knowledge or experiences) [8]. Long-term memory is often implemented using vector databases, where key facts or past events are stored and retrieved by similarity search. By reintegrating past data into the LLM prompt, the agent can recall relevant information across sessions, avoid repetition, and support coherent long-term planning. Effective memory management enables adaptive, learning-driven behavior.
Table 3: Comparison of representative LLM-based Agentic AI frameworks across key design axes.
Framework (Year) Core LLM Planning Memory Tool Use Notable Features
AutoGPT [70] GPT-4 Self-looped CoT Vector DB OS shell + web Fully autonomous goal loop
BabyAGI [65] GPT-3.5/4 Task queue In-memory Web search Minimal task generator
GPT Engineer [71] GPT-4V Spec-to-code File cache Python REPL End-to-end code generation pipeline
LangGraph [72] Model-agnostic FSM via graph Persistent nodes Custom modules Visual orchestration of agent graphs
AutoGen [18] GPT-4 Multi-agent PDDL JSON/DB API calls Modular, reusable agent templates
MRKL [73] LLaMA/GPT Prompt router N/A Math + search tools Neuro-symbolic expert routing
Reflexion [74] GPT-3.5/4 Retry–reflect loop Episodic buffer Same as base agent Verbal self-improvement via reflection
MetaGPT [75] GPT-4 SOP workflow YAML state Git CLI Structured roles for software engineering
Voyager [76] GPT-4 Auto skill tree Task DB Minecraft API Lifelong open-ended learning in environments
WebVoyager [77] LLaVA-1.6 ReAct JSON store Browser actions Multimodal web interaction via vision and text
HuggingGPT [78] ChatGPT Task-plan-select Log store Hugging Face models External model orchestration by LLMs
CAMEL [79] GPT-4 Role-play CoT Dialogue log Chat only Multi-agent role simulation via dialogue
ChatDev [80] GPT-4 Chat chain File repo Unix tools Simulated software development workflow
CrewAI [81] Any Declarative plan Optional DB Python modules Lightweight, LangChain-free framework
AgentVerse [82] Model-agnostic Config graph Redis/KV store Plugin API Multi-agent simulation and task solving
OpenAgents [83] Model-agnostic Agent scripts MongoDB Web plugins Open platform with public hosting
SuperAGI [84] GPT-4 DAG workflow Postgres Pinecone Concurrent agents for production use
Tool-Use Interface. To extend its capabilities beyond text generation, the agent is equipped with a tool-use interface [85]. This layer allows the invocation of external tools such as web search, APIs, code interpreters, or databases. The available tools are specified in the agent prompt with command schemas. When the LLM determines a tool is needed, it emits a structured command, which is executed externally. The result is fed back into the LLM as a new observation. This mechanism enables the agent to access real-time information, perform computations, and interact with external systems dynamically. Tool-use frameworks like Modular Reasoning, Knowledge, and Language (MRKL) [73] illustrate this design by routing queries to different expert modules (symbolic or learned) with an LLM as the router. Similarly, approaches like Toolformer [85] train the model to insert API calls in its text generation.
Perception and Environment Interface. For agents interacting with dynamic environments, such as web interfaces, simulated worlds, or physical systems, an observation-action interface is essential [86]. Perception modules translate raw inputs (e.g., sensor data, images, or textual states) into representations the LLM can process. Conversely, the agent’s chosen actions are executed within the environment, and the resulting state changes are returned to the agent as observations. This loop supports sense-plan-act cycles that continue until the task is completed or halted. In robotic or multimodal settings, the interface may include additional sensory models (e.g., vision transformers), but the core flow remains consistent.
Integration and Autonomy. These modules together form a closed-loop architecture. The LLM plans and reasons over tasks, guided by memory and tools, and interfaces with the environment to execute decisions and observe results [16]. Each iteration updates the agent’s context, enabling self-refinement through dynamic feedback. This integrated design empowers Agentic AI systems to operate autonomously, pursue long-range goals, and exhibit adaptive behavior across dynamic environments.
Table  3 presents representative systems and shows how each one maps these five design axes (LLM core, planning/ reasoning, memory, tool-use, environment interface) into practice. For example, AutoGPT pairs a GPT-4 core with a self-looped chain-of-thought planner and a vector-database memory to serve as both tool layer and environment interface.
5 The TRiSM Framework for Agentic AI
The TRiSM Framework for Agentic AI is presented in this section.
Table 4: TRiSM pillars in Agentic AI systems with aligned governance references.
Pillar Primary Objective Key Practices & Agentic-LLM Considerations Exemplar Standards / Metrics
Explainability Transparent multi-agent decisions SHAP, LIME, counterfactuals; reasoning-trace logs; cross-agent dependency mapping NIST AI RMF “Explain” [87]
ISO/IEC 24029-1 [88]
Metrics: fidelity, sufficiency, human-trust score
ModelOps Lifecycle governance of agents, models, orchestration Version-controlled prompts; CI/CD w/ multi-agent sims; drift monitoring; rollback policies ISO/IEC 42001 (AI MS) [89]
SRE MTTR/MTTD; $\Delta$-accuracy (model drift); pipeline pass-rate
Application Security Prevent prompt injection, spoofing, lateral exploits Prompt sanitation; least-privilege scopes; encrypted comms; sandboxed tool calls; anomaly detectors OWASP Top-10 for LLM Apps [90]
Jailbreak-success%; mean-time-to-detect; exploit CVSS
Model Privacy Safeguard sensitive data and shared memory Differential-privacy training; data minimization; homomorphic encryption; secure enclaves; audit logs GDPR Art. 25 [91]
HIPAA §164 [92]
Metrics: $\epsilon$-DP budget, leakage rate, access-audit pass-rate
5.1 Governance for Agentic LLM Systems
AI TRiSM is a comprehensive governance framework designed to ensure that AI systems are trustworthy, robust, and compliant with safety standards [29]. Originally highlighted in industry guidelines for AI governance [93], TRiSM provides a structured approach to manage the unique challenges of LLM-based “ Agentic AI" systems (herein we refer to as Agentic LLM systems). The TRiSM framework addresses the trust concerns in Agentic AI by focusing on four key pillars: Explainability, Model Operations (ModelOps), Application Security, and Model Privacy [59, 29]. Figure 4 illustrates how these four TRiSM pillars map onto the core architectural layers of an LLM-based Agentic AI system.
Figure 4: TRiSM-Aligned Architecture for LLM-Driven Agentic AI Systems. The system is organised into four vertical layers: Trust & Explainability, ModelOps, Planning + Reasoning, and Model Privacy, each supported by a cross-cutting Monitoring & Governance Plane that enforces policies, logging, and compliance triggers. External tools are accessed through a single Secure API gateway, while the user interface feeds back into both the Trust layer and the governance plane to enable end-to-end oversight.
The TRiSM framework aligns closely with several key international AI governance efforts that shape best practices for safety, accountability, and transparency in AI systems. For example, the EU AI Act [94] (notably Article 14) mandates clear obligations for transparency, traceability, and human oversight in high-risk systems, principles directly reflected in TRiSM Explainability and ModelOps pillars. The NIST AI Risk Management Framework (AI RMF) [95] emphasizes Explainability, Governability, and Validity as core risk mitigation strategies, providing a conceptual match to TRiSM’s structured controls across lifecycle and system interaction points. Additionally, the OECD AI Principles advocate for transparency, robustness, and accountability [96]each echoed in TRiSM attention to secure tool invocation, memory traceability, and multi-agent auditability. These global frameworks reinforce the need for Agentic AI systems to support real-time monitoring, enforce role-based access, maintain verifiable logs, and provide intelligible outputs. As shown in Figure 5, TRiSM pillars provide operational scaffolding for implementing regulatory expectations in LLM-based multi-agent contexts.
Alignment of TRiSM Pillars with Governance Frameworks Explainability: Aligned with the EU AI Act (Art. 13–14) [94] and the NIST AI RMF “Explain” function [95]. This pillar promotes transparent, rationale, human-interpretable logging, and role-based interpretability [97]. ModelOps (Lifecycle Governance): Anchored in the NIST RMF “Govern” function [95] and ISO/IEC 42001 [98]. It emphasizes lifecycle controls, model versioning, rollback policies, and CI/CD pipelines with safety gates. Application Security: Guided by OWASP LLM Top-10 [90] and OECD AI Principles on Robustness [96], it covers access controls, prompt sanitization, sandboxing, and audit trail mechanisms. Model Privacy: Aligned with GDPR (Art. 25) [91], HIPAA §164 [92], and OECD AI Principles [96], this pillar ensures data minimization, encryption, differential privacy budgeting, and access-audit pass thresholds.
Figure 5: Governance alignment of TRiSM pillars with international frameworks such as the EU AI Act, NIST RMF, ISO/IEC 42001, and OECD Principles. Each pillar operationalizes specific compliance strategies to ensure transparency, resilience, and trust in LLM-based systems.
5.2 Explainability in Multi-Agent Decision Making
Explainability refers to making the inner workings and decisions of AI agents interpretable to humans. In the context of AMAS, explainability is paramount for building user trust, as outcomes often emerge from complex interactions among multiple agents rather than a single model’s prediction [99]. However, achieving explainability in multi-agent LLM systems is challenging because one must interpret not only individual model decisions, but also the inter-agent dynamics that lead to final outcomes.
Established explainable AI techniques provide a starting point. For instance, LIME [100] and Shapley Additive Explanations (SHAP) [101] can be adapted to analyze LLM decisions. These techniques identify which features or input factors most influenced an agent’s output, offering insight into why a particular action or response was taken. In an Agentic LLM setting, a “feature” might be a component of the agent’s input context or a signal from another agent.
Beyond local explanations, counterfactual analysis is increasingly important for multi-agent explainability. Counterfactual techniques examine how the system’s behavior would change if certain conditions were altered [102], for example, if a particular agent’s contribution were removed or modified. This approach, rooted in causal inference, helps isolate each agent’s role in collaborative decision-making. Recent research on explainable AMAS suggests logging the intermediate reasoning steps (e.g. CoT prompts or dialogue between agents) can further enhance transparency. Human auditors can also help with a trace of how agents arrived at a decision, e.g., which agent contributed what information and why, to produce narrative explanations for its outcomes.
5.3 ModelOps : Lifecycle Management for LLM Agents
ModelOps is the discipline of managing AI models through their entire lifecycle, from development and deployment to monitoring, maintenance, and eventual retirement [103]. It extends the principles of MLOps (Machine Learning Operations) to focus specifically on model governance and reliable operation in production. Within Agentic AI systems, ModelOps encompasses not just individual models, but the orchestration of multiple agents and the supporting infrastructure that keeps them running safely [59]. Effective ModelOps is important to maintain consistency, performance, and regulatory compliance as LLM agents evolve or are added to a system.
Agentic AI systems require rigorous lifecycle governance because their behavior can change with model updates, prompt adjustments, or environmental variations. A cornerstone of ModelOps in this context is version control, i.e., tracking and managing versions of each agent’s model and prompt configurations. Additionally, robust CI/CD pipelines (Continuous Integration/Continuous Deployment) are employed to automatically test agents’ performance and safety whenever a model is fine-tuned or an agent’s logic is modified. Before deployment, multi-agent simulations and unit tests validate that new agent behaviors do not introduce regressions or unsafe interactions. This aligns with best practices in LLM operations (LLMOps) [104], which integrate MLOps principles tailored to the challenges of LLMs. In line with TRISM objectives, robust ModelOps thereby ensures that an Agentic AI system remains reliable and maintainable.
5.4 Application Security for LLM Agents
To mitigate threats such as prompt injection, spoofing, and impersonation, the TRiSM security pillar recommends a defense-in-depth approach tailored to LLM agents. This includes prompt hygiene [25] , where agents treat inputs defensively by sanitizing and filtering prompts, and applying guardrails or content policies to detect and refuse suspicious instructions. Techniques like prompt hardening, such as secure prefixes or validation steps, can help reduce susceptibility to injection attacks [105]. It also involves strong authentication and access controls, ensuring that each agent and human user is securely authenticated, and that least-privilege principles are enforced to limit autonomous agent actions [106]. Finally, continuous monitoring plays a critical role: if an agent begins issuing anomalous requests or deviates from typical behavior patterns, automated systems can flag the incident for review or trigger a shutdown to prevent further risk.
Recent frameworks, such as LangChain [107], AutoGen [108], and CrewAI [81] introduce the idea of trust scores or reputation among agents, where agents verify each other’s outputs and cross-check decisions to catch inconsistencies or signs of compromise. Moreover, training LLM agents with adversarial robustness in mind (e.g. fine-tuning on adversarial examples, employing adversarial training regimes) [25] can improve their resistance to malicious inputs.
5.5 Model Privacy and Data Protection
The Model Privacy pillar addresses the protection of sensitive data within AI agent systems, ensuring that the use of personal or confidential information complies with privacy regulations and ethical norms [27]. Since multiple agents may share information with each other (e.g. via a shared memory store or message passing) to collaborate. Without strict privacy controls, there is a risk that an agent could inadvertently expose private data to unauthorized parties or that sensitive information could “leak” through the language model’s outputs. Therefore, TRiSM’s privacy pillar recommends organizations to institute measures that safeguard data throughout the AI lifecycle, from training and inference to inter-agent communication.
To preserve privacy in multi-agent LLM systems, several technical measures are employed. Differential Privacy (DP) adds calibrated noise during model training to ensure that no individual data point can significantly influence the output, thus preventing memorization of personal information [109]. Data anonymization and minimization further enhance privacy by limiting shared information to only what is essential, often in aggregated or pseudonymized formats, such as using “age bracket 30–40” instead of precise birthdates [110]. Secure Multi-Party Computation (SMPC) allows agents to jointly compute functions without revealing their private inputs, making it particularly useful in cross-organizational scenarios like collaborative fraud detection [111]. Similarly, Homomorphic Encryption (HE) permits computation directly on encrypted data; with Fully Homomorphic Encryption (FHE), even plaintext queries and responses remain inaccessible to the agent itself [112]. Trusted Execution Environments (TEEs), or secure enclaves, offer hardware-based isolation, ensuring that even privileged users cannot access data during processing, which is an essential safeguard for secure memory sharing and execution [113]. Finally, the implementation of model privacy policies and compliance frameworks—including data retention limits, audit logs, and adherence to regulations such as GDPR or HIPAA, ensures that agents operate within legal and ethical boundaries [114].
Table 4 condenses each pillar into its governing goal.
6 Threats and Risks in LLM-based Agentic Multi-Agent Systems (AMAS)
In this section, we present the main threats in Agentic AI and also discuss a taxonomy of emergent risk classes.
Figure 6: Overview of Key Risks and Threats in Agentic AI Systems. The left panel outlines system-level threats unique to agentic architectures, including Autonomy Abuse, Persistent Memory, and Agent Orchestration. The right panel presents a taxonomy of emergent risk classes : Adversarial Attacks, Data Leakage, Agent Collusion, and Emergent Behaviors that arises from multi-agent interactions, memory persistence, and decentralized decision-making.
6.1 Threats in Agentic AI
Agentic AI systems introduce a distinct set of security and reliability concerns compared to traditional single-agent LLM architectures [115]. Below, we discuss these threats.
Autonomy abuse. The foremost threat is autonomy abuse, where agents with significant decision-making authority might misinterpret objectives or implement harmful plans due to erroneous reasoning or manipulated inputs [116]. Unlike deterministic models, agentic systems dynamically generate actions, complicating efforts to define and enforce safe operational states [117].
Persistent memory. Another threat is the persistent memory, which, while crucial for context retention, introduces unique vulnerabilities through potential adversarial injections [118, 119]. Such contamination can propagate subtly via shared memory, especially in the absence of detailed version control and robust audit mechanisms [116].
Agent orchestration. This risk involves central or distributed control mechanisms for role assignment and workflow mediation. A compromised orchestrator could distort task distribution or misroute information, triggering cascading failures, issues exemplified by documented vulnerabilities in MetaGPT [19] and AutoGen [18]. These orchestration vulnerabilities differentiate Agentic AI systems from conventional stateless, single-threaded LLM deployments.
6.2 Taxonomy of Risks
We propose a taxonomy of risks associated with autonomous multi-agent systems (AMAS), categorizing them into four primary classes: adversarial attacks, data leakage, agent collusion, and emergent behaviors, as shown in Figure 6. These categories encapsulate the diverse challenges arising from the complex interactions, persistent memory, and autonomous decision-making inherent in AMAS. Below, we delineate each risk category, highlighting their mechanisms, implications, and illustrative examples.
Adversarial Attacks. Autonomous agents are susceptible to a range of adversarial manipulations, including prompt injections, gradient-based attacks, and engineered reasoning traps. These vulnerabilities are exacerbated in AMAS due to the cascading effects of agent interactions, where a single compromised agent can propagate malicious influences across the system [120, 121]. For instance, the role-swapping attack observed in ChatDev demonstrates how adversaries can exploit inter-agent dependencies to subvert system objectives, leading to unauthorized access or manipulated outputs [17]. Such attacks underscore the need for robust defenses, including adversarial training and real-time monitoring, to mitigate risks in dynamic, multi-agent environments.
Data Leakage. The integration of persistent memory and extensive inter-agent communication in AMAS heightens the risk of unintended disclosure of sensitive information [122, 123]. In domains such as financial services and human resources, where confidentiality is paramount, inadequate boundary enforcement and ineffective data sanitization protocols can lead to significant breaches [124, 125]. For example, an agent processing proprietary financial data may inadvertently share sensitive information with other agents or external systems due to insufficient access controls. Addressing data leakage requires rigorous data isolation mechanisms, encryption, and continuous auditing to ensure compliance with privacy regulations.
Agent Collusion and Mode Collapse. Coordination mechanisms designed to enhance collaboration among agents can inadvertently foster undesirable outcomes, such as groupthink or mode collapse, where agents reinforce mutual errors or converge on suboptimal solutions [126, 127]. This phenomenon, often termed agent collusion, arises when agents prioritize consensus over critical evaluation, leading to echo chambers that amplify biases or inaccuracies. For instance, in multi-agent systems tasked with decision-making, repeated reinforcement of flawed assumptions can degrade system performance. Mitigating these risks necessitates diversity in agent design, robust feedback loops, and mechanisms to detect and correct convergent errors.
Emergent Behaviors. The complex interplay of agents, memory components, tools, and tasks in AMAS can give rise to unpredictable behaviors that elude conventional testing and validation frameworks [128, 129]. Such emergent behaviors may manifest as unintended shortcuts, where agents optimize for efficiency at the expense of critical verification processes, or as the suppression of contradictory information, compromising system reliability. Notable examples include blockchain-based systems, where agents bypassed consensus protocols to expedite transactions [130], and audio verification systems, where agents overlooked subtle discrepancies in authentication signals [131]. These cases highlight the need for adaptive monitoring to anticipate and address emergent risks.
6.3 Real-World Use Cases
Several real-world and research-based examples illustrate the tangible impact of these risks in deployed or experimental Agentic AI systems.
Prompt Leakage in Agentic AI Systems. Instances of prompt leakage have been observed in LLM-based agent frameworks such as AutoGPT [132], where recursive prompt augmentation and insufficient memory controls led to the unintentional exposure of sensitive information. In one reported scenario, sensitive tokens were stored in persistent memory and later surfaced in planning summaries or external logs [133]. Such vulnerabilities underscore the critical importance of implementing memory sanitization, access controls, and prompt boundary protections to safeguard Agentic AI systems from cascading information leaks [134].
Collusive Failure in ChatDev. In a collaborative code generation session involving planner, coder, and tester agents within the ChatDev framework, an error in a shared planning module led to the propagation of a faulty design [80]. Due to the absence of external ground-truth or objective feedback loops, all agents validated each other’s outputs, resulting in a feedback loop of erroneous confirmations. This scenario highlights the necessity of incorporating diverse information sources and adversarial checks within agent loops to prevent such collusive failures [24].
Simulation Attack in Swarm Robotics. In a simulated swarm robotics experiment utilizing LLM-based planning strategies, an agent was provided with a misleading environmental assumption, leading to a coordination failure characterized by spatial congestion and task incompletion [135]. This incident demonstrates the potential vulnerabilities in real-world deployments, particularly in critical infrastructure or logistics, where such failures could have significant consequences. The case highlights the importance of robust validation mechanisms and the integration of diverse information sources to ensure reliable swarm behavior [136].
Memory Poisoning in Multi-Agent Chatbots. In a multi-agent customer support system, a customer-facing agent injected sarcastic feedback into a persistent feedback buffer. This buffer was later utilized by the policy improvement agent to adapt dialogue strategies, resulting in responses with inappropriate tones. This incident highlights the importance of implementing validation filters, sentiment monitoring, and robust feedback loop governance in self-adapting systems to prevent such memory poisoning vulnerabilities [137].
System Prompt Drift in Autonomous Memory Agents. In experiments with agents using system-level memory (e.g., LangGraph or BabyAGI), over time, system prompts began drifting due to self-appended contextual memory that wasn’t properly versioned or validated. This led to hallucinated goals and emergent behaviors misaligned with initial intentions [138].
7 Trust and Explainability in Agentic AI
Table 5: Trust-Enabling Mechanisms in Agentic AI Systems
Mechanism Purpose Example System(s) Trust Contribution
Transparency via Reasoning Traceability Make agent decisions observable through intermediate steps or explanations SciAgent [139], MetaGPT [140] Enhances interpretability and debuggability
Status Reporting and Intent Disclosure Provide real-time updates or clarification of progress and errors ChatDev [17], AutoGen [108] Mimics human team transparency; reduces uncertainty
Human-in-the-Loop (HITL) Enable user oversight over high-stakes or irreversible actions ChemCrow [141], enterprise dashboards Prevents unsafe execution; reinforces control
Behavioral Consistency & Bounded Autonomy Ensure agents operate within predefined roles and limits Adaptive tutors [142, 143] Improves predictability; prevents overreach
Social Trust Cues Use polite language, uncertainty and expressions GPT-4 safety-tuned profiles [144, 145] Enhances user comfort and trust in uncertain scenarios
Agentic AI systems are highly autonomous agents capable of making decisions and taking actions without continuous human oversight. These systems also pose unique challenges and opportunities for human trust. Two key factors that influence trust in Agentic AI are the transparency of the agent’s decision-making processes and the ability to explain or justify its actions in human-understandable terms. We show the relationship between trust mechanisms, explainability, and evaluation facets in Figure 7 and discuss next.
7.1 Building Trust in Autonomous Agents
In Agentic AI systems, building trust is foundational to user acceptance, system reliability, and responsible deployment, especially as these systems begin to make autonomous decisions in critical domains such as healthcare, finance, and scientific research [146]. Unlike traditional software agents, autonomous LLM-based agents are characterized by self-directed reasoning, adaptive memory, and dynamic collaboration, which make their operations opaque and often unpredictable [147]. Establishing trust in such systems, therefore, requires a combination of technical transparency, user feedback integration, and robust oversight mechanisms.
Trust-Enabling Mechanisms Transparency • Status Reporting • HITL • Bounded Autonomy • Social Cues Explainability Techniques LIME • SHAP • Decision Provenance • Attention Maps • Prompt Audits Evaluation Facets Trustworthiness • Explainability • User-Centric • Coordination • Composite
Figure 7: Relationship between trust mechanisms, explainability techniques, and evaluation facets in Agentic AI.
Transparency and Decision Traceability. Transparency is one of the core enablers of trust. For users to understand and evaluate agentic decisions, the reasoning chains, decision states, and action triggers of agents must be made observable. Several Agentic AI systems are now integrating decision traceability through mechanisms such as CoT prompting and self-explanation modules. For example, SciAgent [139] generates scientific summaries and provides justifications by linking outputs to source documents retrieved via retrieval-augmented generation (RAG). Similarly, MetaGPT[140] structures its reasoning using role-based outputs, where each agent (e.g., planner, coder) explicitly states the logic behind its task execution, creating modular interpretability.
Status Reporting and Progress Visibility. Clear reporting of intent and intermediate status is also essential. Human collaborators often require updates about what an agent is doing, why a task is taking longer, or how an agent interprets ambiguous instructions. Tools like AutoGen[108] and ChatDev[17] have incorporated structured chat interfaces where agents summarize their intermediate progress, decisions, and encountered errors.
Human-in-the-Loop Oversight. Human oversight and intervention mechanisms further reinforce trust. Allowing human users to review, edit, or approve agent-generated outputs not only prevents missteps but also signals that the system respects user authority [148, 149, 150]. Many systems adopt a human-in-the-loop (HITL) paradigm [151], where agents request confirmation before executing high-risk or irreversible actions.
Behavioral Consistency and Bounded Autonomy. Trust requires predictability. Agents should follow defined roles, output in expected formats, and remain within delegated authority. For example, in enterprise AI platforms used for automated data analytics, agents may generate insights or dashboards but defer publishing until a domain expert reviews the material. Similarly, in autonomous research assistants like ChemCrow [141], agents pause to allow chemists to validate proposed reactions or data pipelines before proceeding, reinforcing safe deployment.
Social Trust Cues and Language Behaviors. Beyond system-level mechanisms, behavioral consistency and guided autonomy are crucial. If an agent behaves unpredictably or inconsistently, even if technically correct, users are less likely to trust it. Behavioral alignment mechanisms such as predefined role protocols can help output style consistency [152, 16, 153]. For example, in adaptive educational platforms using AI tutors, agents may be allowed to revise lesson plans but not change grading criteria, preserving institutional trust boundaries [142, 143]. Also, social trust cues, such as polite language, turn-taking, and cooperative gestures, have shown promise in reinforcing user trust even in non-expert settings [154]. Studies have found that users trust agents more when they express uncertainty ("I’m not sure, but here’s what I found") rather than overconfidence [144, 145]. This has been implemented in models like GPT-4 when configured with safety-tuned instruction sets, improving reliability perception without undermining capability. These mechanisms form a layered trust strategy for Agentic AI, as summarized in Table 5.
7.2 Explainability Techniques in AMAS
Emergent behavior of AMAS poses a challenge: the interpretability of an individual agent does not necessarily imply the interpretability of the overall system. In platforms such as ChatDev [17] or AutoGen [18], agents emulate specialized roles (e.g., engineer, reviewer), and tracing a final action back to its source agent is often non-trivial. While local explainability techniques such as LIME [100] and SHAP [101] remain foundational, their application in Agentic AI is limited due to decentralized agent contexts and shared memory. Researchers have thus proposed composite frameworks that combine local attribution with global decision traceability, such as decision provenance graphs [155], causal influence chains [156], and multi-agent SHAP adaptations [157]. These techniques aim to surface the system-level dynamics that underlie emergent agent behaviors.
Symbolic and Hybrid Architectures. Another promising direction is the use of inherently interpretable modules such as rule-based planners and decision trees within hybrid architectures [5]. These agents offer built-in explainability but retain the generative capabilities of LLMs for broader context understanding. Such designs are increasingly used in domains where structure and interpretability are prioritized, such as educational AI or mission planning.
Lightweight Interpretability: Attention and Prompts. Attention map visualizations have been used to highlight focus areas in multi-modal language agents [158, 159, 160], offering lightweight but informative views into model behavior. Prompt audit trails logging prompt history, agent actions, and response metadata have also gained traction [161, 162]. These mechanisms support system debugging, safety evaluations, and human-in-the-loop fine-tuning in multi-agent environments.
Table 6: Explainability Techniques in Agentic AI Systems
Technique Purpose Challenges in Agentic AI Example Use Cases
LIME [100] Local post-hoc explanation using surrogate models Fails to capture cross-agent dependencies in multi-agent setups Feature-level auditing in fraud detection
SHAP [101] Attribution of prediction to input features via Shapley Difficult to aggregate across agents with divergent contexts Financial risk analysis, regulatory explanations
Decision Provenance Graphs [155, 163] Trace inter-agent decision paths and data flows Can be complex to interpret at scale Collaborative document generation, planning chains
Multi-agent SHAP [157] Attribution extended to agents and shared memory High complexity, lack of standard frameworks Strategic gameplay, collaborative writing
Symbolic/Rule-Based Agents [5] Built-in interpretability via logic or rule-based models Limited flexibility compared to LLMs Hybrid planning systems, educational tutors
Attention Maps [158, 160] Visualize focus of language/image models Only reflects internal focus, not logic or intent Multimodal VQA, image captioning
Prompt Audit Trails [161, 162] Log input prompts, outputs, and agent actions Requires infrastructure, may miss implicit state changes Debugging multi-agent sequences, fine-tuning
Despite these advancements, achieving robust explainability in Agentic AI systems remains an open research problem. Many techniques focus on isolated predictions or modules and fail to capture system-level dynamics. Future work should prioritize longitudinal interpretability across agent interactions, causal reasoning pipelines, and interactive querying interfaces that support transparency in real time. A comparative summary of current techniques is shown in Table 6.
7.3 Evaluation Metrics for Agentic AI Systems
Agentic AI systems demand a comprehensive evaluation across multiple dimensions beyond traditional accuracy. We outline five key categories of metrics: trustworthiness, explainability, user-centered performance, coordination, and composite scores, each capturing a distinct aspect of an Agentic AI system performance and its real-world implications. Below, we discuss these metrics and summarize in Table 7.
Trustworthiness. This criteria evaluates the reliability, safety, and ethical alignment of the AI agent. A trustworthy agent consistently produces correct and unbiased results, adheres to constraints, and avoids harmful or unpredictable behavior. Metrics for trustworthiness include success rates on tasks under varying conditions (measuring robustness), violation rates of safety or ethical guidelines (which should be minimal), and calibration of the agent’s confidence (how well the agent’s self-reported confidence aligns with actual accuracy) [164]. Some approaches combine such factors into an overall trust score. For example, one model defines a trustworthiness score $T$ as
$T=\frac{C+R+I}{S}\,,$ (1)
where $C$ is the agent’s credibility (accuracy and correctness of its outputs), $R$ is reliability (consistency of performance over time), $I$ is the level of user alignment or rapport (analogous to “intimacy” in trust modeling), and $S$ is self-orientation (the degree to which the agent pursues its own goals over the user’s goals). A higher $T$ indicates an agent that is accurate, consistent, user-aligned, and not self-serving, which corresponds to greater trustworthiness. In practice, achieving high trustworthiness means the agent behaves predictably and transparently in accordance with ethical AI principles (such as fairness and accountability).
Explainability. Explainability metrics assess how well the agent’s decisions can be understood and traced by humans [165]. These metrics focus on the clarity and completeness of the rationale the agent provides for its actions. For instance, one can measure the explanation coverage (the percentage of decisions or outputs that come with an adequate explanation) and the fidelity of explanations (how accurately the explanation reflects the true reasoning or model logic).
Consistency of explanations for similar scenarios is another important metric: the agent should explain comparable decisions in a similar way, indicating a stable reasoning process. Quantitatively, methods like OpenXAI [166] provide a suite of metrics to evaluate explanation quality across dimensions such as faithfulness, stability, and fairness of explanations. High explainability builds user trust, as users can follow why the agent made a decision, and it aids debugging by revealing the agent’s internal decision process. In regulated domains (e.g., healthcare or finance), explainability is often essential for compliance and user acceptance.
User-Centered Performance. User-centered metrics capture how effectively the AI agent interacts with and satisfies the end-user’s needs. These criteria emphasize the user’s experience and outcomes [167]. Key metrics include user satisfaction ratings, typically collected via surveys or feedback after interaction, which reflect whether the user’s goals were met and how comfortable they felt with the Agentic AI system behavior.
Additionally, interaction metrics like the number of back-and-forth clarification queries needed (fewer indicates the agent understood the user well) and the coherence or naturalness of the conversational flow contribute to user-centered evaluation. Human-in-the-loop evaluations are often employed here: for example, user studies might rate the agent on criteria such as helpfulness, clarity and naturalness of language, and adherence to user instructions. Ultimately, a user-centered Agentic AI system should align its actions with user intent and preferences.
Table 7: Summary of key evaluation metrics for Agentic AI systems. Each category addresses a different facet of an autonomous agent’s performance and behavior.
Metric Aspect Evaluation Focus (Examples of Metrics)
Trustworthiness Reliability and safety of the agent’s behavior. Example metrics: task success rate across diverse scenarios; frequency of rule or safety violations (lower is better); calibration of confidence vs. outcomes; fairness and bias indices.
Explainability Transparency and interpretability of the agent’s decisions. Example metrics: percentage of outputs with an accompanying explanation; explanation fidelity to actual model reasoning; consistency of explanations for similar cases; human interpretability ratings of explanations.
User-Centered User satisfaction and alignment with user needs. Example metrics: user satisfaction score (post-interaction survey); rate of fulfilling user-defined goals; number of clarification questions needed (lower is better); naturalness and coherence of dialog from the user’s perspective.
Coordination Effective collaboration in multi-agent or modular systems. Example metrics: multi-agent task completion rate; communication overhead (e.g., messages exchanged); consistency of shared plans or beliefs among agents; synergy score quantifying complementary actions.
Composite Overall performance across metrics. Example metrics: weighted aggregate score encompassing trust, explainability, user-centric, and coordination measures; specialized composite indices (e.g., TUE for tool-use efficacy); cross-domain benchmark results (aggregate success across diverse tasks, as in AgentBench).
Coordination. In scenarios where an Agentic AI system consists of multiple cooperating agents or modular components, coordination metrics gauge how effectively these parts work together. Good coordination means agents share information, divide labor without conflict or redundancy, and converge on solutions efficiently. Quantitative measures include the team success rate on collaborative tasks (whether the group of agents achieves the overall goal) and communication efficiency metrics (e.g., the number of messages or iterations required among agents to reach a decision, with fewer often indicating more efficient interaction).
One specific example is the Component Synergy Score (CSS), which counts or weights effective interactions between agents, reflecting how well each agent’s actions complement the others (a higher CSS means agents are synergistic rather than working on individual basis). MAS such as ChatDev and MetaGPT provide practical testbeds for these metrics: they orchestrate specialized agents (e.g., different roles in a software engineering pipeline) that must cooperate to complete complex projects.
Evaluations on such frameworks examine whether agents maintain a consistent shared plan, handle inter-agent dependencies smoothly, and recover from misunderstandings. For instance, if one agent generates a plan and another executes it, a coordination metric would assess if the executing agent follows the planner’s intent correctly and whether both agents remain in agreement throughout the process. High coordination scores indicate that the agentic system functions as a cohesive unit, which is crucial for complex tasks beyond the capability of any single agent.
Composite Metrics. Composite metrics aggregate multiple evaluation aspects into a single overall score. These are useful for summarizing agents’ performance holistically, especially when comparing different systems. A composite metric is often a weighted combination of the above categories, for example:
$M_{\text{composite}}=w_{T}M_{T}+w_{E}M_{E}+w_{U}M_{U}+w_{C}M_{C}\,,$ (2)
where $M_{T},M_{E},M_{U},M_{C}$ are the normalized scores (on a common scale) for trustworthiness, explainability, user-centered performance, and coordination, respectively, and $w_{T},w_{E},w_{U},w_{C}$ are weights reflecting the relative importance of each aspect for a given application. The choice of weights $w_{i}$ can be domain-specific (for instance, in healthcare applications, trustworthiness and explainability might be weighted more heavily than raw efficiency).
An example of a specialized composite metric in agentic contexts is the Tool Utilization Efficacy (TUE) score, which combines sub-metrics evaluating how correctly and efficiently an agent uses external tools (including proper tool selection and correct parameter usage in tool calls) into one measure. By condensing multiple criteria, composite metrics enable high-level comparison and benchmarking of Agentic AI systems. For instance, AgentBench [168] is a comprehensive benchmark that evaluates agents across a diverse range of tasks and environments (from operating system manipulation to web shopping), effectively providing a composite performance profile of an agent. Such aggregated scores highlight if an agent performs strongly across the board or if it excels in some dimensions while underperforming in others. It is important to interpret composite scores in light of their components: a single number can mask specific weaknesses (e.g., an agent might achieve a high overall score by doing well in task completion and coordination, yet still have poor explainability). Therefore, composite metrics are most informative when accompanied by a breakdown of the agent’s performance per category.
8 Security and Privacy in LLM-based Agentic Multi-Agent Architectures
In this section, we discuss the security and privacy in AMAS.
Table 8: Governance Dimensions for Agentic AI Systems
Governance Component Description Example Tools or Methods
Regulatory Compliance Aligning with legal and industry standards for responsible AI [169] NIST AI RMF [87], EU AI Act [170], GDPR [91], HIPAA [92]
Auditability and Logging Capturing traceable records of agent actions and decisions [171] Decision provenance, immutable logs, blockchain-based traceability
Policy Enforcement Enforcing system-level operational, ethical, and security constraints [53, 172] RBAC/ABAC policies [173], formal logic rules, dynamic sandboxing [60, 174, 175]
Human Oversight Human-in-the-loop [151] decision checkpoints and override capabilities Human review agents, interactive dashboards, compliance checkpoints
Risk Monitoring Continuous identification and mitigation of systemic or emergent risks [176] Model risk scanners, out-of-distribution detectors, reinforcement learning audit modules
Explainability Governance Ensuring interpretability standards are met in decision pipelines [177] LIME/SHAP [178, 179, 180], decision trees [181, 182], interpretable surrogate models [183, 184, 185]
Adaptive Governance Updating governance mechanisms based on model evolution and deployment context [144, 146] Governance-as-code [114, 186], auto-updated rule engines [187, 188, 189], learning-based policy adaptation [126, 190]
Incident Response and Recovery Procedures for handling, logging, and recovering from failures or breaches Real-time alerting, kill-switch mechanisms, secure rollback protocols
8.1 Security Mechanisms
Agentic AI systems, composed of loosely coupled yet collaboratively functioning LLM-based agents, introduce an expanded attack surface relative to conventional AI agents [191]. Ensuring the security of such systems necessitates a multi-layered defense architecture that addresses data protection, execution integrity, inter-agent communication, and model robustness [192]. Among the foundational techniques employed are encryption [193, 194], access control [195, 171], adversarial defense [196, 197], and runtime monitoring[198, 199] each adapted to the unique demands of decentralized multi-agent environments.
Encryption. plays an important role in safeguarding data exchanged between agents, especially when sensitive or regulated content (e.g., healthcare records, financial data) is involved [194, 61, 200]. Agentic workflows often include inter-agent handoff of partially processed results, models, or prompts. Implementations such as SSL/TLS, homomorphic encryption [201, 200], and secure enclaves (e.g., Intel SGX) are increasingly integrated into Agentic AI pipelines to ensure confidentiality across message-passing protocols.
Access control. becomes crucial when orchestrators or shared memory modules manage permissions for agents with distinct capabilities and responsibilities. For instance, in systems like AutoGen[18] and CrewAI [81] where agents take on specialized roles (e.g., summarizer, planner, coder) [202], enforcing principle-of-least-privilege access prevents privilege escalation and unauthorized tool invocation [18, 198, 203]. Agent-based access control policies, often aligned with Role-Based Access Control (RBAC) [204, 205] and Attribute-Based Access Control (ABAC) [206, 205] paradigms, can dynamically restrict which agents may access sensitive APIs, files, or memory buffers, based on contextual trust levels [5].
Adversarial robustness. is a growing concern as LLM-based agents are susceptible to prompt injection, manipulation through poisoned tool outputs, or coordination disruption via malformed intermediate results [207, 208]. Recent studies have shown that multi-agent LLM frameworks can be destabilized by adversarially crafted outputs from one compromised agent propagating misleading information to others [208, 192, 115, 55, 116, 209]. Adversarial training methods, such as input perturbation, reward shaping, and contrastive learning, can partially mitigate these vulnerabilities [210, 115]. Integrating safety constraints and verifying tool responses before execution are also effective mitigation strategies.
Runtime monitoring. systems support the detection of anomalous agent behaviors, especially in high-stakes domains like automated healthcare or cybersecurity [211]. Log-based auditing, anomaly detection with LSTM or autoencoder-based detectors, and trust scoring among agents are becoming essential components of real-time surveillance layers [212, 213, 125]. For example, Microsoft’s Copilot governance layers monitor anomalous agent behavior across sessions to ensure compliant execution and flag potentially harmful interactions.
As Agentic AI systems continue to expand into mission-critical domains, developing standardized, scalable security mechanisms will be paramount. Future approaches must include zero-trust frameworks, secure multi-party computation, and formal verification of inter-agent protocols to ensure safe and resilient operation across decentralized autonomous agent collectives.
8.2 Privacy-Preserving Techniques
The decentralized and interactive nature of LLM-based Agentic AI systems introduces new challenges for preserving privacy, especially as agents continuously communicate, access external data sources, and store episodic or shared memory. To ensure data confidentiality and protect personally identifiable information (PII), Agentic AI systems must adopt robust privacy-preserving techniques such as differential privacy, data minimization, and secure computation.
Differential privacy (DP). DP offers mathematically grounded guarantees by injecting statistical noise into outputs, ensuring that individual user contributions cannot be re-identified. In multi-agent LLM systems, DP can be applied during training or at inference-time when agents exchange information. For example, Google’s implementation of DP in federated learning frameworks can be extended to distributed agentic systems, where agents collaboratively train or fine-tune local models without exposing raw data [214, 215]. DP-SGD and privacy budgets ($\epsilon$-differentials) can regulate information exposure during policy updates or collaborative planning in real-time decision-making agents.
Data minimization. Agentic AI systems can mitigate exposure risks by limiting the scope, granularity, and duration of data collected or retained during task execution. For instance, temporary memory buffers used in systems like ChatDev or ReAct-based pipelines are cleared once subgoals are completed, preventing persistent storage of unnecessary user data. Furthermore, anonymization and pseudony [215] techniques can help remove identifying features before data is passed between agents or stored in shared memory repositories.
Secure computation. Techniques including secure multi-party computation (SMPC), homomorphic encryption, and trusted execution environments (TEEs) enable agents to perform computations over encrypted or obfuscated data without compromising privacy. In scenarios where agents collaborate across different organizational boundaries (e.g., federated medical agents or cross-silo industrial agents), SMPC allows joint computations such as diagnostics or anomaly detection without data leakage. Homomorphic encryption, while computationally expensive, is increasingly being explored to allow arithmetic operations on encrypted vectors used in RAG workflows.
Privacy-by-design principles are becoming central to the engineering of next-generation agentic systems. Architectures now embed user consent layers, configurable privacy settings, and memory redaction modules that allow end-users or system administrators to control what agents can remember or share. As Agentic AI expands into domains such as personalized education, healthcare, and finance, ensuring privacy-respecting behaviors will be essential for regulatory compliance (e.g., GDPR, HIPAA) and public trust.
8.3 Compliance and Governance
As Agentic AI systems grow in capability and autonomy, ensuring regulatory compliance and instituting robust governance mechanisms become imperative. Unlike traditional AI agents, Agentic AI systems operate with greater autonomy, persistent memory, and complex decision flows necessitating layered oversight to manage legal, ethical, and societal implications. Effective governance in this context spans three critical dimensions: adherence to regulatory standards, system-level auditability, and enforceable policy frameworks.
Regulatory standards. Standards for regulation provide the baseline requirements that all AI systems, including agentic architectures, must meet. Frameworks such as the NIST AI Risk Management Framework (AI RMF) and the EU AI Act define principles for trustworthy AI including transparency, accountability, and fairness. These standards are especially relevant for LLM-based AMAS that interact in high-stakes domains like healthcare, finance, defense, or transportation. For example, the EU AI Act classifies certain autonomous systems as “high-risk,” requiring continuous risk monitoring of such systems.
Auditability. This is critical in ensuring transparency and facilitating post-hoc accountability. Each decision, plan, or interaction within an agentic system should be logged with timestamps, context, agent role, and justification. Techniques such as decision provenance and action traceability enable this, allowing regulators or internal auditors to reconstruct how decisions were reached. For instance, in systems like AutoGen or MetaGPT, where agents assume specialized roles (e.g., researcher, coder, reviewer), audit trails can capture role-specific actions and flag inconsistencies, bias amplification, or security violations. Blockchain-based audit logs are also being explored to ensure immutability and verifiability of multi-agent interactions.
Policy enforcement. Policy governs what Agentic AI system can and cannot do, and under which conditions. These policies must be codified into the orchestration layers or meta-agent governance modules that manage agent interactions. Examples include enforcing memory expiration policies to avoid data retention violations or restricting access to external tools based on role and authentication level. Role-based access control (RBAC) and attribute-based access control (ABAC) are essential for enforcing differentiated privileges across agent sub-components. Furthermore, real-time monitoring systems can halt or flag agent activity that deviates from pre-specified ethical constraints or operational bounds, using formal verification tools like TLA+ or symbolic execution engines.
Emerging best practices also include the creation of AI governance boards, the adoption of governance-as-code platforms, and the integration of adaptive governance layers that evolve as the Agentic AI systems scale or change context. These practices are designed to meet not only current standards like ISO/IEC 42001 for AI management systems, but also prepare for future regulatory evolutions. Below, we summarize the core governance components required for secure and compliant Agentic AI deployment in Table 8.
9 Discussion
We discuss the broader implications of our findings below:
9.1 Technical Implications of TRiSM for LLM Agent Design
Rather than treating LLM agents as black-box decision-makers, TRiSM encourages instrumenting them with continuous oversight “guardrails” [216]. For example, there is discussion on designing specialized “guardian agents” within AMAS [40]. In this paradigm, such agents serve as proactive monitors that filter sensitive data and establish baselines of normal behavior, while operator agents dynamically enforce policies at runtime (e.g. blocking disallowed actions such as outputting personally identifiable information).
Prior research highlights risks like “excessive agency” [106] where an LLM given too much autonomy or tool access can produce unintended harmful actions (for instance, via hallucination or misinterpreted goals). TRiSM-driven Agentic AI design mitigates these risks by limiting agent autonomy within well-defined safety bounds. Likewise, emerging threats specific to Agentic AI systems such as prompt injection [23] attacks, memory poisoning [137], or cascading hallucinations [217] highlights the need for more built in risk controls. By incorporating anomaly detection and policy-checking modules, an LLM agent can detect deviations from normal behavior and either alert humans or automatically neutralize the threat (e.g. masking a sensitive datum or stopping an unsafe action) .
9.2 Ethical and Societal Ramifications of Multi-Agent AI
Beyond technical matters, deploying networks of autonomous LLM agents raises pressing ethical and societal questions. Applying TRiSM in this context emphasizes principles of accountability, human oversight, and fairness [169], all of which are vital for public trust in AI systems. Governance in TRiSM recommends that organizations retain clear responsibility for their AI actions, and do not obscure blame behind algorithmic “black boxes”. This implies implementing audit trails and explicable decision logs so that any harmful or biased outcome can be traced and attributed [161]. Recent guidance on trustworthy AI [172] often highlights accountability and explainability as key pillars of trust.
In practice, this review sheds some light on how each autonomous agent decisions should be transparent enough to be understood and challenged by human reviewers when necessary [218]. Human oversight is another ethical imperative tightly coupled with accountability. TRiSM does not seek to eliminate humans from the loop; rather, it provides a structured way for humans and AI agents to collaborate under a defined governance framework [219].
In MAS setups, humans can monitor agent swarms in real time, pause or shut down agents exhibiting anomalies, and adjust policies on the fly [126]. The risk of “user complacency”, trusting an autonomous agent too much, has been noted as a hazard [220]. TRiSM governance counteracts this by formalizing oversight roles and ensuring no AI operates without appropriate human or regulatory supervision.
9.3 Alignment with Emerging AI Regulations and Standards
The principles embedded in TRiSM align closely with emerging regulatory frameworks for AI. This convergence means that adopting TRiSM-based governance can help organizations meet new legal obligations and industry standards. For example, the European Union’s AI Act [94] (set to fully apply in 2026) mandates rigorous risk management, transparency, data governance, and human oversight for “high-risk” AI systems. These are precisely the capabilities that a TRiSM approach promotes.
The EU AI Act also stresses accuracy, robustness, and cybersecurity for high-risk AI, qualities that TRiSM security management component is designed to ensure (through adversarial resilience, access control, etc.). Similarly, international AI governance standards are emerging that mirror TRiSM’s principles. ISO/IEC 42001:2023, the first global standard for AI management systems, highlights requirements such as transparency, accountability, bias mitigation, safety, and privacy in AI development. By implementing TRiSM, organizations essentially put in place the processes that ISO 42001 [89] and similar standards call for (e.g. leadership oversight, documented risk controls, ongoing monitoring and improvement cycles). Another example is the U.S. NIST AI Risk Management Framework [87], which emphasizes many of the same concepts: identifying risks, embedding governance, and cultivating trustworthiness in AI.
9.4 Limitations and Current Research Gaps
While the TRiSM-based approach appears promising, we also discuss some limitations and open challenges in current research. First, the classical AI domains (vision, NLP) that have standard test suites, there is no consensus on how to measure an “AI agent” ability to operate safely under TRiSM principles. This makes it difficult to compare different governance strategies or to track progress objectively. We encourage future work to develop evaluation frameworks, possibly extending from adversarial attack simulations [210] or “red-teaming” exercises [221] that can stress-test Agentic AI systems and score their resilience.
Secondly, there is a lack of real-world validation for many TRiSM-inspired controls. Much of the existing literature and tooling for LLM agent safety has been demonstrated in laboratory settings or on narrowly scoped tasks [31]. It remains uncertain how these governance mechanisms perform in complex, open-ended real-world environments.
Another critical gap is adversarial robustness. As we improve defenses, attackers will inevitably adapt [222]. Recent findings show that LLM-based systems remain vulnerable to attacks (for example, hidden prompt injections or subtle data poisoning) that can bypass superficial guardrails [25]. TRiSM solutions today are not foolproof against these. We identify evolving threats and adversarial attacks as ongoing obstacles to trustworthy AI [40].
Finally, organizational and human factors present limitations: implementing TRiSM requires interdisciplinary expertise (AI specialists, security experts, ethicists, legal advisors) and clear governance structures. Many organizations lack the necessary skill sets or frameworks, making TRiSM adoption superficial or inconsistent. Without a strong organizational commitment, even the best technical framework can falter.
9.5 Future Roadmap for Agentic AI TRiSM
Table 9: Future Roadmap for Agentic AI: Eight Strategic Domains
Scalable Agent Architectures Secure Multi-Agent Collaboration
Explainable Multi-Agent Decisions Human-Centered Trust Design
Autonomous Lifecycle Management Ethics & Governance
Benchmarking & Evaluation Cognitive Capability Expansion
Drawing from our findings and best practices from multiple disciplines, we propose several actionable directions for future research and implementation, as shown in Figure LABEL:fig:roadmap_mindmap. These recommendations span both technical system design improvements and governance-level policy initiatives:
Develop Standardized Evaluation Benchmarks. The community should create open benchmarks and challenge environments to test multi-agent AI governance. For instance, a suite of scenario-based tasks (with built-in threats and ethical dilemmas) could be used to evaluate how well a TRiSM-governed agent system performs relative to one without such controls. This will enable direct comparisons and drive progress on measurable metrics of trust (e.g. frequency of prevented failures or fairness outcomes).
Advance Adversarial Robustness Techniques. Future system design must anticipate a continually evolving threat landscape. Techniques from cybersecurity (e.g. adversarial training, AI model “penetration testing” [223], and formal verification) should be integrated into the LLM agent development pipeline. Cross-disciplinary collaboration with security experts can yield LLM-specific hardening methods, such as dynamic prompt anomaly detectors or robust tool APIs that constrain agent actions. Additionally, creating red-team/blue-team exercises for AMAS , akin to cyber wargames [224] , can help discover vulnerabilities in a controlled way before real adversaries do.
Human-Centered Oversight Tools. We encourage designing better interfaces and protocols for human oversight of Agentic AI systems. Borrowing from human-computer interaction [225] and cognitive engineering [226], researchers could devise dashboards that visualize an agent society’s state, flag important decisions, and allow intuitive human intervention (pausing agents, rolling back actions, etc.).
Regulatory Sandboxes and Compliance-by-Design: Policymakers and industry should collaborate to create regulatory sandboxes for multi-agent AI trials. These would be controlled environments where innovators can deploy Agentic AI systems under supervision, demonstrating TRiSM controls to regulators. Insights from such pilots can inform refinements in both technical standards and regulations. Moreover, adopting a compliance-by-design mindset is key: future AI system designs should bake in the requirements of frameworks like the EU AI Act and ISO 42001 from the start rather than retrofit them.
Cross-Domain Best Practices and Ethical Governance: There is much to learn from other high-stakes domains. For example, the safety engineering field (e.g. aerospace, automotive) has mature practices for redundant controls and failure mode analysis; these could inspire analogous practices in AI agent design. Likewise, ethics boards in biomedical research provide a template for AI ethics committees that review agent behaviors and approve high-risk deployments. We advocate establishing multidisciplinary governance boards that include ethicists, legal experts, domain specialists, and community representatives to oversee significant deployments of autonomous AI.
10 Conclusion
The rapid advancement of Agentic AI, driven by LLM-based multi-agent systems, marks a transformative shift in autonomous systems, enabling sophisticated collaboration, planning, and decision-making across high-stakes domains such as healthcare, finance, and public services. However, this evolution introduces complex challenges in trust, risk, and security management (TRiSM) that demand a tailored framework to ensure safe, transparent, and accountable operations. This review has systematically explored these challenges, proposing a TRiSM framework structured around four pillars: Governance, Explainability, ModelOps, and Privacy/Security, specifically adapted for LLM-based Agentic Multi-Agent Systems (AMAS).
Key contributions include a novel risk taxonomy capturing unique threats like prompt injection, memory poisoning, and collusive failures, alongside two new metrics: the Component Synergy Score (CSS) and Tool Utilization Efficacy (TUE), which provide practical tools for assessing inter-agent collaboration and tool efficiency. The review also highlights state-of-the-art techniques, such as LIME, SHAP, and decision provenance, to enhance explainability, while advocating for robust security measures like prompt hygiene, sandboxing, and homomorphic encryption to mitigate vulnerabilities.
Looking forward, the proposed research roadmap emphasizes scalable, verifiable, and regulation-compliant Agentic AI systems. Future work should prioritize advancing adversarial robustness, refining governance protocols, and developing standardized benchmarks for trustworthiness and coordination. By aligning Agentic AI development with TRiSM principles, researchers and practitioners can foster systems that are not only innovative but also responsible, ensuring their safe and ethical integration into society.
Declaration of Competing Interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Appendix
Table A.1: Key terminologies for LLM-based agentic AI systems.
Term Definition
Agentic AI System A multi-agent architecture powered by large language models (LLMs), where autonomous agents collaborate, plan, and execute tasks over extended horizons with persistent memory and dynamic role assignment.
Autonomy Model The mechanism by which an agent decides and acts without direct human intervention, often using goal-driven planning and chain-of-thought reasoning.
Chain of Thought (CoT) A prompting strategy in which an LLM generates intermediate reasoning steps before producing a final answer or action, enhancing interpretability and multi-step planning.
Counterfactual Analysis An interpretability technique that examines how altering certain inputs or agent contributions would change the overall system outcome, revealing causal dependencies among agents.
Explainability The capacity of an AI system (or individual agent) to produce human-understandable justifications or rationales for its decisions and actions, often via LIME, SHAP, or decision provenance graphs.
Foundation Model (LLM) A pretrained large language model (e.g., GPT-4, LLaMA) that serves as the “brain” of each agent, providing generative capabilities, reasoning, and tool-calling support.
Shared Memory (Persistent Memory) A centralized or distributed store (often a vector database) where agents write and retrieve contextual information, enabling long-term planning and consistency across iterations.
ModelOps The practice of managing AI models (and agent prompts) throughout their lifecycle—development, deployment, monitoring, and retirement—with version control, CI/CD testing, and drift detection.
Application Security Safeguards and best practices (e.g., prompt sanitation, authentication, sandboxing) designed to protect agentic systems from prompt injection, identity spoofing, and lateral exploits.
Model Privacy Techniques (e.g., differential privacy, homomorphic encryption, secure enclaves) that ensure sensitive data—either during training or inter-agent communication—remains protected in multi-agent workflows.
Prompt Injection A security exploit in which an attacker crafts input containing hidden instructions that corrupt an agent’s reasoning or propagate malicious commands through agent interactions (“prompt infection”).
Retrieval-Augmented Generation (RAG) A framework where agents query an external knowledge store (e.g., vector database) to fetch relevant documents or facts, then condition their LLM responses on that retrieved context.
Role-Specialized Coordination An architectural pattern in which each agent is assigned a specific function (e.g., planner, verifier, coder) and collaborates via structured communication protocols to achieve complex tasks.
Decision Provenance Graph A graph-based representation that traces data flows and decision steps across multiple agents, enabling post-hoc auditing and system-level interpretability.
Tool-Use Interface The mechanism by which an agent issues structured commands (e.g., API calls, code execution) to external services or environments and incorporates the results back into its reasoning.
Trust Score A composite metric that quantifies an agent’s reliability, alignment with user goals, and consistency over time, often combining accuracy, safety-violation rates, and calibration of confidence.
Composite Metric An aggregate evaluation score (e.g., a weighted sum of trustworthiness, explainability, user-centered performance, and coordination metrics) used to benchmark different agentic systems.
Figure A.1: Traditional AI Agent vs. Agentic AI.