### v3

# TRiSM for Agentic AI: A Review of Trust, Risk, and Security Management in LLM-based Agentic Multi-Agent Systems

Abstract
Agentic AI systems, built upon large language models (LLMs) and deployed in multi-agent configurations, are redefining intelligence, autonomy, collaboration, and decision-making across enterprise and societal domains. This review presents a structured analysis of Trust, Risk, and Security Management (TRiSM) in the context of LLM-based Agentic Multi-Agent Systems (AMAS). We begin by examining the conceptual foundations of Agentic AI and highlight its architectural distinctions from traditional AI agents. We then adapt and extend the AI TRiSM framework for Agentic AI, structured around four key pillars: Explainability, ModelOps, Security, Privacy and Governance, each contextualized to the challenges of multi-agent LLM systems. A novel risk taxonomy is proposed to capture the unique threats and vulnerabilities of Agentic AI, ranging from coordination failures to prompt-based adversarial manipulation. To support practical assessment in Agentic AI works, we introduce two novel metrics: the Component Synergy Score (CSS), which quantifies the quality of inter-agent collaboration, and the Tool Utilization Efficacy (TUE), which evaluates the efficiency of tool use within agent workflows. We further discuss strategies for improving explainability in Agentic AI , as well as approaches to enhancing security and privacy through encryption, adversarial robustness, and regulatory compliance. The review concludes with a research roadmap for the responsible development and deployment of Agentic AI, outlining critical directions to align emerging systems with TRiSM principles for safe, transparent, and accountable operation.
keywords:
Agentic AI, LLM-based Multi-Agent Systems, TRiSM, AI Governance, Explainability, ModelOps, Application Security, Model Privacy, AI Agents, Trustworthy AI, Risk Management, AI Safety, Privacy-Preserving AI, Adversarial Robustness, Human-in-the-Loop
1 Introduction
1.1 Background and Motivation
The global market for Artificial Intelligence (AI) agents is projected to grow from \$5.4 billion in 2024 to \$7.6 billion in 2025  [1]. By mid-2025, over 70% of enterprise AI deployments are expected to involve multi-agent or action-based systems, reflecting a dramatic shift from traditional single-agent or rule-based conversational models [2]. An AI agent, typically defined as a computational entity that perceives its environment and takes actions to achieve goals  [3], has undergone a fundamental transformation. While early agents were task-specific and deterministic, modern systems have evolved significantly ,increasing operational autonomy by over 70% on benchmark tasks [4]. These Agentic AI systems now integrate large language models (LLMs), external tools, and persistent memory to support complex planning and coordinated decision-making [5].
This rapid evolution presents new opportunities but also brings challenges in coordination, control, transparency, and risk assessment across heterogeneous agents. In particular, as the shift toward agentic AI architectures amplifies and evolves, longstanding concerns around safety, fairness, interpretability, and systemic harms resurface. While such issues have been central to machine learning (ML) systems broadly, they become more promiment in Agentic AI, where autonomous agents operate in high-stakes domains like healthcare, law, and public services, areas where their decisions can carry profound and lasting societal consequences.
From Isolated Agents to Agentic AI The autonomous software agents date back to the 1990s [6], however, the advent of LLM-driven multi-agent ecosystems [7] since 2023 has made “Agentic AI”, a distinctly new and rapidly expanding research focus [5]. Traditional agents primarily focussed on automating limited tasks, such as information retrieval [8], data summarization [9], and dialogue response [10]. Typically driven by single step logic [11] or scripted rules, these traditional agents lacked the deep reasoning, adaptability, and persistence required for complex, multi-step problem solving .
In contrast, Agentic AI consists of collaborative agents with specialized roles (e.g., planner, coder, analyst), and enabled by LLMs and tool use [12]. These systems dynamically decompose tasks, share context, and pursue high-level goals across long timelines [13]. This shift reflects not only a change in technology but also a fundamental increase in complexity and autonomy, marking the emergence of machine collectives capable of exhibiting emergent and decentralized behavior.
1.2 Trust, Risk, and Security Challenges in Agentic AI
The rise of Agentic AI introduces new challenges that extend beyond those of traditional ML systems. For example, unlike deterministic agents [14], LLM-based multi-agent systems exhibit emergent and often opaque decision-making, increasing the risk of cascading errors, biases, and unintended behaviors. These characteristics necessitate a robust approach to have trust and ensuring trustworthiness, in high-stakes domains such as healthcare, finance, and law. Below, we discuss some of these challanges:
  1. 1.
    Trust and Trustworthiness In the context of Agentic AI, trust refers to a user willingness to rely on an AI system, while trustworthiness denotes whether the system consistently behaves in a safe, fair, and predictable manner, thereby deserving that trust. This distinction is critical in high-stakes settings, where undue reliance on an untrustworthy system can lead to severe consequences, such as misdiagnoses in healthcare or biased rulings in legal applications. The stochastic nature of LLM reasoning [15] further complicates trustworthiness, as randomness in output generation, while promoting creativity and ensemble diversity, introduces inconsistency and non-repeatability. These factors hinder traceability, verification, and compliance auditing, all of which are essential for trustworthy systems.
  2. 2.
    Security Risks in Multi-Agent Systems Security risks escalate as autonomous agents gain access to tools, external APIs, and persistent memory, amplifying the potential for privacy breaches, adversarial misuse, or regulatory violations in the absence of robust human oversight [16]. For example, a multi-agent supply chain optimizer may autonomously coordinate between procurement and logistics agents but inadvertently leak sensitive data or violate compliance protocols (e.g., GDPR or HIPAA) if safeguards are inadequate. The distributed and cooperative nature of multi-agent systems introduces unique vulnerabilities that traditional security measures often fail to address.
  3. 3.
    Specific Threats: Prompt Injection, Spoofing, and Impersonation LLM-based agents are susceptible to novel security threats exploiting their language-based interfaces and cooperative behaviors. One prominent threat is prompt injection [17], where attackers design inputs with hidden or malicious instructions to manipulate agent behavior. Recent studies [18] describe “prompt infection” as a phenomenon where malicious prompts propagate from one agent to another, akin to a computer virus spreading across networks [19]. Such corrupted outputs can lead to cascading effects, including data leaks, fraudulent transactions, misinformation, or coordinated misbehavior within an agent society.
    Another critical vulnerability involves spoofing and impersonation [20]. Spoofing means faking an entity’s identity or credentials (e.g., using a false name or address) to deceive a system, while impersonation involves actively pretending to be a specific agent or user to gain trust, access, or privileges. In multi-agent systems, where agents rely on credentials or tokens for authentication during coordination, an adversary could steal an API key or mimic a trusted peer to issue unauthorized commands or access sensitive information.
  4. 4.
    Privacy Risks in Multi-Agent LLM Systems LLM-based agents often need to handle user data, proprietary business information, or other sensitive inputs to fulfill their tasks [21]. In a multi-agent context, this challenge is amplified by the fact that agents may share information with each other (e.g. via a shared memory store or message passing) to collaborate.
Traditional evaluation and safety frameworks, designed for deterministic agents AI systems, are inadequate for addressing the autonomy and control of Agentic AI. To address these gaps, frameworks like AI TRiSM (Artificial Intelligence Trust, Risk, and Security Management) [22, 23] proposes lifecycle-level controls, including explainability, secure model orchestration, and privacy management. AI TRiSM is a framework to help organizations build AI systems that people can trust. The goal is to focus not just on technical performance, but also on ethical values like fairness, transparency, and security. The framework brings together practices that support governance, strengthen reliability and robustness, and protect data and make sure that AI systems are not only effective, but also aligned with human values.
Figure 1: Taxonomy of TriSM for Agentic AI systems presented in this review. It includes five major sections: Fundamentals of Agentic AI Systems ( Agentic AI vs. Traditional AI Agents, LLM-Based Architecture), The TRISM Framework (Governance, Explainability, ModelOps, Application Security, Privacy), Threats and Risks , Trust and Explainability, and Evaluation Metrics (including Human-Centered and Composite Metrics).
1.3 Problem Statement
Agentic AI systems increasingly rely on the coordination of multiple LLM-powered agents to perform complex, long-horizon tasks. However, no unified framework exists for managing the trust, risk, and security across dynamic as well as multi-agent workflows. Recent accidents, such as, an autonomous research agent citing unverifiable medical sources or collaborative swarms generating inconsistent forecasts due to memory drift, highlight that traditional safeguards are inadequate [24]. These traditional systems are vulnerable to orchestration breakdowns, agent collusion, data leakage, and opaque decision-making pathways. To ensure safety and reliability in Agentic AI systems, a structured TRiSM perspective is therefore urgently needed.
1.4 Scope and Objectives
Objectives This study is a review article that maps the TRiSM (Trust, Risk, and Security Management) [22] framework to LLM-powered, multi-agent “Agentic AI” systems. We structure the discussion around four foundational pillars of TRiSM [25]: explainability, security, lifecycle governance, and privacy. The objective of this review is to offer researchers and practitioners a coherent and actionable reference for the design, deployment, and governance of trustworthy Agentic AI ecosystems.
For the purposes of this study, we define Agentic AI as Multi-Agent Systems (MAS) [26] powered by LLMs that exhibit autonomous planning, tool use, memory retention, and emergent reasoning capabilities, with or without human supervision. These LLM-based agents can dynamically adapt workflows based on interactions with tools, users, and other agents, distinguishing them from traditional symbolic or rule-based agents.
Scope This paper is a conceptual survey and synthesis. It does not propose a new architecture for Agentic AI, nor does it include synthetic benchmark evaluations. Instead, it integrates current perspectives on AI TRiSM to inform future research and the responsible development of Agentic AI systems. The scope of this paper is confined to LLM based Agentic AI systems and AMAS  [^3] Throughout this paper, AMAS designates multi-agent systems built on LLMs.
Necessity of this Survey Agentic AI research has largely centered on agent modeling, planning, and task coordination, while comparatively overlooking critical dimensions such as adversarial robustness, lifecycle governance, and system-level explainability. As these systems are increasingly deployed in high-stakes domains - such as healthcare, scientific research, and finance the absence of an integrated TRiSM framework exposes stakeholders to opaque decision-making and unmanaged risks. This survey addresses that gap by synthesizing TRiSM principles in the context of LLM-driven multi-agent systems and providing actionable guidance for researchers, engineers, and policymakers.
To the best of our knowledge, few existing surveys systematically explore LLM-based multi-agent systems through the lens of trust, risk, and security. Table 1 underscores this gap by comparing prior reviews with respect to their coverage of TRiSM dimensions.
1.5 Contributions
Our key contributions in this review are:
  • We propose the first TRiSM framework for Agentic AI systems. This framework is explicitly tailored to systems in which several LLM-powered agents collaborate and cooperate. The framework is organized around pillars: Explainability, ModelOps, Security, Privacy and Governance around autonomous decision to guide the development and use of safe and ethical Agentic AI systems. (see Fig. 1 and Sec. 6).
  • We offer a systematic taxonomy of risks that emerge when multiple agents interact (Sec. 5). This taxonomy discusses prompt injection, memory poisoning, collusive failure, emergent misbehavior, and related threats commonly seen in AMAS.
  • We introduce an evaluation dimensions for Agentic AI systems that covers Trustworthiness, Explainability, User-centered Performance, Coordination, and Composite Assessment (Sec. 8). To support this template, we define two new metrics: the Component Synergy Score (CSS), which measures the quality of inter-agent collaboration, and the Tool Utilization Efficacy (TUE), which measures how correctly and efficiently agents invoke external tools. Our aim is to achieve high performance in both dimensions, accuracy and usability, without sacrificing one.
  • We contribute to the management and control of Agentic AI by mapping existing TRiSM techniques such as explainability methods, prompt hygiene, decision provenance, sandboxing, and ModelOps pipelines, to AMAS environments. We critically analyze their limitations and adaptation needs, providing a foundation for re-engineering these techniques to fit the demands of interacting LLM agents (Secs. 79).
  • Building on the identified research gaps, we outline a research agenda that prioritizes scalable, verifiable, and regulation-compliant Agentic AI (Sec. 11).
2 Related Work
Table 1: Comparison of Related Surveys on LLM-Based Multi-Agent Systems and TRiSM Aspects. Existing work offers limited coverage of trust, risk, and security in Agentic AI.
[…]
Guo et al. (2024) [27] ✓ (simulated env’ts) ~ (research challenges)
Chen et al. (2025) [28] ~ (task/simulation focus) ~ (future research areas)
Yan et al. (2025) [29] ✓ (mentions diverse) ~ (future work directions)
Tran et al. (2025) [30] ✓ (networks, QA, etc.) ~ (open challenges)
Lin et al. (2025) [31] ~ (creative tasks) ~ (roadmap for research)
Fang et al. (2025) [32] ~ ~ ✓ (health, finance cited) ~ (technical “next steps”)
Xi et al. (2025) [33] ✓ (single-/multi-agent,
human-AI coop)
~ (open problems)
Luo et al. (2025) [34] ~ ~ ✓ (science, games, etc.) ~ (open challenges)
Zou et al. (2025) [35] ~ ✓ (e.g., finance,
healthcare)
~ (challenges & opp.)
Wang et al. (2025) [36] ~ ~ ~ (healthcare domain) ~ (future opportunities)
This Survey (2025) ✓ (high-stakes domains) ✓ (tailored guidance)
Legend: ✓= explicitly addressed; ~= partially/indirectly addressed; ✗= not addressed.
[…]
We present the related work in this section.
LLM-Based Multi-Agent Surveys (Technical Focus)
Prior surveys on LLM-driven AMAS predominantly focus on technical dimensions such as system architectures, agent capabilities, and domain-specific applications of Agentic AI. However, there remains a gap in addressing governance and policy aspects. For example, reviews like [27] and [28] delve into simulated environments, inter-agent communication protocols, and performance benchmarks. These works mainly detail how agents coordinate or scale but fall short in examining adversarial threats, such as prompt injection attacks or agent collusion, lifecycle governance (e.g., continuous monitoring or regulatory compliance), or system-wide explainability, especially for emergent behaviors in complex systems.
More specialized surveys, such as [29], focus on natural language communication, while [30] details collaboration mechanisms, exploring interaction protocols and practical use cases in domains like question-answering systems, 5G networks, and Industry 5.0. Although [29] highlights security risks associated with communication channels, and [30] discusses deployment contexts, both remain narrowly centered on coordination strategies. There is a need for deeper analysis of robustness against adversarial manipulation (e.g., Byzantine failures), governance frameworks for multi-agent deployments, and explainability challenges at scale.
Trustworthy/Responsible AI Surveys (Broad Trust Focus)
There are some review articles that address TRiSM in AI, such as [32], tackle ethical and risk-related issues including alignment with human values, discuss fairness in decision-making [37], and defenses against privacy attacks (e.g., data leakage). However, these works typically target general ML systems and overlook the unique dynamics of LLM-based multi-agent systems. For instance, while they may advocate for privacy-preserving techniques like differential privacy, they do not address privacy risks in inter-agent data sharing or the distributed nature of LLM-driven systems. Moreover, critical TRiSM aspects [38] like governance (e.g., auditing autonomous agents) and explainability (e.g., tracing decisions across multiple agents) are often absent in recent discussions.
Current Study (Multi-Agent Systems and LLMs)
This review distinguishes itself by mapping TRiSM principles into the study of AMAS. Unlike prior studies (see Table 1), which offer limited insight into governance and policy, this work operationalizes TRiSM principles in practice. It provides actionable guidance for researchers, engineers, and policymakers, addressing critical gaps in both the technical AMAS literature and the broader responsible AI landscape.
3 Literature Review Methodology
To review the literature related to TRiSM as applied in Agentic AI systems, we adopted a structured methodology informed by best practices in systematic reviews [39, 40]. This section outlines our research objectives, data sources, inclusion criteria, and classification strategy.
(a) Agentic AI papers per year (2022–2025).
(b) TRiSM pillars papers by topic.
Figure 2: (a) Annual trend in Agentic AI research, showing a sharp rise post-2023. (b) Distribution of TRiSM-related publications across governance and safety topics.
Research Objectives The review was guided by the following four specific questions:
  • RQ1: What are the key trust, risk, and security challenges posed by LLM-enabled Agentic AI systems? How do current approaches establish and measure the trustworthiness of Agentic AI outputs and decisions?
  • RQ2: What technical and governance strategies have been proposed to address these challenges?
  • RQ3: How do existing approaches map onto the TRiSM pillars: explainability (including trustworthiness), model operations, security, and privacy?
  • RQ4: What gaps remain in current research, and what directions are promising for future work?
Search Strategy To conduct a review on Agentic AI and TRiSM, we systematically searched major digital libraries, including IEEE Xplore, ACM Digital Library, SpringerLink, arXiv, ScienceDirect, and Google Scholar. The search encompassed publications from January 2022 to May 2025, with select seminal works prior to 2022 included to provide foundational context. We utilized Boolean search queries with an expanded set of keywords to capture the interdisciplinary intersection of LLM-based multi-agent systems and TRiSM considerations. The search terms were structured as follows:
(“ Agentic AI” OR “multi-agent systems” OR “multi-agent LLMs” OR “AI agents” OR “autonomous agents” OR “intelligent agents” OR “collaborative AI” OR “distributed AI” OR “LLM-based agents” OR “agent coordination” OR “agent-based modeling” OR “swarm intelligence”)
AND (“trust” OR “trustworthiness” OR “risk” OR “security” OR “safety” OR “governance” OR “oversight” OR “compliance” OR “explainability” OR “interpretability” OR “transparency” OR “privacy” OR “data protection” OR “robustness” OR “accountability” OR “ethical AI” OR “fairness” OR “adversarial attacks” OR “regulatory compliance” OR “bias mitigation” OR “system reliability”)
Inclusion and Exclusion Criteria We applied the following criteria during the screening process:
Inclusion Criteria:
  • Studies published between 2022 and 2025, to capture developments following the introduction of LLMs (mostly post-ChatGPT).
  • Studies that explicitly discuss Agentic AI, Multi-Agents, or LLM-powered AI systems in the context of at least one TRiSM dimension: trust, risk, security, governance, explainability, or privacy.
  • Peer-reviewed articles, preprints, or reports from credible sources (e.g., arXiv, NeurIPS, IEEE, ACM, Nature, or governmental standards bodies).
Exclusion Criteria:
  • Papers focused solely on traditional, rule-based, or symbolic agents without any integration of LLMs or emergent coordination.
  • Studies addressing only low-level ML components (e.g., training optimization, architecture design) without reference to agentic behavior or TRiSM concerns.
  • Non-English papers or those lacking sufficient metadata (e.g., missing abstracts or publication details).
After screening titles and abstracts, we shortlisted 250 papers. A subsequent full-text review narrowed the selection to 180 primary studies that directly addressed one or more TRiSM pillars. In addition to peer-reviewed literature, we also incorporated relevant technical reports, blogs, and website links to contextualize and map the emerging landscape of Agentic AI within the TRiSM framework. An analysis on the literature is shown in Figure 2.
Quality Assurance To ensure the reliability of included studies, we conducted a quality assessment adapted from established systematic review guidelines [41]. Each paper was evaluated against the following criteria: (1) Does the study clearly define its objectives in the context of agentic or LLM-based systems? (2) Are the methods, experiments, or architectural designs well-documented and reproducible? (3) Does the paper substantively engage with at least one of the trust, risk, security, explainability, or privacy aspects? (4) Does the study offer empirical, theoretical, or normative contributions relevant to Agentic AI governance? Each criterion was rated as low, medium, or high. Studies scoring “low” in more than two dimensions were excluded or flagged for contextual relevance only. Quality ratings were assigned independently by four reviewers, with discrepancies resolved through discussion.
4 Fundamentals of Agentic AI Systems
In this section, we present the fundamentals of Agentic AI systems.
Table 2: Traditional AI Agents vs. Agentic AI Systems
Dimension Traditional AI Agents Agentic AI Systems
Autonomy Model × Reactive or deliberative (e.g., cognitive architectures) Goal-driven planning & adaptation
$\bullet$ Fixed action sequences or symbolic reasoning $\bullet$ Recursive self-improvement
Cognitive Foundation × Symbolic logic, FSMs, or cognitive models Foundation models (LLMs/LIMs)
$\bullet$ Hand-coded or structured KBs $\bullet$ Emergent reasoning
Intelligence Scope × Narrow & task-specific Broad & compositional
$\bullet$ Single-domain focus $\bullet$ Cross-domain transfer
Reasoning Approach × Deterministic or multi-step symbolic Multi-step & contextual
$\bullet$ Rule-based or production rules $\bullet$ Chain-of-Thought (CoT)
Collaboration × Isolated or limited coordination Role-specialized coordination
$\bullet$ Manual decomposition $\bullet$ Automated hierarchy
Temporal Context × Episodic, often stateless Persistent memory
$\bullet$ Session-bound or episodic memory $\bullet$ VectorDB/LTM
Orchestration × Hard-coded or fixed workflows Dynamic meta-agents
$\bullet$ Sequential or predefined goals $\bullet$ Conflict resolution
Tool Utilization × Static or domain-specific tools Planned API invocation
$\bullet$ Handcrafted interfaces $\bullet$ Function calling
Context Awareness × Bounded or heuristic context Memory-augmented
$\bullet$ Limited or cognitive context $\bullet$ RAG architectures
Learning Mechanism × Limited or rule-based learning Emergent learning
$\bullet$ Fixed rules or chunking $\bullet$ Fine-tuning/RL
Human-AI Interaction × Scripted or cognitive modeling Natural language interfaces
$\bullet$ Limited or human-like reasoning $\bullet$ Conversational adaptability
Exemplars $\triangleright$ MYCIN  $\triangleright$ ELIZA  $\triangleright$ SOAR  $\triangleright$ ACT-R $\triangleright$ AutoGen  $\triangleright$ ChatDev  $\triangleright$ MetaGPT  $\triangleright$ AgentVerse
= Core strength  × = Fundamental limitation  $\bullet$ = Implementation characteristic  $\triangleright$ = Modern framework  $\triangleright$ = Classical system
4.1 Traditional AI Agents vs Agentic AI
Traditional agents operate through predefined rules [42], heuristic workflows [43, 44], or deterministic logic [45], and perform well in narrow, well-defined environments. Cognitive Agents [46], a subset of traditional agents, model human-like reasoning by integrating perception, memory, and decision-making, often using architectures like SOAR [47] or ACT-R [48]. However, they remain constrained by static knowledge bases and limited adaptability [46].
In contrast to traditional AI agents, Agentic AI systems leverage foundation models (mainly LLMs) to achieve adaptive goal-oriented behavior through multi-agent coordination and emergent reasoning capabilities. Agentic AI systems fundamentally redefine autonomy through three core innovations: (1) Multi-agent coordination: Specialized agents, such as planners and verifiers, collaborate using structured protocols [13, 49]. (2) Persistent context: Memory architectures maintain task state across workflows, ensuring continuity [50]. (3) Dynamic meta-orchestration: Systems delegate tasks and resolve conflicts dynamically [51].
This Agentic AI architecture supports longitudinal task execution and cross-domain generalization, overcoming the context fragmentation inherent in traditional approaches [52]. For example, systems like AutoGen [53] enable LLM-backed agents to decompose complex tasks, such as software development, through multi-step reasoning [54]. This capability of Agentic AI represents a significant shift from reactive to proactive problem-solving systems [49]. A comparison of traditional AI agents and Agentic AI systems is presented in Table 2.
4.2 LLM-Based Agentic AI System Architecture
Agentic Multi-Agent System (AMAS) represent an emerging paradigm in AI where multiple LLM-powered agents operate semi-autonomously, interact with external tools, and collaborate to achieve complex tasks. As illustrated in Figure 3, a typical AMAS architecture comprises several key components that together form a flexible yet highly dynamic ecosystem. At the core are multiple LLM-Based Agents, each capable of reasoning, planning, and tool invocation [55]. These agents access a Shared Toolchain Interface [56] to execute code, perform searches, or interact with domain-specific APIs.
Figure 3: Architecture of LLM-based Agentic Multi-Agent System (AMAS), illustrating key functional layers: Perception/Input Layer (text, image, audio processing), Cognitive/Reasoning Layer (goal-setting, planning, decision-making), Action/Execution Layer (digital and physical task execution), Learning Module (supervised and reinforcement learning), Communication and Collaboration (agent messaging and coordination), Data Storage & Retrieval (centralized/distributed databases), and Monitoring & Governance (ethical oversight, observability, and compliance mechanisms). The modular design highlights adaptive intelligence and inter-agent synergy.
Communication and coordination are facilitated through a Communication Middleware [57] that allows agents to share goals, observations, or intermediate results. A Task Manager or Orchestrator [58] governs high-level planning and delegates subtasks to agents based on their roles or specializations. Agents can read from and write to a World Model or Shared Memory [27], which stores contextual knowledge, system state, or evolving task data. Human oversight is supported through a Human-in-the-Loop Interface [44], enabling users to prompt, correct, or halt agent behavior.
To ensure accountability, a Trust and Audit module monitors agent actions, logs tool usage, and generates behavioral traces [59]. However, this modular and distributed structure introduces significant TRiSM [25] challenges. With multiple autonomous agents accessing external resources, the Security Gateway becomes critical for enforcing access controls, authentication, and sandboxing [60].
Likewise, a dedicated Privacy Management Layer is essential to prevent leakage of sensitive or personally identifiable information [61], especially when data traverses multiple agents or tools. Finally, an Explainability Interface must provide interpretable rationales for multi-agent decisions, supporting transparency and trust calibration [62]. Together, these architectural elements make AMAS powerful yet complex, raising unique and urgent questions about how to ensure their trustworthiness, mitigate systemic risks, and secure them against adversarial behaviors. Below, we discuss the architecture of AMAS.
Language Model Core (Agent Brain). At the center of an Agentic AI system lies a LLM serving as the primary decision-making controller or “brain” [63]. The core LLM is initialized with a user goal and a structured agent prompt (defining its role, capabilities, and tool access). It then generates step-by-step decisions or actions, interpreting instructions, producing reasoning traces, and selecting next steps in either natural language or structured action formats. In many agent frameworks, such as AutoGPT [64], Baby AGI [65] and GPT Engineer [66], the LLM governs the full control loop, orchestrating the overall system behavior.
Planning and Reasoning Module. To handle complex goals, an explicit planning mechanism decomposes tasks into manageable sub-goals. This can be done internally via chain-of-thought (CoT) or tree-of-thoughts prompting (ToT) [67], where the model performs intermediate reasoning before arriving at a final decision. Reasoning is also one aspect of cognitive functionalities [68]. Some implementations employ external planning systems by translating goals into structured planning languages and using classical planners for long-term decision-making. Planning is often interleaved with execution and feedback: the agent refines its plan based on outcomes, alternating between reasoning, acting, and integrating observations. Techniques like Reasoning and Acting (ReAct) [69] exemplify this loop by guiding the agent through repeated reasoning–action–observation cycles, improving performance on complex tasks.
Memory Module. Agentic AI systems integrate memory to maintain context across iterations. This includes short-term memory (recent interactions held within the prompt context) and long-term memory (accumulated knowledge or experiences) [7]. Long-term memory is often implemented using vector databases, where key facts or past events are stored and retrieved by similarity search. By reintegrating past data into the LLM prompt, the agent can recall relevant information across sessions, avoid repetition, and support coherent long-term planning. Effective memory management enables adaptive, learning-driven behavior.
Table 3: Comparison of representative LLM-based Agentic AI frameworks across key design axes.
Framework (Year) Core LLM Planning Memory Tool Use Notable Features
AutoGPT [70] GPT-4 Self-looped CoT Vector DB OS shell + web Fully autonomous goal loop
BabyAGI [65] GPT-3.5/4 Task queue In-memory Web search Minimal task generator
GPT Engineer [71] GPT-4V Spec-to-code File cache Python REPL End-to-end code generation pipeline
LangGraph [72] Model-agnostic FSM via graph Persistent nodes Custom modules Visual orchestration of agent graphs
AutoGen [53] GPT-4 Multi-agent PDDL JSON/DB API calls Modular, reusable agent templates
MRKL [73] LLaMA/GPT Prompt router N/A Math + search tools Neuro-symbolic expert routing
Reflexion [74] GPT-3.5/4 Retry–reflect loop Episodic buffer Same as base agent Verbal self-improvement via reflection
MetaGPT [75] GPT-4 SOP workflow YAML state Git CLI Structured roles for software engineering
Voyager [76] GPT-4 Auto skill tree Task DB Minecraft API Lifelong open-ended learning in environments
WebVoyager [77] LLaVA-1.6 ReAct JSON store Browser actions Multimodal web interaction via vision and text
HuggingGPT [78] ChatGPT Task-plan-select Log store Hugging Face models External model orchestration by LLMs
CAMEL [79] GPT-4 Role-play CoT Dialogue log Chat only Multi-agent role simulation via dialogue
ChatDev [80] GPT-4 Chat chain File repo Unix tools Simulated software development workflow
CrewAI [81] Any Declarative plan Optional DB Python modules Lightweight, LangChain-free framework
AgentVerse [82] Model-agnostic Config graph Redis/KV store Plugin API Multi-agent simulation and task solving
OpenAgents [83] Model-agnostic Agent scripts MongoDB Web plugins Open platform with public hosting
SuperAGI [84] GPT-4 DAG workflow Postgres Pinecone Concurrent agents for production use
Tool-Use Interface. To extend its capabilities beyond text generation, the agent is equipped with a tool-use interface [85]. This layer allows the invocation of external tools such as web search, APIs, code interpreters, or databases. The available tools are specified in the agent prompt with command schemas. When the LLM determines a tool is needed, it emits a structured command, which is executed externally. The result is fed back into the LLM as a new observation. This mechanism enables the agent to access real-time information, perform computations, and interact with external systems dynamically. Tool-use frameworks like Modular Reasoning, Knowledge, and Language (MRKL) [73] illustrate this design by routing queries to different expert modules (symbolic or learned) with an LLM as the router. Similarly, approaches like Toolformer [85] train the model to insert API calls in its text generation.
Perception and Environment Interface. For agents interacting with dynamic environments, such as web interfaces, simulated worlds, or physical systems, an observation-action interface is essential [86]. Perception modules translate raw inputs (e.g., sensor data, images, or textual states) into representations the LLM can process. Conversely, the agent’s chosen actions are executed within the environment, and the resulting state changes are returned to the agent as observations. This loop supports sense-plan-act cycles that continue until the task is completed or halted. In robotic or multimodal settings, the interface may include additional sensory models (e.g., vision transformers), but the core flow remains consistent.
Integration and Autonomy. These modules together form a closed-loop architecture. The LLM plans and reasons over tasks, guided by memory and tools, and interfaces with the environment to execute decisions and observe results [12]. Each iteration updates the agent’s context, enabling self-refinement through dynamic feedback. This integrated design empowers Agentic AI systems to operate autonomously, pursue long-range goals, and exhibit adaptive behavior across dynamic environments.
Table  3 presents representative systems and shows how each one maps these five design axes (LLM core, planning/ reasoning, memory, tool-use, environment interface) into practice. For example, AutoGPT pairs a GPT-4 core with a self-looped chain-of-thought planner and a vector-database memory to serve as both tool layer and environment interface.
5 Threats and Risks in LLM-based Agentic Multi-Agent Systems (AMAS)
In this section, we present the main threats in AMAS and also discuss a taxonomy of emergent risk classes.
Figure 4: Overview of Key Risks and Threats in Agentic AI Systems. The left panel outlines system-level threats unique to agentic architectures, including Autonomy Abuse, Persistent Memory, and Agent Orchestration. The right panel presents a taxonomy of emergent risk classes : Adversarial Attacks, Data Leakage, Agent Collusion, and Emergent Behaviors that arises from multi-agent interactions, memory persistence, and decentralized decision-making.
5.1 Threats in Agentic AI
Agentic AI systems introduce a distinct set of security and reliability concerns compared to traditional single-agent LLM architectures [87]. Below, we discuss some of these major threats.
  • Autonomy abuse. The foremost threat is the autonomy abuse, where agents with significant decision-making authority might misinterpret objectives or implement harmful plans due to erroneous reasoning or manipulated inputs . Unlike deterministic models, these Agentic AI systems dynamically generate actions, complicating efforts to define and enforce safe operational states .
  • Persistent memory. Another threat is the persistent memory, which, while crucial for context retention, introduces unique vulnerabilities through potential prompt injections. Such contamination can propagate subtly via shared memory, especially in the absence of detailed version control and robust audit mechanisms .
  • Agent orchestration. This risk involves central or distributed control mechanisms for role assignment and workflow mediation. A compromised orchestrator could distort task distribution or misroute information, triggering cascading failures, issues exemplified by documented vulnerabilities in MetaGPT and AutoGen . These orchestration vulnerabilities differentiate Agentic AI systems from conventional stateless, single-threaded LLM deployments.
  • Goal Misalignment and Exposure. Agentic AI systems often operate with internalized or user-specified goals. Improper scoping or exposure of goal representations can lead to prompt extraction, misalignment, or competitive agent behaviors, particularly when third-party tools or APIs are invoked dynamically during task completion.
  • Tool Misuse and External API Exploits. Agents capable of invoking external tools (e.g., via APIs, code execution, or web access) present novel attack surfaces. Misuse of these tools, intentional or accidental, can trigger financial costs, Distributed Denial of Service (DDoS) behavior, or legal violations. Sandboxing and invocation tracing remain limited in most open-source frameworks.
  • Multi-Agent Collusion or Drift. Emergent coordination among agents, while useful for collaboration, can also lead to collusion, competitive drift, or polarization of outputs, especially in self-reinforcing memory loops or agent swarms. Such behavior complicates interpretability and may undermine safety alignment goals.
5.2 Taxonomy of Risks
We propose a taxonomy of risks associated with AMAS, categorizing them into four primary classes: adversarial attacks, data leakage, agent collusion, and emergent behaviors, as shown in Figure 4. These categories encapsulate the diverse challenges arising from the complex interactions, persistent memory, and autonomous decision-making inherent in AMAS. Below, we delineate each risk category, highlighting their mechanisms, implications, and illustrative examples.
  • Adversarial Attacks. Autonomous agents are susceptible to a range of adversarial manipulations, including prompt injections, gradient-based attacks, and engineered reasoning traps. These vulnerabilities are exacerbated in AMAS due to the cascading effects of agent interactions, where a single compromised agent can propagate malicious influences across the system [88]. For instance, the role-swapping attack observed in ChatDev demonstrates how adversaries can exploit inter-agent dependencies to subvert system objectives, leading to unauthorized access or manipulated outputs [13]. Such attacks underscore the need for robust defenses, including adversarial training and real-time monitoring, to mitigate risks in dynamic, multi-agent environments.
  • Data Leakage. The integration of persistent memory and extensive inter-agent communication in AMAS heightens the risk of unintended disclosure of sensitive information. In domains such as financial services and human resources, where confidentiality is paramount, inadequate boundary enforcement and ineffective data sanitization protocols can lead to significant breaches [89]. For example, an agent processing proprietary financial data may inadvertently share sensitive information with other agents or external systems due to insufficient access controls. Addressing data leakage requires rigorous data isolation mechanisms, encryption, and continuous auditing to ensure compliance with privacy regulations.
  • Agent Collusion and Mode Collapse. Coordination mechanisms designed to enhance collaboration among agents can inadvertently foster undesirable outcomes, such as groupthink or mode collapse, where agents reinforce mutual errors or converge on suboptimal solutions [90]. This phenomenon, often termed agent collusion, arises when agents prioritize consensus over critical evaluation, leading to echo chambers that amplify biases or inaccuracies. For instance, in multi-agent systems tasked with decision-making, repeated reinforcement of flawed assumptions can degrade system performance. Mitigating these risks necessitates diversity in agent design, robust feedback loops, and mechanisms to detect and correct convergent errors.
  • Emergent Behaviors. The complex interplay of agents, memory components, tools, and tasks in AMAS can give rise to unpredictable behaviors that elude conventional testing and validation frameworks [91]. Such emergent behaviors may manifest as unintended shortcuts, where agents optimize for efficiency at the expense of critical verification processes, or as the suppression of contradictory information, compromising system reliability. Notable examples include blockchain-based systems, where agents bypassed consensus protocols to expedite transactions [92], and audio verification systems, where agents overlooked subtle discrepancies in authentication signals . These cases highlight the need for adaptive monitoring to anticipate and address emergent risks.
5.3 Real-World Examples
Several real-world examples shows the consequences risks in deployed or experimental Agentic AI systems.
Prompt Leakage in Agentic AI Systems. Instances of prompt leakage have been observed in LLM-based agent frameworks such as AutoGPT [93], where recursive prompt augmentation and insufficient memory controls led to the unintentional exposure of sensitive information. In one reported scenario, sensitive tokens were stored in persistent memory and later surfaced in planning summaries or external logs [94]. Such vulnerabilities underscore the critical importance of implementing memory sanitization, access controls, and prompt boundary protections to safeguard Agentic AI systems from cascading information leaks.
Collusive Failure in ChatDev. In a collaborative code generation session involving planner, coder, and tester agents within the ChatDev framework, an error in a shared planning module led to the propagation of a faulty design [80]. Due to the absence of external ground-truth or objective feedback loops, all agents validated each other’s outputs, resulting in a feedback loop of erroneous confirmations. This scenario highlights the necessity of incorporating diverse information sources and adversarial checks within agent loops to prevent such collusive failures [18].
Simulation Attack in Swarm Robotics. In a simulated swarm robotics experiment utilizing LLM-based planning strategies, an agent was provided with a misleading environmental assumption, leading to a coordination failure characterized by spatial congestion and task incompletion [95]. This incident demonstrates the potential vulnerabilities in real-world deployments, particularly in critical infrastructure or logistics, where such failures could have significant consequences. The case highlights the importance of robust validation mechanisms and the integration of diverse information sources to ensure reliable swarm behavior [96].
Memory Poisoning in Multi-Agent Chatbots. In a multi-agent customer support system, a customer-facing agent injected sarcastic feedback into a persistent feedback buffer. This buffer was later utilized by the policy improvement agent to adapt dialogue strategies, resulting in responses with inappropriate tones. This incident highlights the importance of implementing validation filters, sentiment monitoring, and robust feedback loop governance in self-adapting systems to prevent such memory poisoning vulnerabilities [97].
System Prompt Drift in Autonomous Memory Agents. In experiments with agents using system-level memory (e.g., LangGraph or BabyAGI), over time, system prompts began drifting due to self-appended contextual memory that wasn’t properly versioned or validated. This led to hallucinated goals and emergent behaviors misaligned with initial intentions [98].
Next, we discuss the AI TRiSM prism framework and map its core principles to a AMAS.
6 The TRiSM Framework for Agentic AI
6.1 Overview of AI TRiSM and Its Core Components
The TRiSM Framework, short for Trust, Risk, and Security Management, offers a structured lens for evaluating and governing AI systems, especially those that exhibit autonomous, agentic behaviors. In the context of Agentic AI, where LLMs are deployed as self-directed agents capable of perception, reasoning, and interaction, the TRiSM framework provides a critical foundation for ensuring robustness, transparency, and responsible deployment.
Table 4: TRiSM pillars in Agentic AI systems with aligned governance references.
Pillar Primary Objective Key Practices & Agentic-LLM Considerations Exemplar Standards / Metrics
Explainability Transparent multi-agent decisions SHAP, LIME, counterfactuals; reasoning-trace logs; cross-agent dependency mapping NIST AI RMF “Explain” [99]
ISO/IEC 24029-1 [100] , ISO/IEC 42005:2025 [101]
Metrics: fidelity, sufficiency, human-trust score
ModelOps Lifecycle governance of agents, models, orchestration Version-controlled prompts; CI/CD w/ multi-agent sims; drift monitoring; rollback policies ISO/IEC 24029-1 [100] , ISO/IEC 42005:2025 [101]
SRE MTTR/MTTD; $\Delta$-accuracy (model drift); pipeline pass-rate
Application Security Prevent prompt injection, spoofing, lateral exploits Prompt sanitation; least-privilege scopes; encrypted comms; sandboxed tool calls; anomaly detectors OWASP Top-10 for LLM Apps [102]
Jailbreak-success%; mean-time-to-detect; exploit CVSS
Model Privacy Safeguard sensitive data and shared memory Differential-privacy training; data minimization; homomorphic encryption; secure enclaves; audit logs GDPR Art. 25 [103]
HIPAA §164 [104]
Metrics: $\epsilon$-DP budget, leakage rate, access-audit pass-rate
  • AI Trust Management Establishing trust in AI systems necessitates a foundational commitment to transparency, accountability, and fairness. Trust management focuses on embedding mechanisms that enable AI models to generate interpretable and justifiable outputs, thereby facilitating meaningful human oversight [50]. This involves the adoption of explainability tools, the implementation of governance protocols, and alignment with ethical standards to address systemic biases and ensure equitable outcomes across diverse user groups.
  • AI Risk Management Effective risk management within AI systems entails the systematic identification, analysis, and mitigation of potential harms arising from the deployment and operation of these technologies. This includes technical risks such as algorithmic bias [37], model errors, and data leakage, as well as societal risks involving fairness, safety, and misuse [32]. Structured risk assessment frameworks enable stakeholders to anticipate failure modes and design mitigation strategies that enhance system robustness and reduce the likelihood of adverse impacts.
  • AI Security Management Security management in the context of AI is concerned with safeguarding models, data, and infrastructure from adversarial threats and unauthorized access. This encompasses both conventional cybersecurity practices, such as encryption, authentication, and access control, and AI-specific considerations like adversarial example detection, model extraction prevention, and secure model deployment pipelines [19]. Continuous monitoring and threat modeling are critical for identifying vulnerabilities and ensuring the resilience and integrity of AI systems over time
Originally highlighted in industry guidelines for AI governance [105], TRiSM provides a structured approach to managing the unique challenges of LLM-based “Agentic AI” systems. The TRiSM framework addresses trust concerns in Agentic AI by focusing on four key pillars: Explainability, Model Operations (ModelOps), Application Security, and Model Privacy [59, 23]. Figure 5 illustrates how these four TRiSM pillars map onto the core architectural layers of an LLM-based Agentic AI system.
Figure 5: TRiSM-Aligned Architecture for LLM-Driven Agentic AI Systems. The system is organised into four vertical layers: Trust & Explainability, ModelOps, Planning + Reasoning, and Model Privacy, each supported by a cross-cutting Monitoring & Governance Plane that enforces policies, logging, and compliance triggers. External tools are accessed through a single Secure API gateway, while the user interface feeds back into both the Trust layer and the governance plane to enable end-to-end oversight.
6.2 Mapping Pillars of TRiSM with Agentic AI
Alignment of TRiSM Pillars with Governance Frameworks Explainability: Aligned with the EU AI Act (Art. 13–14) [106] and the NIST AI RMF “Explain” function [107]. This pillar promotes transparent, rationale, human-interpretable logging, and role-based interpretability [108]. ModelOps : Anchored in the NIST RMF “Govern” function [107] and ISO/IEC 24029-1 [100] , ISO/IEC 42005:2025 [101]. It emphasizes lifecycle controls, model versioning, rollback policies, and CI/CD pipelines with safety gates. Application Security: Guided by OWASP LLM Top-10 [102] and OECD AI Principles on Robustness [109], it covers access controls, prompt sanitization, sandboxing, and audit trail mechanisms. Model Privacy: Aligned with GDPR (Art. 25) [103], HIPAA §164 [104], and OECD AI Principles [109], this pillar ensures data minimization, encryption, differential privacy budgeting, and access-audit pass thresholds. Governance: Grounded in overarching regulatory frameworks such as GDPR, CCPA [110], and the OECD AI Principles [109], this pillar supports institutional oversight, policy compliance, and cross-system accountability. It enforces auditable workflows, verifiable role-based access, regulatory traceability, and human oversight, particularly in multi-agent and high-risk AI deployments.
Figure 6: Governance alignment of TRiSM pillars with international frameworks such as the EU AI Act, NIST RMF, ISO/IEC 42001, 42005, and OECD Principles. Each pillar operationalizes specific compliance strategies to ensure transparency, resilience, and trust in LLM-based systems.
Explainability in Multi-Agent Decision Making
Explainability refers to making the inner workings and decisions of AI agents interpretable to humans. In the context of LLM-based multi-agent systems, it is paramount for building user trust because outcomes often emerge from complex interactions among multiple agents rather than a single model’s prediction [111]. However, achieving explainability in such systems is particularly challenging , one must interpret not only individual model decisions, but also the inter-agent dynamics that lead to the final outcomes [112]. In other words, explaining why a multi-agent system behaved a certain way requires tracing which agent did what and how their collaboration produced an emergent result, a task that traditional single-model explainability techniques do not directly address. Several approaches are being explored to enhance explainability in LLM-driven multi-agent systems:
  • Feature Attribution (LIME/ SHAP): Established explainable AI techniques provide a starting point. For instance, LIME) [113] and Shapley Additive Explanations (SHAP) [114] can be adapted to analyze LLM decisions. These techniques identify which features or input factors most influenced an agent’s output, offering insight into why a particular action or response was taken. In an Agentic LLM setting, a “feature” might be a component of the agent’s input context or a signal from another agent. By highlighting the most influential parts of each agent’s input, methods like LIME/SHAP help clarify individual agent decisions in the multi-agent mix [115].
  • Counterfactual Analysis Beyond local feature attributions, “what-if” analyses are increasingly important for multi-agent explainability. Counterfactual techniques examine how the system’s behavior would change if certain conditions or agent behaviors were altered [116]. For example, one can remove or modify a particular agent’s contribution and then observe how the collective outcome changes. This approach, rooted in causal inference, helps isolate each agent’s role in collaborative decision-making.
  • Reasoning Trace Logging: Instead of treating the agents as black boxes, an explicit trace of their intermediate reasoning can greatly enhance transparency. Logging the step-by-step reasoning process – for example, an agent’s chain-of-thought (CoT) prompts or the dialogue exchanged between agents – creates an audit trail of how the decision emerged [69]. By examining this trace, one can see how agents influence each other, which agent contributed what, and how partial conclusions build up to the final decision. This visibility into the multi-agent process addresses a key blind spot of standard XAI, effectively “opening up” the black box of inter-agent interactions.
  • Natural Language Explanations: Finally, the generative nature of LLMs can be leveraged to turn the above insights and logs into human-friendly explanations [117]. Instead of expecting end-users to read raw CoT logs or analyze feature weights, an “explainer” agent (which could be another LLM) or a human analyst can compile the reasoning trace into a narrative. In fact, an important aspect of explainability is communicating the reasons behind decisions in an intelligible way. By using natural language, the system can answer user questions like “Why did the agents decide to do X?” in a conversational manner.
ModelOps : Lifecycle Management for LLM Agents
ModelOps (Machine Learning Operations) is the discipline of managing AI models through their entire lifecycle, from development and deployment to monitoring, maintenance, and eventual retirement [118]. Within AMAS, ModelOps covers not just individual models, but the orchestration of multiple agents and the supporting infrastructure that keeps them running safely [59]. Effective ModelOps helps maintain consistency, performance, and regulatory compliance as LLM agents evolve or are added to a system. AMAS require rigorous lifecycle governance because their behavior can change with model updates, prompt adjustments, or environmental variations. A cornerstone of ModelOps in this context is version control, i.e., tracking and managing versions of each agent’s model and prompt configurations. Additionally, robust CI/CD pipelines (Continuous Integration/Continuous Deployment) are employed to automatically test agents’ performance and safety whenever a model is fine-tuned or an agent’s logic is modified. Before deployment, multi-agent simulations and unit tests validate that new agent behaviors do not introduce regressions or unsafe interactions.
Application Security for LLM Agents
To mitigate threats such as prompt injection, spoofing, and impersonation, the TRiSM security pillar recommends a defense-in-depth approach tailored to LLM agents. This includes prompt hygiene [19] , where agents treat inputs defensively by sanitizing and filtering prompts, and applying guardrails or content policies to detect and refuse suspicious instructions. Techniques like prompt hardening, such as secure prefixes or validation steps, can help reduce susceptibility to injection attacks [119]. It also involves strong authentication and access controls, ensuring that each agent and human user is securely authenticated, and that least-privilege principles are enforced to limit autonomous agent actions [120]. Finally, continuous monitoring plays a critical role: if an agent begins issuing anomalous requests or deviates from typical behavior patterns, automated systems can flag the incident for review or trigger a shutdown to prevent further risk.
Recent frameworks, such as LangChain [121], AutoGen [122], and CrewAI [81] introduce the idea of trust scores or reputation among agents, where agents verify each other’s outputs and cross-check decisions to catch inconsistencies or signs of compromise. Moreover, training LLM agents with adversarial robustness in mind (e.g. fine-tuning on adversarial examples, employing adversarial training regimes) [19] can improve their resistance to malicious inputs.
Model Privacy and Data Protection
The model privacy addresses the protection of sensitive data within AI agent systems, ensuring that the use of personal or confidential information complies with privacy regulations and ethical norms [21]. Since multiple agents may share information with each other (e.g. via a shared memory store or message passing) to collaborate. Without strict privacy controls, there is a risk that an agent could inadvertently expose private data to unauthorized parties or that sensitive information could “leak” through the language model’s outputs. To preserve privacy in AMAS, several technical measures can be employed. For example, Differential Privacy (DP) adds calibrated noise during model training to ensure that no individual data point can significantly influence the output, thus preventing memorization of personal information [123]. Data anonymization and minimization further enhance privacy by limiting shared information to only what is essential, often in aggregated or pseudonymized formats, such as using “age bracket 30–40” instead of precise birthdates [124]. Secure Multi-Party Computation (SMPC) allows agents to jointly compute functions without revealing their private inputs, making it particularly useful in cross-organizational scenarios like collaborative fraud detection [125]. Similarly, Homomorphic Encryption (HE) permits computation directly on encrypted data; with Fully Homomorphic Encryption (FHE), even plaintext queries and responses remain inaccessible to the agent itself [126].Trusted Execution Environments (TEEs), or secure enclaves, offer hardware-based isolation, ensuring that even privileged users cannot access data during processing, which is an essential safeguard for secure memory sharing and execution [127].
Governance
Governance of AI systems necessitates compliance with evolving legal and policy frameworks, including the General Data Protection Regulation (GDPR) [103] and the California Consumer Privacy Act (CCPA) [110]. Effective governance extends beyond regulatory adherence to cover proactive management of legal, ethical, and societal risks. This includes establishing transparent accountability structures, maintaining auditable documentation, and ensuring the protection of data subject rights throughout the AI lifecycle.
Several global frameworks provide guiding principles that align with the TRiSM pillars. The EU AI Act [106], particularly Article 14, mandates transparency, traceability, and human oversight in high-risk systems; principles embedded within TRiSM’s Explainability and ModelOps dimensions. Similarly, the NIST AI Risk Management Framework (AI RMF) [107] identifies: Explainability, Governability, and Validity as key components of AI risk mitigation, which directly correspond to TRiSM’s lifecycle control mechanisms. The OECD AI Principles [109] reinforce the importance of robustness, transparency, and accountability, echoing TRiSM’s attention to secure tool invocation, memory traceability, and multi-agent auditability.
Together, these frameworks underscore the need for governance architectures in Agentic AI systems that support real-time monitoring, role-based access, verifiable logs, and intelligible outputs. As illustrated in Figure 6, TRiSM offers a practical scaffolding for operationalizing these requirements in LLM-based multi-agent environments.
7 Trust and Explainability in Agentic AI Systems
Table 5: Trust-Enabling Mechanisms in Agentic AI Systems
Mechanism Purpose Example System(s) Trust Contribution
Transparency via Reasoning Traceability Make agent decisions observable through intermediate steps or explanations SciAgent [128], MetaGPT [129] Enhances interpretability and debuggability
Status Reporting and Intent Disclosure Provide real-time updates or clarification of progress and errors ChatDev [13], AutoGen [122] Mimics human team transparency; reduces uncertainty
Human-in-the-Loop (HITL) Enable user oversight over high-stakes or irreversible actions ChemCrow [130], enterprise dashboards Prevents unsafe execution; reinforces control
Behavioral Consistency & Bounded Autonomy Ensure agents operate within predefined roles and limits Adaptive tutors [131, 132] Improves predictability; prevents overreach
Social Trust Cues Use polite language, uncertainty and expressions GPT-4 safety-tuned profiles [133, 134] Enhances user comfort and trust in uncertain scenarios
Two key factors that influence trust in Agentic AI are the transparency of the agent’s decision-making processes and the ability to explain or justify its actions in human-understandable terms. Figure 7 illustrates the layered relationship between trust mechanisms, explainability techniques, and evaluation facets. Trust-enabling mechanisms, such as transparency and human-in-the-loop (HIIL) oversight, provide the foundation for user confidence. These feed into explainability techniques like decision provenance graphs, which make agent behaviors interpretable. Finally, evaluation (Section 8) facets ensure that trust and explainability are quantifiable, enabling iterative improvements in system design.
7.1 Building Trust in Autonomous Agents
In Agentic AI systems, building trust is foundational to user acceptance, system reliability, and responsible deployment, especially as these systems begin to make autonomous decisions in critical domains such as healthcare, finance, and scientific research [135]. Unlike traditional software agents, autonomous LLM-based agents are characterized by self-directed reasoning, adaptive memory, and dynamic collaboration, which make their operations opaque and often unpredictable [136]. Establishing trust in such systems, therefore, requires a combination of technical transparency, user feedback integration, and robust oversight mechanisms.
Trust-Enabling Mechanisms Transparency • Status Reporting • HITL • Bounded Autonomy • Social Cues Explainability Techniques LIME • SHAP • Decision Provenance • Attention Maps • Prompt Audits Evaluation Facets Trustworthiness • Explainability • User-Centric • Coordination • Composite
Figure 7: Relationship between trust mechanisms, explainability techniques, and evaluation facets in Agentic AI.
Transparency and Decision Traceability. Transparency is one of the core enablers of trust. For users to understand and evaluate AMAS decisions; the reasoning chains, decision states, and action triggers of agents must be made observable. Several Agentic AI systems are now integrating decision traceability through mechanisms such as CoT prompting and self-explanation modules. For example, SciAgent [128] generates scientific summaries and provides justifications by linking outputs to source documents retrieved via retrieval-augmented generation (RAG). Similarly, MetaGPT[129] structures its reasoning using role-based outputs, where each agent (e.g., planner, coder) explicitly states the logic behind its task execution, creating modular interpretability.
Status Reporting and Progress Visibility. Clear reporting of intent and intermediate status is also essential. Human collaborators often require updates about what an agent is doing, why a task is taking longer, or how an agent interprets ambiguous instructions. Tools like AutoGen[122] and ChatDev[13] have incorporated structured chat interfaces where agents summarize their intermediate progress, decisions, and encountered errors.
Human-in-the-Loop Oversight. Human oversight and intervention mechanisms further reinforce trust. Allowing human users to review, edit, or approve agent-generated outputs not only prevents missteps but also signals that the system respects user authority [137, 138, 139]. Many systems adopt a human-in-the-loop (HITL) paradigm [140], where agents request confirmation before executing high-risk or irreversible actions.
Behavioral Consistency and Bounded Autonomy. Trust requires predictability in models’ behaviour. Agents should follow defined roles, output in expected formats, and remain within delegated authority. For example, in enterprise AI platforms used for automated data analytics, agents may generate insights or dashboards but defer publishing until a domain expert reviews the material. Similarly, in autonomous research assistants like ChemCrow [130], agents pause to allow chemists to validate proposed reactions or data pipelines before proceeding, reinforcing safe deployment.
Social Trust Cues and Language Behaviors. Beyond system-level mechanisms, behavioral consistency and guided autonomy are crucial. If an agent behaves unpredictably or inconsistently, even if technically correct, users are less likely to trust it. Behavioral alignment mechanisms such as predefined role protocols can help output style consistency [141, 12, 142]. For example, in adaptive educational platforms using AI tutors, agents may be allowed to revise lesson plans but not change grading criteria, preserving institutional trust boundaries . Also, social trust cues, such as polite language, turn-taking, and cooperative gestures, have shown promise in reinforcing user trust even in non-expert settings [143]. Studies have found that users trust agents more when they express uncertainty ("I’m not sure, but here’s what I found") rather than overconfidence [134]. This has been implemented in models like GPT-4 when configured with safety-tuned instruction sets, improving reliability perception without undermining capability. These mechanisms form a layered trust strategy for Agentic AI, as summarized in Table 5.
7.2 Explainability Techniques in AMAS
Emergent behavior of AMAS poses a challenge: the interpretability of an individual agent does not necessarily imply the interpretability of the overall system. In platforms such as ChatDev [13] or AutoGen [53], agents emulate specialized roles (e.g., engineer, reviewer), and tracing a final action back to its source agent is often non-trivial. While local explainability techniques such as LIME [113] and SHAP [114] remain foundational, their application in Agentic AI is limited due to decentralized agent contexts and shared memory. Researchers have thus proposed composite frameworks that combine local attribution with global decision traceability, such as decision provenance graphs [144], causal influence chains [145], and multi-agent SHAP adaptations [146]. These techniques aim to surface the system-level dynamics that underlie emergent agent behaviors.
Symbolic and Hybrid Architectures. Another promising direction is the use of inherently interpretable modules such as rule-based planners and decision trees within hybrid architectures [5]. These agents offer built-in explainability but retain the generative capabilities of LLMs for broader context understanding. Such designs are increasingly used in domains where structure and interpretability are prioritized, such as educational AI or mission planning.
Lightweight Interpretability: Attention and Prompts. Attention map visualizations have been used to highlight focus areas in multi-modal language agents [147, 148, 149], offering lightweight but informative views into model behavior. Prompt audit trails logging prompt history, agent actions, and response metadata have also gained traction [150, 151]. These mechanisms support system debugging, safety evaluations, and human-in-the-loop fine-tuning in multi-agent environments.
Table 6: Explainability Techniques in Agentic AI Systems
Technique Purpose Challenges in Agentic AI Example Use Cases
LIME [113] Local post-hoc explanation using surrogate models Fails to capture cross-agent dependencies in multi-agent setups Feature-level auditing in fraud detection
SHAP [114] Attribution of prediction to input features via Shapley Difficult to aggregate across agents with divergent contexts Financial risk analysis, regulatory explanations
Decision Provenance Graphs [144, 152] Trace inter-agent decision paths and data flows Can be complex to interpret at scale Collaborative document generation, planning chains
Multi-agent SHAP [146] Attribution extended to agents and shared memory High complexity, lack of standard frameworks Strategic gameplay, collaborative writing
Symbolic/Rule-Based Agents [5] Built-in interpretability via logic or rule-based models Limited flexibility compared to LLMs Hybrid planning systems, educational tutors
Attention Maps [147, 149] Visualize focus of language/image models Only reflects internal focus, not logic or intent Multimodal VQA, image captioning
Prompt Audit Trails [150, 151] Log input prompts, outputs, and agent actions Requires infrastructure, may miss implicit state changes Debugging multi-agent sequences, fine-tuning
The integration of trust-enabling mechanisms and explainability techniques is critical for the adoption of Agentic AI systems in high-stakes domains. While transparency, human oversight, and interpretable architectures improve user confidence, challenges in capturing system-level dynamics and scaling explainability remain. Future research should focus on developing standardized frameworks for multi-agent explainability, real-time interactive querying, and longitudinal trust evaluation to ensure robust and responsible deployment. A comparative summary of current techniques is shown in Table 6.
8 Evaluation of Agentic AI Systems
Agentic AI systems demand a comprehensive evaluation across multiple dimensions beyond traditional accuracy. We outline five key categories of metrics: trustworthiness, explainability, user-centered performance, coordination, and composite scores, each capturing a distinct aspect of an AMAS performance and its real-world implications. Below, we discuss these metrics and summarize in Table 7.
Trustworthiness. This criteria evaluates the reliability, safety, and ethical alignment of the AI agent. A trustworthy agent consistently produces correct and unbiased results, adheres to safety constraints, and avoids harmful or unpredictable behavior. Metrics for trustworthiness include success rates on tasks under varying conditions (measuring robustness), violation rates of safety or ethical guidelines (which should be minimal), and calibration of the agent’s confidence (how well the agent’s self-reported confidence aligns with actual accuracy) [153]. Some approaches combine such factors into an overall trust score. We can define a trustworthiness score $T$ for an Agentic AI system as :
$T=\frac{C+R+I}{S}\,,$ (1)
where $C$ is the agent’s credibility (accuracy and correctness of its outputs), $R$ is reliability (consistency of performance over time), $I$ is the level of user alignment or rapport (analogous to “intimacy” in trust modeling), and $S$ is self-orientation (the degree to which the agent pursues its own goals over the user’s goals). A higher $T$ indicates an agent that is accurate, consistent, user-aligned, and not self-serving, which corresponds to greater trustworthiness. In practice, achieving high trustworthiness means the agent behaves predictably and transparently in accordance with ethical AI principles (such as fairness and accountability).
Explainability. Explainability metrics assess how well the agent’s decisions can be understood and traced by humans [154]. These metrics focus on the clarity and completeness of the rationale the agent provides for its actions. For instance, one can measure the explanation coverage (the percentage of decisions or outputs that come with an adequate explanation) and the fidelity of explanations (how accurately the explanation reflects the true reasoning or model logic).
Consistency of explanations for similar scenarios is another important metric: the agent should explain comparable decisions in a similar way, indicating a stable reasoning process. Quantitatively, methods like OpenXAI [155] provide a suite of metrics to evaluate explanation quality across dimensions such as faithfulness, stability, and fairness of explanations. High explainability builds user trust, as users can follow why the agent made a decision, and it aids debugging by revealing the agent’s internal decision process. In regulated domains (e.g., healthcare or finance), explainability is often essential for compliance and user acceptance.
User-Centered Performance. User-centered metrics capture how effectively the AI agent interacts with and satisfies the end-user’s needs. These criteria emphasize the user’s experience and outcomes [156]. Key metrics include user satisfaction ratings, typically collected via surveys or feedback after interaction, which reflect whether the user’s goals were met and how comfortable they felt with the Agentic AI system behavior.
Additionally, interaction metrics like the number of back-and-forth clarification queries needed (fewer indicates the agent understood the user well) and the coherence or naturalness of the conversational flow contribute to user-centered evaluation. Human-in-the-loop evaluations are often employed here: for example, user studies might rate the agent on criteria such as helpfulness, clarity and naturalness of language, and adherence to user instructions. Ultimately, a user-centered Agentic AI system should align its actions with user intent and preferences.
Table 7: Summary of key evaluation metrics for Agentic AI systems. Each category addresses a different facet of an autonomous agent’s performance and behavior.
Metric Aspect Evaluation Focus (Examples of Metrics)
Trustworthiness Reliability and safety of the agent’s behavior. Example metrics: task success rate across diverse scenarios; frequency of rule or safety violations (lower is better); calibration of confidence vs. outcomes; fairness and bias indices.
Explainability Transparency and interpretability of the agent’s decisions. Example metrics: percentage of outputs with an accompanying explanation; explanation fidelity to actual model reasoning; consistency of explanations for similar cases; human interpretability ratings of explanations.
User-Centered User satisfaction and alignment with user needs. Example metrics: user satisfaction score (post-interaction survey); rate of fulfilling user-defined goals; number of clarification questions needed (lower is better); naturalness and coherence of dialog from the user’s perspective.
Coordination Effective collaboration in multi-agent or modular systems. Example metrics: multi-agent task completion rate; communication overhead (e.g., messages exchanged); consistency of shared plans or beliefs among agents; synergy score quantifying complementary actions.
Composite Overall performance across metrics. Example metrics: weighted aggregate score encompassing trust, explainability, user-centric, and coordination measures; specialized composite indices (e.g., TUE for tool-use efficacy); cross-domain benchmark results (aggregate success across diverse tasks, as in AgentBench).
Coordination. In scenarios where an Agentic AI system consists of multiple cooperating agents or modular components, coordination metrics measure how effectively these parts work together. Good coordination means agents share information, divide labor without conflict or redundancy, and converge on solutions efficiently. Quantitative measures include the team success rate on collaborative tasks (whether the group of agents achieves the overall goal) and communication efficiency metrics (e.g., the number of messages or iterations required among agents to reach a decision, with fewer often indicating more efficient interaction).
One specific example is the Component Synergy Score (CSS), which counts or weights effective interactions between agents, reflecting how well each agent’s actions complement the others (a higher CSS means agents are synergistic rather than working on individual basis). MAS such as ChatDev and MetaGPT provide practical testbeds for these metrics: they orchestrate specialized agents (e.g., different roles in a software engineering pipeline) that must cooperate to complete complex projects.
Evaluations on such frameworks examine whether agents maintain a consistent shared plan, handle inter-agent dependencies smoothly, and recover from misunderstandings. For instance, if one agent generates a plan and another executes it, a coordination metric would assess if the executing agent follows the planner’s intent correctly and whether both agents remain in agreement throughout the process. High coordination scores indicate that the agentic system functions as a cohesive unit, which is crucial for complex tasks beyond the capability of any single agent.
Composite Metrics. Composite metrics aggregate multiple evaluation aspects into a single overall score. These are useful for summarizing agents’ performance holistically, especially when comparing different systems. A composite metric is often a weighted combination of the above categories, for example:
$M_{\text{composite}}=w_{T}M_{T}+w_{E}M_{E}+w_{U}M_{U}+w_{C}M_{C}\,,$ (2)
where $M_{T},M_{E},M_{U},M_{C}$ are the normalized scores (on a common scale) for trustworthiness, explainability, user-centered performance, and coordination, respectively, and $w_{T},w_{E},w_{U},w_{C}$ are weights reflecting the relative importance of each aspect for a given application. The choice of weights $w_{i}$ can be domain-specific (for instance, in healthcare applications, trustworthiness and explainability might be weighted more heavily than raw efficiency).
An example of a specialized composite metric in agentic contexts is the Tool Utilization Efficacy (TUE) score, which combines sub-metrics evaluating how correctly and efficiently an agent uses external tools (including proper tool selection and correct parameter usage in tool calls) into one measure. By condensing multiple criteria, composite metrics enable high-level comparison and benchmarking of Agentic AI systems. For instance, AgentBench [157] is a comprehensive benchmark that evaluates agents across a diverse range of tasks and environments (from operating system manipulation to web shopping), effectively providing a composite performance profile of an agent. Such aggregated scores highlight if an agent performs strongly across the board or if it excels in some dimensions while underperforming in others. It is important to interpret composite scores in light of their components: a single number can mask specific weaknesses (e.g., an agent might achieve a high overall score by doing well in task completion and coordination, yet still have poor explainability). Therefore, composite metrics are most informative when accompanied by a breakdown of the agent’s performance per category.
9 Security and Privacy in Agentic AI Systems
In this section, we discuss the security and privacy in AMAS.
Table 8: Governance Dimensions for Agentic AI Systems
Governance Component Description Example Tools or Methods
Regulatory Compliance Aligning with legal and industry standards for responsible AI [158] NIST AI RMF [99], EU AI Act [159], GDPR [103], HIPAA [104]
Auditability and Logging Capturing traceable records of agent actions and decisions [160] Decision provenance, immutable logs, blockchain-based traceability
Policy Enforcement Enforcing system-level operational, ethical, and security constraints [52, 161] RBAC/ABAC policies [162], formal logic rules, dynamic sandboxing [60, 163, 164]
Human Oversight Human-in-the-loop [140] decision checkpoints and override capabilities Human review agents, interactive dashboards, compliance checkpoints
Risk Monitoring Continuous identification and mitigation of systemic or emergent risks [165] Model risk scanners, out-of-distribution detectors, reinforcement learning audit modules
Explainability Governance Ensuring interpretability standards are met in decision pipelines [166] LIME/SHAP [167, 112, 168], decision trees [169, 170], interpretable surrogate models [171, 172, 173]
Adaptive Governance Updating governance mechanisms based on model evolution and deployment context [133, 135] Governance-as-code [174, 175], auto-updated rule engines [176, 177, 178], learning-based policy adaptation [90, 179]
Incident Response and Recovery Procedures for handling, logging, and recovering from failures or breaches Real-time alerting, kill-switch mechanisms, secure rollback protocols
9.1 Security Mechanisms
Agentic AI systems, composed of loosely coupled yet collaboratively functioning LLM-based agents, introduce an expanded attack surface relative to conventional AI agents [180]. Ensuring the security of such systems necessitates a multi-layered defense architecture that addresses data protection, execution integrity, inter-agent communication, and model robustness [181]. Among the foundational techniques employed are encryption , access control , adversarial defense , and runtime monitoring, each adapted to the unique demands of decentralized multi-agent environments.
Encryption. Encryption plays an important role in safeguarding data exchanged between multiple agents, especially when sensitive or regulated content (e.g., healthcare records, financial data) is involved [61]. Agentic workflows often include inter-agent handoff of partially processed results, models, or prompts. Implementations such as SSL/TLS, homomorphic encryption , and secure enclaves (e.g., Intel SGX) are increasingly integrated into Agentic AI pipelines to ensure confidentiality across message-passing protocols.
Access control. Access conrol becomes crucial when orchestrators or shared memory modules manage permissions for agents with distinct capabilities and responsibilities. For instance, in systems like AutoGen and CrewAI where agents take on specialized roles (e.g., summarizer, planner, coder); enforcing principle-of-least-privilege access prevents privilege escalation and unauthorized tool invocation [53]. Agent-based access control policies, often aligned with Role-Based Access Control (RBAC) [182] and Attribute-Based Access Control (ABAC) [182] paradigms, can dynamically restrict which agents may access sensitive APIs, files, or memory buffers, based on contextual trust levels.
Adversarial learning. Adversarial learning is a growing concern as LLM-based agents are susceptible to prompt injection, manipulation through poisoned tool outputs, or coordination disruption via malformed intermediate results. Recent studies have shown that multi-agent LLM frameworks can be destabilized by adversarially designing outputs from one compromised agent propagating misleading information to others [183]. Adversarial training methods, such as input perturbation, reward shaping, and contrastive learning, can partially mitigate these vulnerabilities. Integrating safety constraints and verifying tool responses before execution are also effective mitigation strategies.
Runtime Monitoring. Runtime monitoring systems support the detection of anomalous agent behaviors, especially in high-stakes domains like automated healthcare or cybersecurity . Log-based auditing, anomaly detection with LSTM or autoencoder-based detectors, and trust scoring among agents are becoming essential components of real-time surveillance layers [184]. For example, Microsoft’s Copilot governance layers monitor anomalous agent behavior across sessions to ensure compliant execution and flag potentially harmful interactions.
9.2 Privacy-Preserving Techniques
The decentralized and interactive nature of AMAS introduces new challenges for preserving privacy, especially as agents continuously communicate, access external data sources, and store episodic or shared memory. To ensure data confidentiality and protect personally identifiable information (PII), Agentic AI systems must adopt robust privacy-preserving techniques such as differential privacy, data minimization, and secure computation.
Differential privacy (DP). DP offers mathematically grounded guarantees by injecting statistical noise into outputs, ensuring that individual user contributions cannot be re-identified. In multi-agent LLM systems, DP can be applied during training or at inference-time when agents exchange information. For example, Google’s implementation of DP in federated learning frameworks can be extended to distributed agentic systems, where agents collaboratively train or fine-tune local models without exposing raw data [185]. DP-SGD and privacy budgets ($\epsilon$-differentials) can regulate information exposure during policy updates or collaborative planning in real-time decision-making agents.
Data minimization. Agentic AI systems can mitigate exposure risks by limiting the scope, granularity, and duration of data collected or retained during task execution. For instance, temporary memory buffers used in systems like ChatDev or ReAct-based pipelines are cleared once subgoals are completed, preventing persistent storage of unnecessary user data. Furthermore, anonymization and pseudony [186] techniques can help remove identifying features before data is passed between agents or stored in shared memory repositories.
Secure computation. Techniques including secure multi-party computation (SMPC), homomorphic encryption, and trusted execution environments (TEEs) enable agents to perform computations over encrypted or obfuscated data without compromising privacy. In scenarios where agents collaborate across different organizational boundaries (e.g., federated medical agents or cross-silo industrial agents), SMPC allows joint computations such as diagnostics or anomaly detection without data leakage. Homomorphic encryption, while computationally expensive, is increasingly being explored to allow arithmetic operations on encrypted vectors used in RAG workflows.
Privacy-by-design [187] principles are becoming central to the engineering of next-generation Agentic AI systems. Architectures now embed user consent layers, configurable privacy settings, and memory redaction modules that allow end-users or system administrators to control what agents can remember or share. As Agentic AI expands into domains such as personalized education, healthcare, and finance, ensuring privacy-respecting behaviors will be essential for regulatory compliance (e.g., GDPR, HIPAA) and public trust.
10 Compliance to AI Governance and Policy in Agentic AI Systems
As Agentic AI systems grow more autonomous and capable, ensuring they remain compliant with regulations and governed by robust policies becomes imperative. Unlike traditional AI, these multi-agent or highly autonomous systems operate with persistent memory and complex decision-making loops, which introduces new risks around safety, ethics, and legality. Recent industry surveys indicate that enterprise leaders rank security, trust, compliance, and oversight as top concerns when deploying agentic AI at scale. To address these concerns, governance of agentic AI spans three critical dimensions: adherence to regulatory standards, system-level auditability, and enforceable policy frameworks. Below, we expand on each dimension and highlight emerging best practices for secure and compliant agentic AI deployments.
10.1 Regulatory standards
Regulatory frameworks provide baseline requirements that all AI systems – including agent-based architectures – must meet. Key frameworks outline principles of transparency, accountability, safety, and fairness which are especially pertinent for autonomous multi-agent systems operating in high-stakes domains (e.g. healthcare, finance, transportation). Notable examples include:
  • NIST AI Risk Management Framework (AI RMF) [188], a voluntary U.S. framework emphasizing risk identification, mitigation, and trustworthy AI practices across the AI lifecycle. It calls for governance functions that ensure AI systems are transparent, accountable, and monitored for risks. Adopting the NIST AI RMF “Govern” and “Explain” functions helps organizations embed policies and documentation that align with trustworthiness goals.
  • EU AI Act [106], a comprehensive EU regulation (set to fully apply in 2026) that classifies AI systems by risk level. High-risk AI systems (a category that would likely include certain autonomous or agentic AI in critical fields) are subject to strict requirements: rigorous risk management, human oversight, transparency, and continuous monitoring. For example, the Act explicitly mandates ongoing risk assessment and traceability for high-risk AI, and it requires that such systems be designed with human-in-the-loop or human supervision capabilities (per Article 14 on oversight). Agentic AI developers must therefore build in features for traceability and human override to remain compliant with these upcoming rules.
  • Domain-Specific Regulations In addition to AI-focused standards, Agentic AI systems must comply with sectoral data and safety laws. For instance, an autonomous multi-agent healthcare assistant must adhere to health privacy laws like HIPAA [104], and any system handling personal data of EU citizens must follow GDPR [103] requirements for data protection by design. These regulations often require data minimization and user consent, which Agentic AI systems need to respect (e.g. by expiring or anonymizing stored user data to avoid retention of sensitive information).
10.2 Auditability
Auditability is critical for transparency and post-hoc accountability in agentic AI. Given the complexity of multi-agent decision-making (with potentially emergent or opaque interactions), organizations must be able to reconstruct and scrutinize the chain of events leading to any outcomes. Achieving robust auditability involves implementing the following practices:
  • Comprehensive Logging: Each decision, action, plan, and inter-agent message should be logged with rich metadata – including timestamps, the initiating agent’s identity/role, the context or prompt, and a rationale or justification for the action. Such detailed logs create a decision provenance trail, enabling auditors to trace how and why a particular outcome was reached.
  • Action Traceability: Techniques like decision provenance link each agent’s actions to preceding events or inputs, so one can follow the sequence of steps across agents. In complex agent workflows (e.g. a chain of a Planner agent delegating to a Coder agent, then to a Tester agent), traceability means every hand-off and decision node is recorded and can be reviewed in context. This helps identify where an error or policy violation might have occurred.
  • Role-Specific Audit Trails: Modern agentic frameworks often assign specialized roles to different agents (for example, in AutoGen or MetaGPT, agents may take on roles like Researcher, Coder, or Reviewer to collaborate on tasks). The audit system should capture actions at the granularity of each role. By tagging log entries with the agent role, one can detect anomalies or biases in specific functions – for instance, flagging if a “Reviewer” agent consistently approves outputs that the “Researcher” agent flagged as doubtful, or if a certain role’s actions lead to security issues
  • Immutable Logging: To ensure the integrity of audit logs (so that records cannot be tampered with after the fact), organizations are exploring append-only, tamper-proof logging mechanisms. For example, using blockchain or other immutable ledger technology to record multi-agent interactions can provide verifiable audit trails. This means any decision or message recorded in the system is cryptographically secured against alteration, which is valuable for compliance checks and forensic analysis in case of an incident.
Robust auditability not only helps with internal accountability but is also increasingly expected by regulators. Transparent record-keeping allows external auditors or regulators to verify compliance with regulations and ethical standards. In the event of a problematic outcome, detailed audit logs ensure that responsibility can be determined and corrective action taken. In fact, governance frameworks stress that organizations should implement such audit trails and explicable decision logs so that any harmful or biased result can be traced back and attributed appropriately. Overall, auditability in agentic AI builds the foundation for trust – both for internal stakeholders and for users or regulators – by making the system’s behavior observable and explainable after the fact.
10.3 Policy Enforcement
While regulations and audits provide oversight and after-the-fact analysis, policy enforcement mechanisms work in real-time to keep an Agentic AI system’s behavior within allowed bounds. Policy in this context refers to the codified rules and constraints that the AI agents must obey – covering operational limits, ethical guidelines, and security restrictions set by developers or governance boards. In practice, effective policy enforcement in agentic AI includes:
  • Orchestrator-Level Governance: Agentic AI systems typically have an orchestration layer or a meta-controller that coordinates multiple agents. This layer should be programmed to enforce global policies on what agents can do and under what conditions. For example, a policy might dictate that any plan an agent comes up with must receive approval from a supervisory agent (or a human) if it involves deleting or altering important data. The orchestration logic would then require that checkpoint before execution, overriding agent autonomy when necessary.
  • Memory and Data Retention Policies: To comply with privacy rules and avoid unchecked growth of an agent’s knowledge base, policies can enforce memory expiration or sanitization. Agents might be required to forget or redact certain information after a task is complete or after a time limit, preventing long-term retention of sensitive data. This guards against data misuse and helps comply with regulations like GDPR’s “data minimization” principle. For instance, if an agent system stores user queries or personal details, a policy could mandate purging those from the agent’s memory after fulfilling the query (unless explicitly allowed to retain).
  • Tool and Resource Access Control: Agentic AI often involves agents calling external tools, APIs, or databases to accomplish tasks. Role-Based Access Control (RBAC) and Attribute-Based Access Control (ABAC) are crucial for limiting what each agent (or agent role) is permitted to do. Under RBAC, each agent role has permissions aligned with its responsibilities – e.g. a “Coder” agent might be allowed to call a code compilation service but not to fetch customer data; a “Data Analyst” agent might read a database but not invoke external network calls. ABAC can add context-sensitive rules (attributes like time of day, the specific project, or the agent’s trust level) to further restrict actions.
  • Dynamic Sandboxing and Execution Constraints: Each agent’s actions (especially those that can change the environment or have side effects) should run in a sandbox or constrained context where possible. For example, if an agent is meant to write to a file, sandboxing can ensure it only writes to a designated directory and not to system-critical files. Similarly, agents might be prohibited from executing certain operations (like sending network requests to unapproved domains) through whitelists/blacklists coded into the environment. Dynamic sandboxing means these execution rules can adjust based on the agent’s role or current task, providing flexible yet controlled operation.
  • Real-time Monitoring and Intervention: Even with policies encoded, agents might encounter novel situations. Real-time monitoring systems serve as a safety net – they watch agent behavior and can halt or flag any activity that deviates from predefined ethical constraints or operational bounds. For instance, if an agent’s output is about to violate a known policy (say, it’s about to generate disallowed content or leak PII), a monitoring module can catch this and either censor the output or stop the agent’s process. Some approaches involve using formal verification and runtime checking tools (like TLA+ specifications or symbolic execution) to prove certain unsafe actions will not occur, or to continuously verify that invariants hold during execution.
10.4 Emerging Governance Practices
Because the field of Agentic AI is evolving rapidly, governance strategies are also adapting. Beyond the foundational controls above, organizations are adopting new best practices to strengthen oversight and ensure long-term compliance as these AI systems scale. Some of the emerging practices include:
  • AI Governance Boards: Many organizations are establishing dedicated governance boards or committees for AI oversight. These boards typically include stakeholders from legal, ethics, security, and domain experts. Their role is to continuously review the Agentic AI system’s objectives, assess risks, and update policies or strategies as needed. For example, a governance board might conduct regular audits of an AI’s decisions for bias or fairness, ensure alignment with the latest regulations, and approve any major changes to the system’s capabilities. This human layer of oversight creates accountability at the organizational level and sends a signal that the AI’s operations are being closely supervised by responsible parties.
  • Governance-as-Code: Inspired by the infrastructure-as-code paradigm in DevOps, governance-as-code means encoding governance policies and ethical rules directly into the software and configuration that drive the AI system. Instead of governance being only a manual or paperwork process, it becomes part of the system’s logic. For Agentic AI, this could involve using policy engines or rule-based systems that automatically check each agent action against a set of coded rules. For instance, a rule repository might specify “Agent of type X cannot use tool Y without permission Z” and this is enforced by the orchestration code.
  • Adaptive Governance Layers: Agentic AI systems may change over time – they might learn new skills, add more agents, or move into new deployment contexts. Adaptive governance refers to frameworks that evolve along with the AI system. This can include auto-updating rule sets or ML-driven policy adjustments. For example, if an agent starts operating in a new country, an adaptive governance layer might automatically incorporate that country’s AI regulations into the policy set. Or if a system update gives agents new capabilities, the governance rules adapt (or prompt human governors to adapt them) to cover those capabilities.
  • Standards Alignment and Certification: To have trust, organizations are aligning their Agentic AI governance with emerging international standards. A prime example is ISO/IEC 42001:2023 [189], and ISO/IEC 42005:2025 [101] the management system standard, which specify requirements for effective AI governance and risk management . ISO 42001 and 42005 emphasizes many of the same principles discussed above – such as transparency, accountability, bias mitigation, safety, and privacy in AI development and deployment. Forward-looking organizations are therefore proactively building those capabilities (e.g. implementing rigorous risk management pipelines, bias audits, etc.) so that by the time new laws fully kick in, their agentic AI will be in natural compliance.
By integrating these enforcement measures, Agentic AI systems can operate with much tighter alignment to intended guidelines and ethical norms. Policy enforcement translates high-level governance principles into day-to-day operational safeguards, reducing the chance of an autonomous agent straying outside its allowed domain or causing a compliance breach. Below, we summarize the core governance components required for secure and compliant Agentic AI deployment in Table 8.
11 Discussion
In this section, we present our findings and outline directions for future work.
11.1 Technical Implications of TRiSM for Agentic AI Design
Rather than treating LLM agents as black-box decision-makers, TRiSM encourages instrumenting them with continuous oversight “guardrails” [190]. For example, there is discussion on designing specialized “guardian agents” within AMAS [38]. In this paradigm, such agents serve as proactive monitors that filter sensitive data and establish baselines of normal behavior, while operator agents dynamically enforce policies at runtime (e.g. blocking disallowed actions such as outputting personally identifiable information).
Prior research highlights risks like “excessive agency” [120] where an LLM given too much autonomy or tool access can produce unintended harmful actions (for instance, via hallucination or misinterpreted goals). TRiSM-driven Agentic AI design mitigates these risks by limiting agent autonomy within well-defined safety bounds. Likewise, emerging threats specific to Agentic AI systems such as prompt injection attacks, memory poisoning , or cascading hallucinations, highlights the need for more built in risk controls. By incorporating anomaly detection and policy-checking modules, an LLM agent can detect deviations from normal behavior and either alert humans or automatically neutralize the threat (e.g. masking a sensitive datum or stopping an unsafe action).
11.2 Ethical and Societal Ramifications of Multi-Agent AI
Beyond technical matters, deploying networks of autonomous LLM agents raises pressing ethical and societal questions. Applying TRiSM in this context emphasizes principles of accountability, human oversight, and fairness [158], all of which are vital for public trust in AI systems. Governance in TRiSM recommends that organizations retain clear responsibility for their AI actions, and do not obscure blame behind algorithmic “black boxes”. This implies implementing audit trails and explicable decision logs so that any harmful or biased outcome can be traced and attributed . Recent guidance on trustworthy AI [161] often highlights accountability and explainability as key pillars of trust.
In practice, this review sheds some light on how each autonomous agent decisions should be transparent enough to be understood and challenged by human reviewers when necessary [191]. Human oversight is another ethical imperative tightly coupled with accountability. TRiSM does not seek to eliminate humans from the loop; rather, it provides a structured way for humans and AI agents to collaborate under a defined governance framework [192].
In AMAS setups, humans can monitor agent swarms in real time, pause or shut down agents exhibiting anomalies, and adjust policies on the fly . Therisk of “user complacency”, trusting an autonomous agent too much, has been noted as a hazard [193]. TRiSM governance counteracts this by formalizing oversight roles and ensuring no AI operates without appropriate human or regulatory supervision.
11.3 Alignment with Emerging AI Regulations and Standards
The principles embedded in TRiSM align closely with emerging regulatory frameworks for AI. This convergence means that adopting TRiSM-based governance can help organizations meet new legal obligations and industry standards. For example, the European Union’s AI Act [106] (set to fully apply in 2026) mandates rigorous risk management, transparency, data governance, and human oversight for “high-risk” AI systems. These are precisely the capabilities that a TRiSM approach promotes.
Deploying Agentic AI systems responsibly requires a multifaceted governance approach. By adhering to established regulatory standards (and keeping abreast of new ones), by ensuring auditability for transparency and accountability, and by enforcing policies through technical and organizational means, we can significantly mitigate the risks posed by autonomous AI agents. Moreover, embracing evolving best practices – from governance boards to adaptive, code-driven policies – will help maintain control over these systems as they grow in complexity. The end goal is to harness the benefits of Agentic AI (increased efficiency, creativity, and problem-solving at scale) while staying within legal, ethical, and safe bounds, thereby preserving public trust and meeting compliance obligations.
11.4 Limitations and Current Research Gaps
While the TRiSM-based approach appears promising, we also discuss some limitations and open challenges in current research. First, the classical AI domains that works on standard test suites, there is no consensus on how to measure an AI Agent vs the Agentic AI ability to operate safely under TRiSM principles. This makes it difficult to compare different governance strategies or to track progress objectively. We encourage future work to develop evaluation frameworks, possibly extending from adversarial attack simulations [194] or “red-teaming” exercises [195] that can stress-test Agentic AI systems and score their resilience.
There is also a lack of real-world validation for many TRiSM-inspired controls. Much of the existing literature and tooling for LLM agent safety has been demonstrated in laboratory settings or on narrowly scoped tasks [25]. It remains uncertain how these governance mechanisms perform in complex, open-ended real-world environments. Another critical gap is adversarial learning. As we improve defenses, attackers will inevitably adapt [196]. Recent findings show that LLM-based systems remain vulnerable to attacks (for example, hidden prompt injections or subtle data poisoning) that can bypass superficial guardrails [19].
Beyond technical aspects, the organizational and human factors present limitations: implementing TRiSM requires interdisciplinary expertise (AI specialists, security experts, ethicists, legal advisors) and clear governance structures. Many organizations lack the necessary skill sets or frameworks, making TRiSM adoption superficial or inconsistent. Without a strong organizational commitment, even the best technical framework can falter.
11.5 Future Roadmap for Agentic AI TRiSM
Agentic AI Roadmap Scalable Agent Architectures Secure Multi-Agent Collaboration Explainable Multi-Agent Decisions Human-Centered Trust Design Autonomous Lifecycle Management Ethics & Governance Benchmarking & Evaluation Cognitive Capability Expansion
Figure 8: Agentic AI Roadmap: Mindmap Representation
Drawing from our findings and best practices from multiple disciplines, we propose several actionable directions for future research and implementation, as shown in Figure 8. These recommendations span both technical system design improvements and governance-level policy initiatives:
The community should create open benchmarks and challenge environments to test AMAS governance. For instance, a suite of scenario-based tasks (with built-in threats and ethical dilemmas) could be used to evaluate how well a TRiSM-governed Agentic AI systemthat performs relative to one without such controls. This will enable direct comparisons and drive progress on measurable metrics of trust (e.g. frequency of prevented failures or fairness outcomes).
Future system design must anticipate a continually evolving threat landscape. Techniques from cybersecurity (e.g. adversarial training, AI model “penetration testing” [197], and formal verification) should be integrated into the LLM agent development pipeline. Cross-disciplinary collaboration with security experts can yield LLM-specific hardening methods, such as dynamic prompt anomaly detectors or robust tool APIs that constrain agent actions. Additionally, creating red-team/blue-team exercises for AMAS , akin to cyber wargames [198] , can help discover vulnerabilities in a controlled way before real adversaries do.
We encourage designing better interfaces and protocols for human oversight of Agentic AI systems. Borrowing from human-computer interaction [199] and cognitive engineering [200], researchers could devise dashboards that visualize an agent society’s state, flag important decisions, and allow intuitive human intervention (pausing agents, rolling back actions, etc.).
Policymakers and industry should collaborate to create regulatory sandboxes for multi-agent AI trials. These would be controlled environments where innovators can deploy Agentic AI systems under supervision, demonstrating TRiSM controls to regulators. Insights from such pilots can inform refinements in both technical standards and regulations.
There is much to learn from other high-stakes domains. For example, the safety engineering field (e.g. aerospace, automotive) has mature practices for redundant controls and failure mode analysis; these could inspire analogous practices in AI agent design. Likewise, ethics boards in biomedical research provide a template for AI ethics committees that review agent behaviors and approve high-risk deployments. We advocate establishing multidisciplinary governance boards that include ethicists, legal experts, domain specialists, and community representatives to oversee significant deployments of autonomous AI.
12 Conclusion
The rapid advancement of Agentic AI, driven by LLM-based multi-agent systems, marks a transformative shift in autonomous systems, enabling sophisticated collaboration, planning, and decision-making across high-stakes domains such as healthcare, finance, and public services. However, this evolution introduces complex challenges in trust, risk, and security management (TRiSM) that demand a tailored framework to ensure safe, transparent, and accountable operations. This review has explored these challenges, proposing a TRiSM framework structured around four pillars: Explainability, ModelOps, Security, Privacy and Governance, adapted for LLM-based AMAS. Key contributions include a novel risk taxonomy capturing unique threats like prompt injection, memory poisoning, and collusive failures, alongside two new metrics: the Component Synergy Score (CSS) and Tool Utilization Efficacy (TUE), which provide practical tools for assessing inter-agent collaboration and tool efficiency. The review also highlights state-of-the-art techniques, such as LIME, SHAP, and decision provenance, to enhance explainability, while advocating for robust security measures like prompt hygiene, sandboxing, and homomorphic encryption to mitigate vulnerabilities. Future work should prioritize advancing adversarial robustness, refining governance protocols, and developing standardized benchmarks for trustworthiness and coordination. By aligning Agentic AI development with TRiSM principles, researchers and practitioners can foster systems that are not only innovative but also responsible, ensuring their safe and ethical integration into society.
Declaration of Competing Interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Appendix
Table A.1: Key terminologies for LLM-based agentic AI systems.
Term Definition
Agentic AI System A multi-agent architecture powered by large language models (LLMs), where autonomous agents collaborate, plan, and execute tasks over extended horizons with persistent memory and dynamic role assignment.
Autonomy Model The mechanism by which an agent decides and acts without direct human intervention, often using goal-driven planning and chain-of-thought reasoning.
Chain of Thought (CoT) A prompting strategy in which an LLM generates intermediate reasoning steps before producing a final answer or action, enhancing interpretability and multi-step planning.
Counterfactual Analysis An interpretability technique that examines how altering certain inputs or agent contributions would change the overall system outcome, revealing causal dependencies among agents.
Explainability The capacity of an AI system (or individual agent) to produce human-understandable justifications or rationales for its decisions and actions, often via LIME, SHAP, or decision provenance graphs.
Foundation Model (LLM) A pretrained large language model (e.g., GPT-4, LLaMA) that serves as the “brain” of each agent, providing generative capabilities, reasoning, and tool-calling support.
Shared Memory (Persistent Memory) A centralized or distributed store (often a vector database) where agents write and retrieve contextual information, enabling long-term planning and consistency across iterations.
ModelOps The practice of managing AI models (and agent prompts) throughout their lifecycle—development, deployment, monitoring, and retirement—with version control, CI/CD testing, and drift detection.
Application Security Safeguards and best practices (e.g., prompt sanitation, authentication, sandboxing) designed to protect agentic systems from prompt injection, identity spoofing, and lateral exploits.
Model Privacy Techniques (e.g., differential privacy, homomorphic encryption, secure enclaves) that ensure sensitive data—either during training or inter-agent communication—remains protected in multi-agent workflows.
Prompt Injection A security exploit in which an attacker crafts input containing hidden instructions that corrupt an agent’s reasoning or propagate malicious commands through agent interactions (“prompt infection”).
Retrieval-Augmented Generation (RAG) A framework where agents query an external knowledge store (e.g., vector database) to fetch relevant documents or facts, then condition their LLM responses on that retrieved context.
Role-Specialized Coordination An architectural pattern in which each agent is assigned a specific function (e.g., planner, verifier, coder) and collaborates via structured communication protocols to achieve complex tasks.
Decision Provenance Graph A graph-based representation that traces data flows and decision steps across multiple agents, enabling post-hoc auditing and system-level interpretability.
Tool-Use Interface The mechanism by which an agent issues structured commands (e.g., API calls, code execution) to external services or environments and incorporates the results back into its reasoning.
Trust Score A composite metric that quantifies an agent’s reliability, alignment with user goals, and consistency over time, often combining accuracy, safety-violation rates, and calibration of confidence.
Composite Metric An aggregate evaluation score (e.g., a weighted sum of trustworthiness, explainability, user-centered performance, and coordination metrics) used to benchmark different agentic systems.
Figure A.1: Traditional AI Agent vs. Agentic AI.