### v5

# TRiSM for Agentic AI: A Review of Trust, Risk, and Security Management in LLM-based Agentic Multi-Agent Systems

Abstract
Agentic AI systems, built upon large language models (LLMs) and deployed in multi-agent configurations, are redefining intelligence, autonomy, collaboration, and decision-making across enterprise and societal domains. This review presents a structured analysis of Trust, Risk, and Security Management (TRiSM) in the context of LLM-based Agentic Multi-Agent Systems (AMAS). We begin by examining the conceptual foundations of Agentic AI and highlight its architectural distinctions from traditional AI agents. We then adapt and extend the AI TRiSM framework for Agentic AI, structured around key pillars: Explainability, ModelOps, Security, Privacy and their Lifecycle Governance, each contextualized to the challenges of AMAS. A risk taxonomy is proposed to capture the unique threats and vulnerabilities of Agentic AI, ranging from coordination failures to prompt-based adversarial manipulation. To make coordination and tool use measurable in practice, we propose two metrics: the Component Synergy Score (CSS), which captures inter-agent enablement, and the Tool Utilization Efficacy (TUE), which evaluates whether tools are invoked correctly and efficiently. We further discuss strategies for improving explainability in Agentic AI, as well as approaches to enhancing security and privacy through encryption, adversarial robustness, and regulatory compliance. The review concludes with a research roadmap for the responsible development and deployment of Agentic AI, highlighting key directions to align emerging systems with TRiSM principles-ensuring safety, transparency, and accountability in their operation.
keywords:
Agentic AI, LLM-based Multi-Agent Systems, TRiSM, AI Governance, Explainability, ModelOps, Application Security, Model Privacy, AI Agents, Trustworthy AI, Risk Management, AI Safety, Privacy-Preserving AI, Adversarial Robustness, Human-in-the-Loop
1 Introduction
AI governance has moved from aspiration to obligation. The EU Artificial Intelligence Act [EU_AI_Act2025] entered into force on 1 August 2024 and applies progressively (e.g., prohibited AI practices and AI literacy obligations from 2 February 2025, and governance rules plus obligations for general-purpose AI models from 2 August 2025) [^3] EU AI Act; meanwhile, organizations are adopting management frameworks such as ISO/IEC 42001:2023 [iso42001] and the NIST AI Risk Management Framework (AI RMF 1.0) [ai2023artificial] to operationalize risk controls. In parallel, enterprise use of AI continues to expand, for example, recent global surveys [McKinsey_StateOfAI2025, Stanford_HAI_AIIndex2025, Gartner_AI_TRiSM2024] report that a substantial majority of organizations use AI in at least one business function. Critically, the emerging paradigm of AI agents, which are systems that autonomously plan, reason, and execute complex tasks, introduces qualitatively different risks than traditional AI applications. These trends demand system-level trust, risk, and security management tailored to agentic systems.
Figure 1: Agentic AI research growth on arXiv (2019–2024), showing multi-agent systems and LLM-based agent publications with marked acceleration following ChatGPT’s release in November 2022.
Figure 2: Google search interest for “AI Agents” vs “Agentic AI” (2022–2024), illustrating rising public awareness with key milestones including ChatGPT launch and Gartner’s recognition of Agentic AI as the top technology trend for 2025.
An AI agent is a computational entity that perceives its environment and acts to achieve goals [russell1997rationality]. What were once task-specific and largely deterministic programs have rapidly evolved into LLM-powered agentic systems with planning, tool use, and persistent memory [sapkota2025ai]. In this paper, we refer to such LLM-based, coordinating agent systems as Agentic Multi-Agent Systems (AMAS). These systems orchestrate multiple specialized agents to address complex tasks, coordinate functions and roles, and adapt workflows through interactions with tools, users, and other agents.
TRiSM for Agentic AI
AMAS exhibit emergent and often opaque behaviors, making them susceptible to cascading errors, biased decisions, and unintended interactions [yang2024multi]. These risks are amplified when agents plan, coordinate, and call external tools. Beyond traditional ML concerns (safety, fairness, and interpretability), Agentic AI must contend with threats from tool integrations, prompt-level attacks, memory poisoning, impersonation, and privacy leakage in collaborative settings [lim2012memory, fang2025trustworthyaisafetybias]. To address these multifaceted system-level risks, AI Trust, Risk, and Security Management (TRiSM) frameworks provide integrated approaches emphasizing governance, explainability, security, privacy, and lifecycle controls [IBM2025AITRiSM, habbal2024artificial].
Integrating TRiSM principles into the design and deployment of Agentic AI systems is increasingly urgent. Figure 1 demonstrates explosive growth in arXiv publications, with multi-agent systems papers increasing from 890 in 2019 to 18,500 in 2024, while LLM-based agent papers surged from near zero to 9,800 following ChatGPT’s release. This academic interest is mirrored by public awareness, as shown in Figure 2, where Google search trends for both “AI Agents” and “Agentic AI” reached peak levels by late 2024. Gartner’s designation of Agentic AI as a top strategic technology trend for 2025, alongside AI governance platforms within its TRiSM framing [^4] Gartner Market Guide for AI TRiSM, and its 2026 emphasis on multiagent systems, digital provenance, and AI security platforms  [^5] Gartner Top Technology Trends 2026 reflect a growing consensus: as AI systems gain autonomy, robust trust, risk mitigation, and security controls become indispensable. We emphasize that a TRiSM perspective is essential to make AMAS deployable in sensitive domains.
1.1 Necessity and Scope of this Survey
Much of the Agentic AI literature prioritizes agent modeling, planning, and collaboration, while comparatively less attention is given to adversarial robustness, lifecycle governance, decision provenance, and system-level explainability. This survey addresses this gap by focusing on AMAS [^6] Throughout, AMAS denotes multi-agent systems built on LLMs. that demonstrate autonomous planning, tool use, memory retention, and emergent reasoning. As deployments expand in healthcare, finance, science, and public services, the absence of an integrated TRiSM perspective exposes stakeholders to opaque decision pathways and unmanaged risks. To our knowledge, no prior survey examines LLM-based multi-agent systems explicitly through the lens of trust, risk, security, and governance. Table 1 situates our contribution relative to prior work.
1.2 Contributions
The contributions of this study are manifold:
  • Unified TRiSM framework for Agentic AI. We present a system-level Trust, Risk and Security Management (TRiSM) framework tailored to AMAS (Section 4). The framework integrates explainability, ModelOps, security, privacy and lifecycle governance, contextualizing each pillar to AMAS workflows and failure modes.
  • Risk taxonomy for AMAS. We synthesize a taxonomy of threats unique to AMAS, including prompt injection, memory poisoning, collusive failure, emergent misbehavior and tool-use abuse, and map these risks to corresponding controls (Section 3).
  • Evaluation template and new metrics. We propose a holistic evaluation template covering trustworthiness, explainability, user-centered performance and inter-agent coordination. To operationalize this template, we introduce the Component Synergy Score (CSS) for measuring collaboration quality across agents and Tool Utilization Efficacy (TUE) for assessing the correctness and efficiency of tool calls (Section 5).
  • Technique mapping and gap analysis. We survey and map existing explainability methods, prompt-hygiene practices, decision-provenance tools, sandboxing strategies and ModelOps pipelines to the AMAS setting, revealing where existing approaches can be adapted and where significant gaps remain (Section 7).
Building on these contributions, we outline a research roadmap for developing scalable, verifiable and regulation-aligned agentic systems. Key directions include adversarial robustness, governance protocols, standardized benchmarks for trustworthiness and coordination, and cross-disciplinary collaboration.
1.3 Related Work
Prior surveys on LLM-based multi-agent systems mainly emphasize architectures, agent capabilities, communication, and domain applications [guo2024largelanguagemodelbased, chen2025surveyllmbasedmultiagentsystem, yan2025selftalkcommunicationcentricsurveyllmbased, tran2025multiagentcollaborationmechanismssurvey]. In parallel, broader trustworthy/responsible AI surveys discuss themes aligned with TRiSM, such as alignment, fairness, and privacy defenses; but typically focus on general ML systems and do not model AMAS-specific dynamics (e.g., inter-agent information sharing, tool-mediated actions, and emergent coordination) [fang2025trustworthyaisafetybias, raza2024exploring]. As a result, key gaps remain in adversarial robustness, governance and accountability for multi-agent deployments, and explainability at scale. In comparison to these works (Table 1), our study addresses these gaps by mapping TRiSM principles to AMAS and operationalizing governance, explainability, security, privacy, and lifecycle controls for multi-agent settings.
Table 1: Comparison with representative related surveys on LLM-based agentic systems.
Survey Threats Lifecycle Explain. TRiSM LLM-Spec. Apps. Actionable
Guo et al. (2024) [guo2024largelanguagemodelbased] ~ ~
Chen et al. (2024) [chen2025surveyllmbasedmultiagentsystem] ~ ~
Yan et al. (2025) [yan2025selftalkcommunicationcentricsurveyllmbased] ~ ~
Tran et al. (2025) [tran2025multiagentcollaborationmechanismssurvey] ~
Lin et al. (2025) [lin2025creativityllmbasedmultiagentsystems] ~ ~
Fang et al. (2025) [fang2025trustworthyaisafetybias] ~ ~ ~
Xi et al. (2025) [xi2025rise] ~ ~
Luo et al. (2025) [luo2025large] ~ ~ ~
Zou et al. (2025) [zou2025llm] ~ ~
Wang et al. (2025) [wang2025surveyllmbasedagentsmedicine] ~ ~
Wang et al. (2025) [11081880] ~
Yu et al. (2025) [yu2025survey] ~
This Survey (2025)
Legend: ✓ = explicitly addressed; ~ = partially addressed; ✗ = not addressed.
1.4 Literature Review Methodology
We conducted a structured literature review to synthesize AI TRiSM principles for AMAS. Following established systematic review guidance [boland2017doing, keele2007guidelines], we describe our data sources, search strategy, screening criteria, and classification schema.
Research Questions (RQs)
The review was guided by the following three questions:
  • RQ1: What trust, risk, and security challenges are specific to AMAS?
  • RQ2: Which governance, explainability, security, privacy, and lifecycle (TRiSM) controls are proposed or adapted for AMAS?
  • RQ3: How are AMAS evaluated (datasets, tasks, metrics), and where are the gaps?
Search Strategy
We systematically searched major digital libraries: IEEE Xplore, ACM Digital Library, SpringerLink, arXiv, ScienceDirect, and Google Scholar. The search covered publications from January 2022 to December 2025, with influential foundational works prior to 2022 included for theoretical context. We used a two-block Boolean query: (agentic/AMAS terms) AND (TRiSM terms).
("agentic AI" OR "multi-agent systems" OR "multi-agent LLMs" OR "AI agents"
OR  "autonomous agents" OR "intelligent agents" OR "collaborative AI" OR
 "distributed AI" OR "LLM-based agents" OR "agent coordination" OR
 "agent-based modeling" OR "swarm intelligence")
AND
("trust" OR "trustworthiness" OR "risk" OR "security" OR "safety" OR
 "governance" OR "oversight" OR "compliance" OR "explainability" OR
 "interpretability" OR "transparency" OR "privacy" OR "data protection" OR
 "robustness" OR "accountability" OR "ethical AI" OR "fairness" OR
 "adversarial attacks" OR "regulatory compliance" OR "bias mitigation" OR
 "system reliability")
Screening and Eligibility
We applied a two-stage screening process (title/abstract, then full text).
Inclusion criteria:
  • Published between 2022–2025 (plus influential foundational papers published before 2022).
  • Focused on agentic AI, multi-agent, or LLM-powered systems and addressed at least one TRiSM aspect (trust, risk, security, governance, explainability, or privacy).
  • Peer-reviewed papers, widely cited preprints, technical reports from reputable venues or organizations (e.g., IEEE/ACM conferences and journals, standards bodies, and government sources).
Exclusion criteria:
  • Traditional rule-based or symbolic agents without LLM-based autonomy or coordination.
  • Papers limited to low-level model training/architecture details with no agentic behavior or TRiSM relevance.
  • Non-English papers or items with insufficient bibliographic information (e.g., missing abstract or venue details).
For empirical studies, we verified that the threat model and evaluation were clearly described and assessed reproducibility when code or data were available. For survey papers, we evaluated coverage, structure, and treatment of risks and governance. After title/abstract screening, we shortlisted 250 papers; full-text screening yielded 180 primary studies covering one or more TRiSM pillars, supplemented by relevant standards and technical reports for regulatory context.
Quality Assurance
To ensure reliability, we adapted quality criteria from established guidelines [kitchenham2004procedures]. Each paper was evaluated on: (1) clear objectives for AMAS; (2) well-documented methods/experiments or architecture (reproducibility where applicable); (3) substantive engagement with at least one of trust, risk, security, explainability, or privacy; (4) empirical, theoretical, or normative contributions relevant to agentic AI governance. Each criterion was rated low, medium, or high; studies rated low on more than two criteria were excluded or flagged as contextual only. Quality ratings were assigned independently by four reviewers; discrepancies were resolved via discussion.
Bibliometric Overview
The results of our screening process reveal the rapid acceleration of agentic AI research. The publication landscape (Figure 3) shows journal articles dominating at 61.2%, with leading venues and publishers detailed in Figure 4. Key research themes, including large language models, multi-agent systems, and security, are visualized in Figure 5a, while Figure 6 confirms the unprecedented growth of research activity in recent years.
(a) Publication trends (1985–2025)
(b) Publication type distribution
Figure 3: Publication landscape: (a) Long-term trends showing accelerating research output over four decades; (b) Distribution of publication types, with journal articles dominating at 61.2%.
(a) Top publication venues
(b) Publisher distribution
Figure 4: Publication outlets: (a) Top venues including NeurIPS, ACM Computing Surveys, and AAAI; (b) Publisher distribution led by Elsevier, Springer, and IEEE.
(a) Research topics and keywords
(b) Publication type evolution over time
Figure 5: Research themes and evolution: (a) Word cloud highlighting key topics including large language models, multi-agent systems, security, and explainability; (b) Temporal distribution of publication types.
Figure 6: Comparison of recent publications (2020–2025) with historical distribution by period, demonstrating the rapid acceleration of agentic AI research in recent years.
Figures 7 and 8 provide a roadmap for the paper, which progresses from architectural foundations to risk analysis, evaluation, and governance.
TRiSM for Agentic AIFundamentalsThreats & RisksTRiSM FrameworkEvaluation
$\circ$ AMAS Architecture
$\circ$ LLM Core
$\circ$ Planning/Reasoning
$\circ$ Memory Module
$\circ$ Tool Interface
$\circ$ Adversarial Attacks
$\circ$ Data Leakage
$\circ$ Agent Collusion
$\circ$ Emergent Behavior
$\circ$ Memory Poisoning
$\circ$ Explainability
$\circ$ ModelOps
$\circ$ Application Security
$\circ$ Model Privacy
$\circ$ Governance
$\circ$ Trustworthiness
$\circ$ CSS Metric
$\circ$ TUE Metric
$\circ$ Compliance
$\circ$ Auditability
Figure 7: Taxonomy of TRiSM for LLM-based Agentic Multi-Agent Systems.
Figure 8: TRiSM taxonomy for Agentic AI (AMAS) presented in this review. The paper progresses through six phases: (i) Fundamentals of Agentic AI: distinguishing AMAS from traditional agents and establishing LLM-based architectures (Section 2); (ii) AI Trust, Risk, and Security Management (TRiSM) framework pillars: Explainability, ModelOps, Application Security, Model Privacy, and Lifecycle Governance, introduced conceptually (Section 4); (iii) Trust management via explainability techniques (Section 4.2.1); (iv) Threats and risks taxonomy with real-world case studies (Section 3); (v) Evaluation metrics encompassing trustworthiness, coordination, and composite measures (Section 5); (vi) Technical implementation of security/privacy mechanisms (Section 6); and (vii) Compliance and governance integration (Section 7).
2 Fundamentals of Agentic AI Systems
In this section, we present the fundamentals of Agentic AI systems.
Table 2: Comparison of Traditional AI Agents and Agentic AI Systems
Dimension Traditional AI Agents Agentic AI Systems
Autonomy Reactive with fixed action sequences Goal-driven with adaptive behavior
Cognitive Foundation Symbolic logic, finite state machines Foundation models (large language models)
Domain Scope Narrow, task-specific applications Broad, cross-domain capabilities
Reasoning Mechanism Rule-based, deterministic inference Multi-step, contextual reasoning (chain-of-thought)
Collaboration Isolated agents, manual task decomposition Role-specialized agents, automated coordination
Memory Architecture Episodic, session-bound storage Persistent memory (vector databases, long-term memory)
Orchestration Hard-coded, predefined workflows Dynamic orchestration via meta-agents
Tool Integration Static, hand-crafted interfaces Dynamic planning with API/function calling
Context Handling Bounded context with heuristic rules Memory-augmented retrieval (RAG)
Learning Paradigm Rule updates, fixed knowledge base Emergent learning (fine-tuning, reinforcement learning)
User Interaction Scripted, template-based dialogue Natural language understanding and generation
Representative Systems MYCIN [daniel1997cadiag], ELIZA [walters2025eliza], SOAR [laird1987soar], ACT-R [taatgen2006modeling] AutoGen [autogen_agents_tutorial], ChatDev [qian2023chatdev], MetaGPT [hong2023metagpt]
2.1 Traditional AI Agents vs. Agentic AI
Traditional agents rely on predefined rules, heuristics, or deterministic logic and typically perform well in narrow, well-specified environments [muscettola1998remote, hannebauer1999formal, govaerts2010development]. Cognitive agents extend this paradigm by modeling human-like functions (perception, memory, decision-making) through cognitive architectures such as SOAR and ACT-R [langley2006cognitive, laird1987soar, anderson1993act]; recent work further categorizes their knowledge into concepts, skills, processes, and motives [langley2025spatial]. However, these approaches were largely developed for physical operational settings (e.g., robotic planning and control), whereas Agentic AI generalizes autonomy to both physical and non-physical environments.
In contrast, Agentic AI systems leverage foundation models (primarily LLMs) to support adaptive, goal-directed behavior via (i) multi-agent coordination, (ii) persistent context through memory, and (iii) dynamic orchestration of roles and tool use [qian2023chatdev, hong2023metagpt, borghoff2025human, schneider2025generative]. This enables long-horizon task execution and cross-domain generalization beyond traditional reactive pipelines [de2025open]. Table 2 summarizes these differences; in this survey, we focus on LLM-based Agentic AI systems.
2.2 Multi-Agentic AI System (AMAS) Architecture
Figure 9: Architecture of LLM-based Agentic Multi-Agent System (AMAS), illustrating key functional layers: Perception/Input Layer (text, image, audio processing), Cognitive/Reasoning Layer (goal-setting, planning, decision-making), Action/Execution Layer (digital and physical task execution), Learning Module (supervised and reinforcement learning), Communication and Collaboration (agent messaging and coordination), Data Storage & Retrieval (centralized/distributed databases), and Monitoring & Governance (ethical oversight, observability, and compliance mechanisms).
As shown in Fig. 9, AMAS coordinate multiple LLM-based agents through a Communication Middleware [weyns2009agent], a Task Manager/Orchestrator [langgraph2024], and a World Model/Shared Memory [guo2024largelanguagemodelbased] that stores system state and task artifacts. Human oversight is provided via a Human-in-the-Loop (HITL) interface [cranshaw2017calendar]. To support trustworthy operation, a Trust and Audit module records actions and tool usage [kaur2024building], while a Security Gateway and Privacy Management Layer enforce access control and reduce leakage across agents and tools [zhong2023towards, feretzakis2024privacy]. An Explainability Interface provides interpretable rationales for multi-agent decisions to support transparency and trust calibration [singh2025explainable].
Communication & Collaboration (Agent Communication and Coordination)
Beyond basic messaging, recent work formalizes AMAS communication through explicit protocols. Google’s Agent-to-Agent (A2A) [^7] A2A Protocol supports discovery, capability advertisement, delegation, and result exchange, improving interoperability and auditability. Related efforts such as the Agent Network Protocol (ANP) [^8] Agent Network Protocol take a network-centric view with standardized interfaces for coordination and trust establishment. From a TRiSM perspective, protocolized communication exposes control points for authentication, authorization, logging, and lifecycle governance. Emerging semantic-aware communication [guo2024survey, wang2025internet] can improve coordination by encoding intent, but introduces risks such as semantic spoofing and privacy leakage.
Cognitive/Reasoning Layer (LLM Core + Planning)
At the center of an AMAS, a foundation model (typically an LLM) acts as the decision controller, taking a user goal plus role and tool specifications as input and generating the next actions (e.g., messages, tool calls, or structured function outputs) [Saravia_LLM_Agents_2024]. In agent frameworks, this controller drives an iterative loop of planning, acting, and updating context based on tool results and feedback.
Planning and Reasoning Module
To achieve complex goals, AMAS employ planning mechanisms that decompose objectives into sub-tasks and interleave reasoning with execution and feedback. This can be implemented through prompting strategies (e.g., CoT/ToT) or through explicit plan–act–observe loops such as ReAct [yao2023reactsynergizingreasoningacting, Saravia_Prompt_Engineering_Guide_Web_2024]. The plan is typically revised as new observations arrive, enabling adaptive long-horizon behavior.
Memory Module
AMAS integrate memory to maintain context across steps and sessions. Short-term memory is maintained in the working context, while long-term memory stores salient facts and past events for later retrieval (often via embedding-based search) [yang2024multi]. Memory improves continuity and reduces repetition, but also increases exposure to stale context and poisoning risks.
Learning Module
Some AMAS include learning components that adapt behavior over time (e.g., from feedback, rewards, or task outcomes), such as updating policies, prompt templates, or routing strategies based on past performance. In multi-agent settings, learning can amplify errors through feedback loops and introduce risks such as drift, poisoning, and emergent misalignment, motivating continuous monitoring and gated updates under TRiSM controls.
Data Storage & Retrieval
AMAS typically combine structured stores for system state and task metadata with vector-based retrieval to support RAG and long-term memory. Because data may be shared across agents and tools, storage and retrieval require strong access control, encryption, and audit logging to reduce leakage and support accountability [he2024security].
Tool-Use Interface
To extend capabilities beyond text generation, AMAS agents use a tool-use interface that enables structured calls to external systems (e.g., web search, APIs, code execution, and databases) [schick2023toolformerlanguagemodelsteach]. Tools and call schemas are specified in the agent configuration; when needed, the LLM emits a structured command that is executed externally, and the returned output is incorporated as a new observation for subsequent reasoning. Tool-routing frameworks such as MRKL [karpas2022mrklsystemsmodularneurosymbolic] exemplify this design by dispatching requests to specialized expert modules under an LLM controller, while training-based approaches (e.g., Toolformer) teach models to generate API calls during generation [schick2023toolformerlanguagemodelsteach].
Perception and Environment Interface
For agents interacting with dynamic environments (e.g., web interfaces, simulated worlds, or physical systems), an observation–action interface is essential [liu2025agent]. Perception modules translate raw inputs (sensor streams, images, or textual state) into representations the LLM can process; the agent’s selected actions are then executed in the environment, and resulting state changes are returned as new observations. This loop supports iterative sense–plan–act behavior until a task is completed or halted. In multimodal or robotic settings, additional perception models (e.g., vision encoders) may be included, while the core interaction pattern remains the same.
Monitoring and Governance Layer
A monitoring and governance layer provides system-wide oversight by collecting logs, traces, and metrics, enforcing policy constraints, and triggering alerts or interventions when anomalous behavior is detected. From a TRiSM perspective, this layer enables auditability and lifecycle governance across agent interactions, tool calls, and memory updates.
Integration and Autonomy
Together, these modules form an integrated architecture in which the LLM coordinates planning, tool use, and memory to act in an environment and update context from feedback [acharya2025agentic]. Table LABEL:tab:agentic_architectures summarizes representative systems and maps them to five design axes (LLM core, planning/reasoning, memory, tool use, and environment interface).
Key Insights
Overall, AMAS combine planning, tool use, memory, and environment interaction into a closed-loop system, which expands the attack surface and complicates accountability. Protocolized communication and monitoring create practical control points for authentication, logging, and governance, motivating TRiSM-oriented analysis for multi-agent deployments.
Table 3: Comparison of representative LLM-based Agentic AI frameworks across key design axes.
Framework Core LLM Planning Memory Tool Use Notable Features
AutoGPT [yang2023auto] GPT-4 Self-looped CoT Vector DB OS shell + web Fully autonomous goal loop
BabyAGI [nakajima_babyagi_archive_2024] GPT-3.5/4 Task queue In-memory Web search Minimal task generator
GPT Engineer [gpt-engineer] GPT-4V Spec-to-code File cache Python REPL End-to-end code generation pipeline
LangGraph [wang2024agent] Model-agnostic FSM via graph Persistent nodes Custom modules Visual orchestration of agent graphs
AutoGen [wu2023autogen] GPT-4 Multi-agent PDDL JSON/DB API calls Modular, reusable agent templates
MRKL [karpas2022mrklsystemsmodularneurosymbolic] LLaMA/GPT Prompt router N/A Math + search tools Neuro-symbolic expert routing
Reflexion [shinn2023reflexionlanguageagentsverbal] GPT-3.5/4 Retry-reflect loop Episodic buffer Same as base agent Verbal self-improvement via reflection
MetaGPT [hong2024metagpt] GPT-4 SOP workflow YAML state Git CLI Structured roles for software engineering
Voyager [wang2023voyageropenendedembodiedagent] GPT-4 Auto skill tree Task DB Minecraft API Lifelong open-ended learning in environments
WebVoyager [he2024webvoyager] LLaVA-1.6 ReAct JSON store Browser actions Multimodal web interaction via vision and text
HuggingGPT [shen2023hugginggpt] ChatGPT Task-plan-select Log store Hugging Face models External model orchestration by LLMs
CAMEL [li2023camel] GPT-4 Role-play CoT Dialogue log Chat only Multi-agent role simulation via dialogue
ChatDev [qian2024chatdevcommunicativeagentssoftware] GPT-4 Chat chain File repo Unix tools Simulated software development workflow
CrewAI [crewAI] Any Declarative plan Optional DB Python modules Lightweight, LangChain-free framework
AgentVerse [chen2023agentverse] Model-agnostic Config graph Redis/KV store Plugin API Multi-agent simulation and task solving
OpenAgents [xie2023openagentsopenplatformlanguage] Model-agnostic Agent scripts MongoDB Web plugins Open platform with public hosting
SuperAGI [superagi] GPT-4 DAG workflow Postgres Pinecone Concurrent agents for production use
Semantic Kernel [microsoft_semantickernel] Model-agnostic Agent planning framework Memory & context mgmt Skill-based plugins, OpenAPI Enterprise focus, .NET integration, security, vector-DB support
OpenAI Swarm [openai_swarm] GPT-4o Agents + handoffs loop Stateless (no LT memory) Python functions, agent handoffs Minimalist, controllable, educational multi-agent orchestration
OpenAI Agents SDK [openai_agents_sdk] Provider-agnostic Agent loops, handoffs, deterministic flows Sessions manage conversation history Tools + guardrails, tracing Multi-agent workflows, safety guardrails, extensibility
Strands Agents [strands_agents] Model-agnostic Agent loop with multi-agent & streaming N/A Python tools, hot-reloading, MCP tools Lightweight SDK, scalable autonomous workflows
LlamaIndex Agents [llamaindex_agents] Model-agnostic Task breakdown + planning Task memory module External tools & parameters Pre-built agent/tool architectures, custom workflows
3 Threats and Risks in Agentic AI
This section summarizes major threats in Agentic AI and organizes emergent AMAS risks into a compact taxonomy.
3.1 Threats in Agentic AI
Agentic AI systems introduce security and reliability concerns beyond single-agent LLM pipelines:
  • Autonomy abuse. Agents with delegated authority may misinterpret objectives or execute harmful plans due to reasoning errors or manipulated inputs.
  • Persistent memory. Long-term memory improves continuity but creates a contamination channel: injected or low-quality content can persist and propagate through shared stores without strong versioning and audit controls.
  • Agent orchestration. Centralized or distributed orchestrators mediate roles and workflows; compromise or misconfiguration can misroute information and trigger cascading failures.
  • Goal misalignment and exposure. Poorly scoped or exposed goal representations can enable prompt extraction, misalignment, or adversarial steering, especially when third-party tools are invoked.
  • Tool misuse and external API exploits. Tool-enabled agents introduce new attack surfaces (e.g., unsafe code execution, costly actions, policy violations); sandboxing and invocation tracing are often inconsistent across open-source frameworks.
  • Multi-agent collusion or drift. Agents can converge on mutually reinforcing errors (groupthink/mode collapse) or drift in self-reinforcing loops, complicating interpretability and safety alignment.
3.2 Risks Taxonomy in AMAS
We group AMAS risks into four classes—adversarial attacks, data leakage, agent collusion, and emergent behaviors—as shown in Fig. 10.
Figure 10: Overview of key risks and threats in agentic AI systems. The left panel outlines system-level threats (e.g., autonomy abuse, persistent memory, orchestration). The right panel summarizes emergent AMAS risk classes arising from multi-agent interaction, memory persistence, and decentralized decision-making.
  • Adversarial attacks. AMAS are vulnerable to prompt injection, engineered reasoning traps, and other manipulations that can cascade across agents; e.g., role-swapping attacks in ChatDev exploit inter-agent dependencies [miao2025autonomous, qian2023chatdev].
  • Data leakage. Shared memory and inter-agent communication increase the chance of unintended disclosure of sensitive information, particularly when boundaries, sanitization, or access controls are weak [tran2025multi].
  • Agent collusion and mode collapse. Collaboration mechanisms can produce groupthink, where agents reinforce the same flawed assumptions and converge on suboptimal outputs [feng2025multi].
  • Emergent behaviors. Complex interactions among agents, tools, and memory can yield unpredictable strategies that evade conventional testing (e.g., shortcutting verification), motivating continuous monitoring and adaptive controls [dawid2025agentic].
Real-World Examples
The following examples illustrate how these risks can appear in practice:
  • Prompt leakage. In AutoGPT-style setups, recursive prompt augmentation and weak memory controls can expose sensitive content when stored tokens later surface in summaries or logs [Euler2023AutoGPTRCE, invicti_prompt_injection_ebook].
  • Collusive failure. In ChatDev-like workflows, shared planning errors can propagate when agents validate each other without external checks or objective feedback [qian2024chatdevcommunicativeagentssoftware].
  • Simulation attack in swarms. Misleading environment assumptions can induce coordination failures (e.g., congestion and task incompletion) in multi-agent planning settings [StrDorFri2024, ligot2022using].
  • Memory poisoning. Feedback buffers can be polluted (e.g., sarcastic or adversarial signals) and later used for policy updates, degrading behavior without validation gates [atkins2023memorypoisoning].
  • System prompt drift. Self-appended context without versioning can shift prompts over time and induce hallucinated goals or misaligned behaviors [liu2023lostmiddlelanguagemodels].
Key Insights
We identified AMAS-specific risks that arise from autonomy, shared memory, orchestration, and tool use, and organized them into four emergent classes (Fig. 10). Across the taxonomy, failures can cascade across agents and persist through memory substrates, while emergent behaviors may evade conventional test suites. These observations motivate defense-in-depth controls spanning prompt hygiene, memory sanitization, orchestrator authentication, tool sandboxing, and continuous monitoring.
4 TRiSM Framework for Agentic Multi-Agent Systems
4.1 Overview of AI TRiSM
The TRiSM Framework offers a structured lens for evaluating and governing AI systems, especially those that exhibit autonomous, agentic behaviors. Below, we discuss core components of TRiSM framework.
  • AI Trust Management Establishing trust in AI systems necessitates a foundational commitment to transparency, accountability, and fairness. Trust management focuses on embedding mechanisms that enable AI models to generate interpretable and justifiable outputs, thereby facilitating meaningful human oversight [borghoff2025human]. This involves the adoption of explainability tools, the implementation of governance protocols, and alignment with ethical standards to address systemic biases and ensure equitable outcomes across diverse user groups.
  • AI Risk Management Effective risk management within AI systems entails the systematic identification, analysis, and mitigation of potential harms arising from the deployment and operation of these technologies. This includes technical risks such as algorithmic bias [raza2024exploring], model errors, and data leakage, as well as societal risks involving fairness, safety, and misuse [fang2025trustworthyaisafetybias]. Structured risk assessment frameworks enable stakeholders to anticipate failure modes and design mitigation strategies that enhance system robustness and reduce the likelihood of adverse impacts.
  • AI Security Management Security management in the context of AI is concerned with safeguarding models, data, and infrastructure from adversarial threats and unauthorized access. This encompasses both conventional cybersecurity practices, such as encryption, authentication, and access control, and AI-specific considerations like adversarial example detection, model extraction prevention, and secure model deployment pipelines [raza2025justhumansneedvaccines]. Continuous monitoring and threat modeling are critical for identifying vulnerabilities and ensuring the resilience and integrity of AI systems over time
Table 4: TRiSM pillars (Part 1): Core controls, techniques, and key risks mitigated.
Pillar Core Controls Techniques / Patterns Key Risks Mitigated
Explainability / Trust CoT logging; layered-CoT; inter-agent traceability; explainer agent; role-based interpretability; prompt audit trails; attention viz; decision provenance Layered-CoT [sanwal2025layered]; LIME/SHAP [ribeiro2016should, lundberg2017unified]; decision-provenance graphs [tupayachi2024towards, bai2024derived]; multi-agent SHAP [tian2025outlook]; attention maps [liu2024ava, hu2024overview]; role-based interpretability [tomsett2018interpretable]; RAG-linked justifications [ghafarollahi2024sciagents]; prompt/action logs [gu2024artificial, chin2025automating] Opaque chains; unverifiable reasoning; hidden inter-agent effects; spurious correlations; user miscalibration
ModelOps (Lifecycle) Versioning; lineage; CI/CD safety gates; rollout/rollback; multi-agent simulation; hierarchical monitoring; usage/cost SLOs; ITSM/IAM integration Prompt & agent-config versioning; pre-deploy safety tests; regression/drift checks; sandbox simulation; hierarchical monitors (agent/group/system) [wang2025megaagent] Silent regressions; drift; bias re-introduction; unsafe changes; cost blowouts; observability gaps
Application Security Prompt hygiene; secure prefixes; sandboxed tools; allowlisted actions; plan-then-execute; least privilege (RBAC); authN/authZ; anomaly detection; red-teaming Prompt-injection patterns [beurer2025design, qiao2024agent]; plan-then-execute; action selector; tool isolation; output filters; cross-agent cross-checks; adversarial training [standen2025adversarial]; HITL holds Prompt injection; tool abuse; data exfiltration; cross-modal injection; lateral movement; jailbreaks
Privacy & Data Protection DP; anonymization; pseudonymization; minimization; MPC; HE/FHE; TEEs (SGX/SEV/TrustZone); encryption; access logs; retention/consent DP budgets [dwork2006calibrating]; k-anonymity [sweeney2002k]; MPC/PUMA [lindell2009secure]; HE/FHE pipelines [gentry2009fully]; TEEs [costan2016intel]; memory scoping; PII detectors Memorization; leakage via shared memory/messages; unauthorized access; weak consent/retention
Governance Human oversight; accountability; auditability; DPIA; policy/change control; role clarity; transparency notices; logging/traceability Oversight playbooks; two-person verification for high-risk; role-oriented modularity; audit trails; DPIA templates; policy gates Unclear responsibility; weak oversight; non-compliance; opaque decisions
Table 5: TRiSM pillars (Part 2): Evaluation facets, example systems, and governing standards.
Pillar Evaluation Facets & Metrics Example Systems Standards / Frameworks
Explainability / Trust Trustworthiness: calibration/consistency; Explainability: fidelity, stability; User-centric: task utility, cognitive load; Coordination: cross-agent attribution; Composite: multi-facet scorecards [raza_fair_2024] SciAgent [ghafarollahi2024sciagents]; MetaGPT [ke2025meta]; ChatDev [qian2023chatdev]; AutoGen [wu2023autogen] EU AI Act Arts 13–14 [eu_ai_act]; NIST AI RMF / GAI profile [ai2023artificial, ai2024artificial]; OECD AI [oecd_ai_catalogue]; ISO/IEC 24029-1 [iso24029-1]
ModelOps (Lifecycle) Trustworthiness: drift/robustness; User-centric: task success, latency; Composite: risk & cost scorecards; incident MTTR/MTTD MegaAgent [wang2025megaagent] ISO/IEC 42001 (AIMS); ISO/IEC 42005 [ISO_IEC_42005_2025]; ISO/IEC 23894 [ISOIEC23894_2023]; NIST AI RMF [ai2023artificial, ai2024artificial]; ModelOps [lefevre2022modelops, Sinha2025ModelOps]
Application Security Trustworthiness: robustness/adversarial resilience; Coordination: agent-of-agent validation; Composite: residual risk index; incident rates HITL review gates [chen2025human]; enterprise dashboards [chen2025human]; ChemCrow safety pauses [bran2023chemcrow] OWASP LLM Top-10 [owasp2024llmtop10, owasp2025agentic]; OECD Robustness [oecd_ai_catalogue]
Privacy & Data Protection Trustworthiness: privacy loss; User-centric: data-minimization utility; Composite: privacy-risk score; audit pass rate Encrypted memories; scoped retrieval; consented logs GDPR Art 25/5/35 [gdpr25]; CCPA/CPRA [oag_ccpa_2024]; HIPAA §164 [hipaa164]; ISO/IEC 27001/27701
Governance User-centric: oversight efficacy; Coordination: traceability across agents; Composite: compliance scorecards HITL confirm/override/stop; role-based review boards EU AI Act Art 14 [eu_ai_act]; NIST AI RMF [ai2023artificial]; OECD AI [oecd_ai_catalogue]; ISO/IEC 42001/42005; ISO/IEC 23894 [ISOIEC23894_2023]
4.2 Mapping Pillars of TRiSM with Agentic AI
Lifecycle Governance Plane (TRiSM Pillar) policies $\cdot$ accountability $\cdot$ compliance $\cdot$ oversight $\cdot$ traceability
Trust & Explainability (TRiSM Pillar) reasoning/provenance traces $\cdot$ explainer agent $\cdot$ user feedback $\cdot$ audit logs ModelOps / Lifecycle Management (TRiSM Pillar) prompt+agent config versioning $\cdot$ CI/CD safety gates $\cdot$ monitoring/rollback $\cdot$ incident workflow Application Security (TRiSM Pillar) prompt hygiene $\cdot$ allowlists $\cdot$ tool sandboxing $\cdot$ least privilege (RBAC) $\cdot$ anomaly detection Model Privacy & Data Protection (TRiSM Pillar) encrypted/scoped memory $\cdot$ data minimization $\cdot$ DP where applicable $\cdot$ access logs $\cdot$ retention/consent Planning & Reasoning (Agent Core Layer; non-TRiSM) multi-agent planning $\cdot$ task decomposition $\cdot$ CoT/ToT/ReAct $\cdot$ inter-agent coordination
User / HITL UI
confirm $\cdot$ override $\cdot$ stop
feedback loop
Secure API Gateway
authN/authZ $\cdot$ egress control
tool isolation $\cdot$ policy checks
External Tools
web $\cdot$ DB $\cdot$ code
enterprise APIs
Figure 11: TRiSM-aligned architecture for Agentic AI systems. The five TRiSM pillars are shown explicitly and in the same order as defined in the framework: Trust & Explainability, ModelOps, Application Security, Model Privacy & Data Protection, with a cross-cutting Lifecycle Governance plane. Planning & Reasoning is included as the agent-core execution layer (non-TRiSM).
Originally highlighted in industry guidelines for AI governance [litan2024ai], TRiSM also provides a structured approach to managing the unique challenges of AMAS. In this work, we focus on five key TRiSM pillars: Explainability, ModelOps, Application Security, Model Privacy, and Governance [kaur2024building, habbal2024artificial]. These five pillars also reflect established risk-surface decompositions in the NIST AI RMF [ai2023artificial], ISO/IEC 42001 [iso42001], and the EU AI Act [eu_ai_act], where explainability, lifecycle management, security, privacy, and governance are treated as distinct and non-overlapping risk domains. We summarize this mapping in Tables LABEL:tab:trism-pillars-1 and LABEL:tab:trism-pillars-2. Figure 11 illustrates how these five TRiSM pillars map onto the core architectural layers of an LLM-based agentic AI system.
Next, we discuss these pillars and how they can help in AMAS.
4.2.1 Explainability in Multi-Agent Decision Making
In AMAS, explainability is less about interpreting a single model and more about understanding how decisions emerge from interacting agents [yossef2024explainable]. In AMAS settings, outcomes often emerge from interactions among specialized agents rather than from a single model prediction [rosenfeld2019explainability]. As a result, explainability must account for both (i) individual agent decisions and (ii) the inter-agent dynamics that produce emergent system-level behavior. Recent work emphasizes that user trust depends on understanding each agent’s contribution and how collaboration is orchestrated [sanwal2025layered]. Traditional single-model XAI methods often overlook this multi-agent dimension, motivating new approaches tailored to AMAS.
Challenges. Explaining multi-agent decisions requires tracing which agent performed which sub-task, what information it used, and how inter-agent exchanges influenced the final outcome [de2025open]. Key challenges include:
  • Emergent behavior: agent interactions can produce behaviors that were not explicitly specified; explaining these patterns requires capturing and summarizing interaction chains.
  • Verification of intermediate steps: vanilla CoT prompting can yield plausible but unverified reasoning. Without systematic verification, small errors may propagate across agents and compound into incorrect final decisions.
Approaches for explainability in AMAS The following techniques adapt classical XAI ideas to AMAS:
  • Feature attribution (LIME/SHAP) for agents: LIME [ribeiro2016should] and SHAP [lundberg2017unified] can be extended to attribute an agent’s output to specific inputs (e.g., context spans, retrieved evidence, or inter-agent messages), clarifying which signals most influenced an agent’s action [bilal2025llms].
  • Counterfactual and causal analysis: “What-if” analysis evaluates how outcomes change when an agent’s contribution or a key condition is modified [gyevnar2023causal]. This helps isolate agent roles and identify causal dependencies in collaborative decision-making.
  • Reasoning-trace logging: Logging intermediate reasoning artifacts (e.g., CoT traces, tool calls, and inter-agent dialogues) creates an auditable trail of how decisions emerged [yao2023reactsynergizingreasoningacting]. Such traces reveal which agent contributed what and how partial conclusions accumulate into final outputs.
  • Natural-language explanations: An explainer agent can synthesize traces and attributions into human-readable narratives that answer questions such as “Why did the system choose action X?” Embedded explainers in workflows (e.g., DSM5AgentFlow) can improve transparency and trust [ozgun2025trustworthy].
  • Role-oriented modular architectures: Assigning explicit roles (e.g., reasoning, verification, summarization, credit assignment) improves inspectability because each stage is separable and reviewable [lo2025ai].
  • Explainable reward critics: In multi-agent RL, LLM-based critics can provide transparent reward decomposition and justification, supporting debugging and governance of agent incentives [nagpal2025leveraging].
Recent directions. Recent work proposes frameworks and techniques that embed explainability directly into multi-agent operation:
  • Layered Chain-of-Thought prompting: Layered-CoT decomposes reasoning into discrete layers, where each layer answers a sub-question [sanwal2025layered]. A specialized Reasoning Agent produces partial reasoning, while Verification and User-Interaction Agents cross-check intermediate outputs (e.g., against knowledge graphs or human feedback). This decomposition can localize errors, reduce contradictions, and promote validation before proceeding.
  • Integrated multi-agent workflows: Some systems integrate explanation into the workflow rather than adding it post hoc. For example, in a DSM-5 mental-health diagnosis workflow, one agent collects patient information while a diagnostician agent retrieves DSM-5 criteria and explicitly links conversational evidence to diagnostic rules, producing a self-documenting rationale [ozgun2025trustworthy].
  • Modular resume-screening systems: In recruitment, multi-agent pipelines separate extraction, evaluation, summarization, and formatting. The evaluation agent retrieves company-specific criteria via RAG, enabling recruiters to trace how candidate evidence maps to scores [lo2025ai].
  • Explainable credit assignment in multi-agent reinforcement learning: A centralized LLM critic can provide per-agent credit assignment and natural-language justification, explaining how global rewards are decomposed across agents [nagpal2025leveraging].
4.2.2 ModelOps : Lifecycle Management for LLM Agents
ModelOps (Machine Learning Operations) extends the MLOps paradigm beyond model training and deployment to include governance, risk management, and continuous oversight across AI systems [lefevre2022modelops]. In production, ModelOps provides end-to-end oversight by enforcing governance and compliance controls, including lineage, ownership, and metadata tracking to support regulatory requirements (e.g., GDPR and HIPAA), as well as policy-driven deployment gates so that only approved models go live with controlled rollouts and rollbacks [Sinha2025ModelOps]. It also enables continuous risk monitoring to detect drift, bias and fairness issues, hallucinations, and robustness regressions.
For AMAS, ModelOps must additionally support prompt and toolchain versioning, enforcement of guardrails and content filtering, fine-tuning lineage tracking, usage and cost observability, and ethical compliance. Effective ModelOps for LLM agents therefore requires:
  • Version control and CI/CD pipelines: Track versions of each agent’s model, prompts, tools, and configuration. Automated pipelines should test safety, performance, and cost impacts whenever agent logic changes, using multi-agent simulations to validate that updates do not introduce regressions or unsafe interactions.
  • Hierarchical monitoring in multi-agent frameworks: Systems such as MegaAgent [wang2025megaagent] decompose tasks into subtasks and dynamically form agent groups. They apply hierarchical monitoring, where agents log actions (e.g., via checklists), group-level administrators oversee agent teams, and system-level supervisors review final outputs. This layered oversight helps ensure that changes in one agent do not compromise system-wide behavior.
  • Risk measurement and oversight policies: The NIST generative AI profile recommends documenting risks, updating measurement approaches regularly, and establishing lifecycle oversight policies. It also highlights techniques such as gradient-based attributions, occlusion tests, and counterfactual prompting to strengthen transparency and explainability [ai2024artificial].
4.2.3 Application Security for LLM Agents
Security in AMAS must address vulnerabilities such as prompt injection, data exfiltration, and the misuse of external tools. OWASP identifies prompt injection as a core risk in LLM applications: attackers manipulate the model with malicious instructions, potentially causing data leakage, unauthorized tool use, or unsafe decisions [owasp2024llmtop10, owasp2025agentic]. Multimodal models further face cross-modal injection, where malicious content in one modality (e.g., text in an image) influences behavior in another. To mitigate these risks:
  • Prompt hygiene and hardening: Sanitize and filter user inputs, apply secure system prefixes, and validate instructions before they reach an agent. Constraining output formats and enforcing guardrails can reduce the attack surface [raza2025justhumansneedvaccines].
  • Defense-in-depth design patterns: Techniques such as adversarial training and user confirmation provide partial protection, but do not guarantee safety. Design patterns [beurer2025design] such as Action-Selector and Plan-Then-Execute restrict capabilities after untrusted input is processed: the agent selects from a predefined set of allowed actions, or produces a plan that cannot be altered by tool outputs, thereby isolating untrusted data from execution paths [qiao2024agent].
  • Least-privilege access and authentication: Enforce strong authentication for both users and agents, and apply role-based access control so agents can invoke only the minimum set of tools required for their task. Continuous monitoring should detect anomalous requests; if an agent deviates from expected behavior, automated systems should flag or suspend it [barbera2025ai].
  • Post-processing and anomaly detection: Apply output filters to detect sensitive or inaccurate content, and update or retrain models to reduce hallucinations. Maintain robust logging and retention policies to support audits and incident response while minimizing exposure of private data.
Recent frameworks also emphasize hierarchical monitoring and cross-checking among agents to detect inconsistencies [wang2025megaagent]. Tooling ecosystems such as LangChain [langchain2024function], AutoGen [autogen_agents_tutorial], and CrewAI [crewAI] operationalize verification through agent-level checks (e.g., reviewers/critics) and, in some cases, trust or reputation signals that encourage agents to validate each other’s outputs and flag potential compromise. Finally, adversarial training with attack examples [standen2025adversarial] and systematic red-teaming [ganguli_red_2022] can further strengthen robustness against malicious inputs.
4.2.4 Model Privacy and Data Protection
Without strict controls, an agent may inadvertently leak sensitive data [ko2025sevensecuritychallengessolved]. A layered privacy strategy should include:
  • Differential Privacy (DP): Adding calibrated noise to training data prevents individual data points from being re‑identified and reduces the risk that models memorize or regurgitate personal information [dwork2006calibrating]. DP should be assessed on a case‑by‑case basis, since heavy noise can harm accuracy.
  • Data anonymization and minimization: Apply robust anonymization and pseudonymization techniques and regularly test them for effectiveness [sweeney2002k]. Limit data collection to what is strictly necessary for the agent’s task and enforce strict retention policies.
  • Secure Multi‑Party Computation (MPC): MPC allows parties to jointly compute a function without revealing their private inputs. In privacy‑preserving LLM services, secure MPC can protect user prompts from the model provider; for example, the PUMA framework uses MPC to ensure that prompts remain hidden during inference, while maintaining reasonable performance (it can evaluate a 7‑billion‑parameter model within minutes) [lindell2009secure].
  • Homomorphic Encryption (HE): HE enables computation on encrypted data and offers cryptographically provable privacy. Recent work demonstrates an encryption‑friendly LLM architecture that uses LoRA fine‑tuning and Gaussian kernels to speed up fully homomorphic encryption (FHE) operations, allowing encrypted inference with performance comparable to plaintext models [gentry2009fully].
  • Trusted Execution Environments (TEEs): TEEs provide hardware‑enforced isolation that protects code and data [costan2016intel]. They partition processors into secure and normal worlds; implementations include Intel SGX, AMD SEV and Arm TrustZone. In secure LLM inference, TEEs ensure that models run only inside authenticated enclaves, with attestation verifying model integrity.
  • Continuous monitoring and access controls: Encrypt data at rest and in transit, restrict access to authorized personnel, and use monitoring to detect data poisoning or unusual behaviour.
4.2.5 Governance
Governance integrates ethical, legal, and societal considerations into the design and operation of AI systems. Below, we summarize how major governance frameworks inform the TRiSM principles.
  • EU AI Act — Human Oversight (Article 14): High-risk AI systems must be designed so that natural persons can effectively oversee them and prevent or minimize risks to health, safety, or fundamental rights [EU_AI_Act2025]. Oversight mechanisms should enable users to understand capabilities and limitations, detect anomalies, interpret outputs, decide not to use the system, or stop operation (e.g., via a “stop” mechanism). In certain high-risk identification contexts, decisions may require independent verification by multiple competent individuals.
  • OECD AI Principles: These intergovernmental principles (updated in 2024) promote human-centric, trustworthy AI [oecd_ai_catalogue]. They emphasize transparency and explainability: AI actors should provide meaningful information about data sources, factors, and logic so affected individuals can understand and, where appropriate, challenge outputs. Robustness, security, and safety require systems to function appropriately under normal and adverse conditions, including mechanisms to override or decommission systems to prevent harm. Accountability requires that AI actors ensure traceability of datasets and decisions and apply systematic risk management throughout the lifecycle.
  • NIST AI Risk Management Framework (AI RMF): The AI RMF and its Generative AI profile define four core functions: Govern, Map, Measure, and Manage [ai2023artificial]. For generative AI, NIST recommends documenting risks and potential impacts, applying interpretability techniques (e.g., gradient-based attributions, occlusion tests, and counterfactual prompting) to improve transparency, and establishing policies for oversight across the AI lifecycle.
  • Auditability and role-based access: AMAS should maintain auditable logs of agent actions, decisions, and data flows, and assign clear roles (e.g., reasoning, verification, summarization) so responsibilities remain transparent [khan2024role]. Role-based access control (RBAC) helps enforce accountability and prevents unauthorized agents from invoking sensitive functions.
  • Global harmonisation: Many jurisdictions (e.g., GDPR, CCPA) require data-subject rights, consent, and privacy by design. Frameworks such as the OECD Principles and NIST AI RMF also promote international cooperation and interoperability.
These governance frameworks (Figure 12) highlight the need for clear accountability structures, human oversight, robust privacy controls, and documented processes in AMAS. TRiSM provides a practical scaffolding to operationalize these requirements by aligning lifecycle management, security, privacy, and governance to support trustworthy agentic AI.
Figure 12: Governance alignment of TRiSM pillars with international frameworks such as the EU AI Act, NIST AI RMF, ISO/IEC 42001, 42005, and OECD Principles. Each pillar operationalizes specific compliance strategies to ensure transparency, resilience, and trust in LLM-based systems.
Alignment of TRiSM Pillars with Governance Frameworks Explainability: Aligned with the EU AI Act (Art. 13–14) [eu_ai_act] and the NIST AI RMF trustworthiness characteristic “Explainability & Interpretability” [ai2023artificial]; also supported by NISTIR 8312 (Four Principles of Explainable AI) [phillips2021four]. Promotes well-justified (rationale) and human-interpretable outputs with traceable logs [tomsett2018interpretable]. ModelOps: Anchored in the NIST AI RMF “Govern/Manage” functions [ai2023artificial] and ISO/IEC 42001:2023 (AI management system) [thiers2024emerging]; secure SDLC controls via NIST SP 800-218/218A [^9] NIST, 2024. Emphasizes lifecycle gates, model versioning, rollback policies, and CI/CD with safety gates. Risk Management & Impact Assessment: EU AI Act (Art. 9, 72–73) [eu_ai_act], ISO/IEC 23894 (AI risk management) [^10] ISO/IEC 23894, and ISO/IEC 42005:2025 (AI impact assessment) [ISO_IEC_42005_2025]. Iterative hazard analysis, impact registers, incident response, and disclosure playbooks. Data Governance & Quality: EU AI Act (Art. 10) [eu_ai_act] mapped to the ISO/IEC 5259 series (data-quality model, measures, processes, governance) [^11] ISO/IEC 5259 series. Data minimization, lineage, coverage/bias checks, and sampling documentation. Robustness & Safety: EU AI Act (Art. 15) [eu_ai_act] and ISO/IEC 24029-1 (robustness assessment) [iso24029-1]. Adversarial/perturbation testing, fault tolerance, fallback/kill-switch triggers, and rollback criteria. Application Security: Guided by OWASP LLM Top-10 [owasp2024llmtop10] and OECD AI Principles on robustness [oecd_ai_catalogue]. Access controls, prompt/input mediation, sandboxing, egress controls, and auditable trails. Model Privacy: Aligned with GDPR (Art. 25) [gdpr25], HIPAA §164 [hipaa164], and OECD AI Principles [oecd_ai_catalogue]. Encryption in transit/at rest, DP budgeting, data minimization, and access-audit pass thresholds. Documentation & Traceability: EU AI Act (Arts. 11–12; Annex IV) [eu_ai_act]. Complete technical documentation, event logs, model/data cards, decision records, and regulator-ready evidence packages. Post-Market Monitoring: EU AI Act (Arts. 72–73) [eu_ai_act]. Telemetry for drift/misuse, continuous evaluations, feedback loops, controlled recalls, and stakeholder notifications. Governance: Grounded in GDPR, CCPA [oag_ccpa_2024], and OECD AI Principles [oecd_ai_catalogue]. Institutional oversight, policy compliance, role-based accountability, regulatory traceability, and human-in-the-loop for high-risk/multi-agent deployments. Fairness & Human Rights: OECD AI Principles [oecd_ai_catalogue] and ISO/IEC TR 24027 (bias) [^12] ISO/IEC TR 24027. Lifecycle bias controls, affected-party impact reviews, and mitigation sign-offs.
Key Insights
This section operationalized the TRiSM framework for AMAS across five pillars: Explainability, ModelOps, Application Security, Model Privacy, and Governance. Multi-agent explainability requires novel techniques including Layered CoT decomposition, decision provenance graphs, and role-based interpretability frameworks to trace distributed decision processes. ModelOps extends lifecycle management through hierarchical monitoring, prompt versioning, and CI/CD safety gates with pre-deployment simulation. Security strategies emphasize defense-in-depth via design patterns (Plan-Then-Execute), least-privilege access control, and cross-agent verification. Privacy preservation leverages differential privacy, secure multi-party computation, homomorphic encryption, and trusted execution environments to protect sensitive data during inter-agent communication. Governance integration with regulatory frameworks (EU AI Act, NIST AI RMF, ISO/IEC 42001) ensures that TRiSM implementation satisfies transparency, accountability, and human oversight requirements for high-stakes deployments.
Next, we discuss the evaluations in AMAS.
5 TRiSM Evaluation of Agentic AI Systems
Accuracy alone is insufficient for Agentic AI, where failures often arise from coordination breakdowns, unsafe tool use, or misaligned autonomy rather than incorrect predictions. We organize evaluation into five dimensions: trustworthiness, explainability, user-centered performance, coordination, and composite reporting, each capturing a distinct facet of real-world performance and impact. We summarize these dimensions in Table 6.
Trustworthiness
This dimension evaluates reliability, safety, and ethical alignment [liu2023gevalnlgevaluationusing]. Core signals include success under distributional shift (robustness), safety/ethical violation rate (lower is better), and calibration quality (e.g., ECE/Brier). To avoid pathological scoring, we define a normalized trustworthiness index:
$T\;=\;\frac{w_{\text{acc}}\,A+w_{\text{rob}}\,R+w_{\text{align}}\,L}{1+\lambda\,V}\,$, (1)
where $A$ is calibrated task success (accuracy adjusted for calibration), $R$ is robustness under shifts, $L$ is alignment/safety compliance (e.g., 1violation rate1{-$\text${violation rate}), all normalized to $[0,1]$, and $V\!\geq\!0$ is a normalized violation/self-serving penalty term (higher is worse). The weights satisfy $w_{\text{acc}}{+}w_{\text{rob}}{+}w_{\text{align}}=1$ and $\lambda\!\geq\!0$ controls how strongly violations penalize trust.
Explainability
Explainability metrics assess how well humans can understand and trace decisions. Useful measures include coverage (fraction of outputs with explanations), faithfulness/fidelity (agreement with the model’s actual decision basis), stability (consistency under small perturbations), and human interpretability ratings. Suites such as OpenXAI [agarwal2024openxaitransparentevaluationmodel] report faithfulness, stability, and fairness-related properties of explanations. High explainability supports debugging and, in regulated domains (e.g., healthcare and finance), is often necessary for compliance.
User-Centered Performance
User-centered metrics capture how effectively the agent satisfies end-user needs [rodden2010measuring]. Signals include post-interaction satisfaction (e.g., CSAT), goal-fulfillment rate, number of clarification turns (lower is better), and perceived coherence/naturalness. Human-in-the-loop (HITL) studies commonly rate helpfulness, clarity, and instruction adherence. A user-centered agent aligns actions with user intent and preferences.
Table 6: Summary of key evaluation dimensions for Agentic AI systems.
Aspect Focus Strengths Limitations
Trustworthiness Reliability, safety, alignment (OOD success, safety violations, calibration, fairness). Deployment readiness; regulatory alignment (EU AI Act); statistically grounded; captures safety failures. Assumes stationarity; fairness trade-offs across groups [raza_fair_2024]; incomplete edge-case coverage.
Explainability Transparency and traceability (coverage, fidelity, stability, human interpretability). Supports debugging and compliance [ai2023artificial]; improves calibrated trust; standardized benchmarks (OpenXAI [agarwal2024openxaitransparentevaluationmodel]). Expensive fidelity checks; faithful $\neq$ interpretable; weak multi-agent attribution.
User-Centered User experience and outcomes (CSAT, task success, clarification turns, NASA-TLX). Reflects real-world utility; identifies usability bottlenecks; validated human factors measures. Subjective and high variance; costly user studies; limited scalability.
Coordination Multi-agent collaboration (task success, communication cost, plan consistency, deadlocks). Captures emergent AMAS behavior; measures coordination efficiency; decomposable contribution metrics. No standard benchmarks; sensitive to orchestration; may penalize useful redundancy.
Composite Aggregate system performance (weighted scores, tool-use efficacy). Single actionable summary; enables cross-system comparison; supports thresholding. Subjective weighting; can mask single-dimension failures; prone to Goodhart effects.
Coordination
In multi-agent or modular systems, coordination measures whether components maintain a shared plan, minimize redundant work, and resolve dependencies. Common signals include team success rate, communication efficiency (messages/tokens and rounds-to-consensus), plan/belief consistency, and deadlock/conflict rate. We additionally report: (i) a Component Synergy Score (CSS) to credit enabling contributions across agents, and (ii) Tool Utilization Efficacy (TUE) to assess correctness and efficiency of tool calls.
Composite Reporting
Composite metrics are often presented as a weighted scalar summary across dimensions:
$M_{\text{comp}}\;=\;\omega_{T}M_{T}+\omega_{E}M_{E}+\omega_{U}M_{U}+\omega_{Co}M_{Co}\,$, (2)
where $M_{T},M_{E},M_{U},M_{Co}\in[0,1]$ are normalized scores for trustworthiness, explainability, user-centered performance, and coordination, respectively, and $\omega_{\bullet}\geq 0$ with $\sum\omega_{\bullet}=1$ encode domain priorities. Since scalarization can hide critical failures, we recommend always reporting the full metric vector:
$\mathbf{M}\;=\;(M_{T},M_{E},M_{U},M_{Co})\,.$ (3)
Model selection should then be based on (i) minimum thresholds (e.g., $M_{T}\geq\tau_{T}$ and $M_{Co}\geq\tau_{Co}$) and/or (ii) Pareto-optimality (non-dominated models) rather than relying only on a single weighted score.
Component Synergy Score (CSS)
Definition: CSS quantifies the quality of inter-agent collaboration by measuring how effectively one agent’s actions enable or enhance a peer’s downstream performance.
Formal expression:
$\text{CSS}=\frac{1}{N_{\text{pairs}}}\sum_{i}\sum_{j\neq i}\Big[\text{Impact}(a_{i}\!\rightarrow\!a_{j})\cdot\text{Quality}(a_{j}\mid a_{i})\Big],$ (4)
where $N_{\text{pairs}}$ is the number of ordered agent pairs $(i,j)$ included in the evaluation. $\text{Impact}(a_{i}\!\rightarrow\!a_{j})$ measures how $a_{i}$ changes $a_{j}$’s subsequent performance (e.g., relative improvement), and $\text{Quality}(a_{j}\mid a_{i})$ is the quality of $a_{j}$’s output conditioned on $a_{i}$’s contribution.
Illustrative example: In a software AMAS with Planner, Coder, and Tester:
  • Planner’s decomposition speeds up Coder by 35% ($\text{Impact}=0.35$).
  • Tester achieves 92% coverage conditioned on Coder’s output ($\text{Quality}=0.92$).
If we score only this link, then $\text{CSS}=0.35\times 0.92\approx 0.32$ (and increases as additional enabling links are included).
Tool Utilization Efficacy (TUE)
Definition: TUE evaluates how correctly and efficiently agents invoke external tools (APIs, databases, code interpreters) within their workflows.
Components:
  • $\text{Sel}\in[0,1]$ — Selection (decision precision/recall for tool use)
  • $\text{Arg}\in[0,1]$ — Argument validity (well-formed parameters)
  • $\text{Exec}\in[0,1]$ — Execution success (no errors/timeouts)
  • $\text{Out}\in[0,1]$ — Outcome correctness (downstream task correctness)
  • $\text{Eff}\in[0,1]$ — Efficiency (normalized inverse of latency/token/call cost)
Linear aggregator:
$\text{TUE}_{\text{lin}}=\alpha_{s}\text{Sel}+\alpha_{a}\text{Arg}+\alpha_{e}\text{Exec}+\alpha_{o}\text{Out}+\alpha_{f}\text{Eff},$ (5)
where $\alpha_{\bullet}\geq 0$ and $\sum\alpha_{\bullet}=1$.
Key Insights
This section defined a multi-dimensional evaluation framework covering trustworthiness (robustness, violation rates, calibration), explainability (coverage, faithfulness, stability), user-centered performance (satisfaction and goal fulfillment), and coordination quality. CSS captures positive inter-agent enablement, while TUE decomposes tool use into selection, argument validity, execution success, outcome correctness, and efficiency. Finally, the framework emphasizes reporting a metric vector and using thresholding/Pareto analysis to reduce the risk that scalar composites mask critical single-dimension failures.
6 Security and Privacy in Agentic AI Systems
Evaluating the security and robustness of AMAS requires standardized benchmarks and well-defined metrics to assess both attack vulnerabilities and defense effectiveness. This subsection reviews current evaluation frameworks and security metrics for agentic AI, aligned with the TRiSM pillars.
6.1 Security Mechanisms
Agentic AI systems (multiple collaborating LLM agents) have a bigger attack surface than traditional agents, so they need layered defenses: protecting data, execution, communication, and model robustness; using encryption, access control, adversarial defenses, and runtime monitoring tailored to multi-agent setups.
Encryption. Encryption plays an important role in safeguarding data exchanged between multiple agents, especially when sensitive or regulated content (e.g., healthcare records, financial data) is involved [feretzakis2024privacy]. Agentic workflows often include the inter-agent handoff of partially processed results, models, or prompts. Implementations such as SSL/TLS, homomorphic encryption , and secure enclaves (e.g., Intel SGX) are increasingly integrated into Agentic AI pipelines to ensure confidentiality across message-passing protocols.
Access control. Access control is highly important when orchestrators or shared memory modules manage permissions for agents with distinct capabilities and responsibilities. For instance, in systems like AutoGen and CrewAI where agents take on specialized roles (e.g., summarizer, planner, coder); enforcing principle-of-least-privilege access prevents privilege escalation and unauthorized tool invocation [wu2023autogen]. Agent-based access control policies, often aligned with Role-Based Access Control (RBAC) [khan2024role] and Attribute-Based Access Control (ABAC) [khan2024role] paradigms, can dynamically restrict which agents may access sensitive APIs, files, or memory buffers, based on contextual trust levels.
Adversarial learning. Adversarial threats are increasing for LLM-based agents, including prompt injection, poisoned tool outputs, and disruption via malformed intermediate results. Recent work shows that compromising one agent can propagate misleading outputs and destabilize an entire multi-agent framework [zhang2024breaking]. Mitigations include adversarial training (e.g., input perturbation, reward shaping, contrastive learning) and practical safeguards such as enforcing safety constraints and validating tool outputs before execution.
Runtime Monitoring. Runtime monitoring systems support the detection of anomalous agent behaviors, especially in high-stakes domains like automated healthcare or cybersecurity. Log-based auditing, anomaly detection with LSTM or autoencoder-based detectors, and trust scoring among agents are becoming essential components of real-time surveillance layers [zhen2025novel]. For example, Microsoft’s Copilot governance layers monitor anomalous agent behavior across sessions to ensure compliant execution and flag potentially harmful interactions.
6.2 Privacy-Preserving Techniques
The decentralized and interactive nature of AMAS introduces new challenges for preserving privacy, especially as agents continuously communicate, access external data sources, and store episodic or shared memory. To ensure data confidentiality and protect personally identifiable information (PII), Agentic AI systems must adopt robust privacy-preserving techniques such as differential privacy, data minimization, and secure computation.
Differential privacy (DP). DP protects individuals by adding controlled noise, and in multi-agent LLM systems it can be used during training or inference to limit what agents can reveal to each other. Using approaches like DP in federated learning, DP-SGD, and an $\epsilon$ privacy budget, agents can collaboratively train, fine-tune, or coordinate while reducing exposure of raw or identifiable data [lazaros2024federated].
Data minimization. Agentic AI systems can mitigate exposure risks by limiting the scope, granularity, and duration of data collected or retained during task execution. For instance, temporary memory buffers used in systems like ChatDev [qian2023chatdev] or ReAct-based pipelines [yao2023reactsynergizingreasoningacting] are cleared once subgoals are completed, preventing persistent storage of unnecessary user data. Furthermore, anonymization and pseudony [luzon2024tutorial] techniques can help remove identifying features before data is passed between agents or stored in shared memory repositories.
Secure computation. Techniques including secure multi-party computation (SMPC) [evans2018pragmatic], homomorphic encryption [acar2018survey], and trusted execution environments (TEEs) [^13] Trusted Execution Environment (Wikipedia) enable agents to perform computations over encrypted or obfuscated data without compromising privacy. In scenarios where agents collaborate across different organizational boundaries (e.g., federated medical agents or cross-silo industrial agents), SMPC allows joint computations such as diagnostics or anomaly detection without data leakage. Homomorphic encryption, while computationally expensive, is increasingly being explored to allow arithmetic operations on encrypted vectors used in RAG workflows.
Privacy-by-design [schaar2010privacy] principles are becoming central to the engineering of next-generation Agentic AI systems. Architectures now embed user consent layers, configurable privacy settings, and memory redaction modules that allow end-users or system administrators to control what agents can remember or share. As Agentic AI expands into domains such as personalized education, healthcare, and finance, ensuring privacy-respecting behaviors will be essential for regulatory compliance (e.g., GDPR, HIPAA) and public trust.
Table 7: Governance controls for agentic AI systems: pillar $\rightarrow$ control $\rightarrow$ operationalization $\rightarrow$ evidence $\rightarrow$ standards.
Pillar Control Operationalization (How) Evidence / Artifacts Standards & Articles
Governance Org. accountability Define owners for models/agents/tools; RACI for changes; approvals for high-impact actions Governance policy; RACI records; change approvals NIST AI RMF (Govern) [tabassi2023ai]; ISO/IEC 42001 [iso42001]
Governance Policy-as-code Encode allow/deny rules in orchestrator; pre-exec checks and human sign-off gates Policy repository; policy test cases; gate logs EU AI Act Arts. 11–14 [eu_ai_act]; ISO/IEC 42001 [iso42001]
Risk Mgmt Risk register & AIA Hazard/threat enumeration; likelihood/impact scoring; AI Impact Assessment (AIA) Risk register; AIA reports; mitigation plan ISO/IEC 23894; ISO/IEC 42005 [ISO_IEC_42005_2025]; EU AI Act Art. 9 [eu_ai_act]
Risk Mgmt Post-market monitoring Drift/jailbreak telemetry; incident thresholds; rollback playbooks PMM plan; incident tickets; root-cause analysis EU AI Act Arts. 72–73 [eu_ai_act]; NIST AI RMF (Manage) [tabassi2023ai]
Data Gov Data minimization & lineage Task/tenant-scoped memory; TTLs/redaction; dataset lineage and licenses Data inventory; lineage graphs; retention logs EU GDPR Art. 25 [gdpr25]; ISO/IEC 5259 series
Data Gov Quality & bias checks Coverage/balance checks; label noise audits; sampling documentation Data quality reports; bias audit results EU AI Act Art. 10 [eu_ai_act]; ISO/IEC TR 24027
Explainability Faithfulness/stability Log rationales; test explanation faithfulness/stability; user-facing summaries XAI eval report; model/data/agent cards NISTIR 8312 [phillips2021four]; NIST AI RMF (Map/Measure) [tabassi2023ai]
Documentation Technical documentation Annex-IV tech docs: purpose, data, performance, oversight, foreseeable misuse Annex-IV dossier; limitations & misuse notes EU AI Act Arts. 11–12; Annex IV [eu_ai_act]
Traceability Provenance & logging Log prompts, plans, tool calls, I/O, timestamps, agent role; sign releases Immutable logs (WORM/tamper-evident); signed artifacts EU AI Act Arts. 12–13 [eu_ai_act]; ISO/IEC 42001 [iso42001]
Human Oversight Human-in-the-loop gates Require human approval for sensitive plans (delete, external write, PII) Approval records; override/rollback logs EU AI Act Art. 14 [eu_ai_act]; NIST AI RMF (Manage) [tabassi2023ai]
Robustness Stress/red-team testing Adversarial prompts; perturbation/fault tests; safety regression suites Red-team report; robustness curves; residual risk EU AI Act Art. 15 [eu_ai_act]; ISO/IEC 24029-1 [iso24029-1]
Security I/O mediation Prompt sanitization; output filtering; sensitive-data (PII/PHI/secrets) detection Filter policies; violation/egress logs OWASP LLM Top-10 [owasp2024llmtop10]; HIPAA §164 [hipaa164]
Security Tool access control RBAC/ABAC per agent/role; per-tool API tokens; argument validation; allowlists Access matrices; token scopes; arg-validator tests NIST SP 800-218/218A; NIST AI RMF (Manage)
Security Isolation/sandboxing Containers/VMs; syscall/network/file egress policies; hardware TEEs for sensitive code Sandbox configs; egress policy; attestation logs NIST SP 800-218A; TEEs best-practice notes
Privacy DP budgeting & consent Differential privacy budgets; consent bases; data-subject rights flows DP budget ledger; DSR tickets; consent records GDPR Art. 25 [gdpr25]; OECD principles [oecd_ai_catalogue]
Supply Chain Provenance & SBOM Track sources and hashes for models/tools/datasets; dependency scanning SBOM; third-party attestations; license checks NIST SSDF (SP 800-218/218A); ISO/IEC 42001
Coordination Role separation & least privilege Distinct agent roles (planner, coder, reviewer); minimal cross-role permissions Role definitions; privilege matrices; audit samples NIST AI RMF (Govern) [tabassi2023ai]; ISO/IEC 42001
Coordination Plan consistency checks Shared plan/belief verification; conflict/deadlock detection; CSS metric Plan snapshots; CSS scores; conflict logs Multi-agent eval practice; robustness [iso24029-1]
Monitoring Anomaly & exfil detection Detect unusual tool chains, volume spikes, strange domains; rate-limit/kill switch Anomaly dashboards; blocked-egress records OWASP LLM Top-10 [owasp2024llmtop10]; NIST AI RMF (Manage)
Transparency User notices & recourse User-facing disclosures, capabilities/limits, and appeal/feedback channels Notices; feedback logs; SLA for appeals OECD principles [oecd_ai_catalogue]; EU AI Act transparency
Evaluation Fairness & calibration Group metrics; ECE/Brier; CI reporting; risk–benefit trade-off docs Fairness report; calibration plots; CI tables NIST AI RMF (Measure) [tabassi2023ai]; ISO/IEC TR 24027
Records Release/rollback discipline Versioned releases; rollback points; emergency disable/recall Release notes; rollback attestations EU AI Act PMM [eu_ai_act]; ISO/IEC 42001
GPAI GPAI obligations Provide model cards, training-data summary, eval reports, and usage restrictions; disclose known limitations GPAI documentation pack; model card; eval report EU AI Act (GPAI duties) [eu_ai_act]; ISO/IEC 42001
Transparency Synthetic content disclosure Label AI-generated/edited media; attach provenance (e.g., C2PA) and user notices Watermarking/provenance config; disclosure logs EU AI Act Art. 52 [eu_ai_act]; C2PA141414c2pa.org
Privacy Cross-border transfers SCCs/TIAs for third-country transfers; data-residency controls; key management SCC/TIA records; DPA; data-map; key-custody logs GDPR Ch. V [gdpr25]
Privacy Data-subject rights ops Verified workflows for access, erasure, rectification, restriction, portability, objection DSR tickets/SLAs; fulfillment proofs; audit samples GDPR Arts. 12–23 [gdpr25]
Operations Change mgmt & release gates Risk-based pre-release checklist; dual-control approvals; staged rollouts; rollback criteria Change tickets; gate results; approvals; rollback points ISO/IEC 42001; NIST AI RMF (Manage) [tabassi2023ai]
Supply Chain Third-party/vendor risk Vendor questionnaires; DPAs; penetration/attestation reports; license compliance Vendor risk register; DPA file; attestation/SBOM NIST SSDF (SP 800-218/218A); ISO/IEC 42001
Human Factors HCD & accessibility Human-centred design reviews; usability risk analysis; accessibility conformance checks HCD review notes; usability test reports; a11y checklists ISO 9241-210151515ISO 9241-210; OECD principles [oecd_ai_catalogue]
Sustainability Energy/CO2 telemetry Track/normalize energy use and CO2; report per release; set budgets Energy/CO2 logs; budget vs. actual; disclosure note Org policy / reporting guidance
6.3 Evaluation Benchmarks and Metrics for Agentic AI Security
Evaluating the security and robustness of AMAS requires standardized benchmarks and well-defined metrics to assess both attack vulnerabilities and defense effectiveness. This subsection reviews current evaluation frameworks and metrics related to agentic AI security, aligned with the TRiSM pillars.
Attack Evaluation Benchmarks
Several benchmark datasets have emerged to evaluate adversarial attacks against LLM-based agents. For prompt injection and jailbreak scenarios, HarmBench [mazeika2024harmbench] and JailbreakBench [chao2024jailbreakbench] provide standardized datasets for evaluating robustness against adversarial prompts, including role-playing, obfuscation, and multi-turn manipulation scenarios particularly relevant to agentic systems. Tool misuse is addressed by ToolBench [xu2023tool] and API-Bank [li2023apibank], which evaluate agents’ ability to use external tools correctly while resisting adversarial instructions that trigger unauthorized access or malicious function calls.
Multi-agent adversarial scenarios are covered by AgentBench [liu2023agentbench] and GAIA [mialon2023gaia], which test resilience against coordinated attacks, impersonation, and information manipulation in multi-agent environments. Recent datasets also evaluate vulnerabilities specific to agentic architectures, including memory poisoning attacks where adversaries inject malicious information into agent memory stores, and context manipulation where attack payloads are embedded in retrieved documents or tool outputs [raza2025justhumansneedvaccines].
Defense Evaluation Metrics
Standard metrics for assessing defense mechanisms have been established across the research community. Attack Success Rate (ASR) measures the percentage of attack attempts that successfully compromise agent behavior, bypass safety constraints, or extract sensitive information, with lower ASR indicating more robust defenses [saha2025breaking]. Robustness Score quantifies the agent’s ability to maintain intended behavior under adversarial conditions, typically measured as the degradation in task performance when subjected to attacks compared to benign scenarios [raza2025humanibench].
Defense Effectiveness Rate captures the proportion of attacks successfully detected and mitigated by defense mechanisms, including input validation, output filtering, and behavioral monitoring systems. For production deployments, False Positive and False Negative Rates are critical, measuring the rate at which legitimate user requests are incorrectly flagged as attacks and actual attacks that evade detection, respectively. Our proposed Composite Security Score (CSS), introduced in Section 5 provides a holistic assessment by combining multiple security dimensions including input validation coverage, tool access control effectiveness, memory integrity, and inter-agent authentication strength.
Table 8 provides a systematic comparison of these benchmarks, including their coverage of attack scenarios, applicable defense metrics, and alignment with TRiSM pillars.
Table 8: Evaluation Benchmarks for Agentic AI Security
Benchmark Attack Coverage Metrics TRiSM Pillars Limitations
HarmBench [mazeika2024harmbench] Prompt injection, jailbreaks ASR, robustness Security, Explainability Limited multi-agent
ToolBench [xu2023tool] Tool misuse, unauthorized API Effectiveness, FP/FN Security, Privacy Tool-focused only
AgentBench [liu2023agentbench] Multi-step attacks, coordination Task success, compliance All pillars Limited adversarial
GAIA [mialon2023gaia] Ambiguous instructions, reasoning Completion, safety Explainability, Governance Early stage
WebArena [zhou2023webarena] Web attacks, malicious content Success, safety Security, Privacy Domain-specific
HELM [liang2022helm] Robustness, fairness, bias Comprehensive ModelOps, Governance Not agent-specific
MLCommons [mlcommons2024] Harmful content, privacy leakage Safety, privacy scores Security, Privacy, Governance Evolving
Key Insights
This section presented a defense-in-depth view of AMAS security spanning data confidentiality (encryption and secure enclaves), access control (RBAC/ABAC with least privilege and dynamic trust), adversarial resilience (prompt/tool-output validation and training-based hardening), and runtime monitoring (auditing, anomaly detection, and trust scoring). Privacy protection was framed as a system property: differential privacy to limit re-identification risk, data minimization through scoped/ephemeral memory and pseudonymization, and secure computation (SMPC, homomorphic encryption, and TEEs) to enable collaboration without raw-data exposure. We also summarized emerging evaluation practice, such as benchmarks for prompt injection, tool misuse, and multi-agent manipulation, together with metrics such as ASR, robustness degradation, defense effectiveness, and FP/FN rates; to quantify both vulnerabilities and mitigation quality in a TRiSM-aligned manner. Overall, security and privacy shift from optional add-ons to foundational requirements for regulated deployments.
7 Compliance with AI Governance and Policy Frameworks for Agentic AI Systems
Recent industry surveys indicate that enterprise leaders deploying Agentic AI at scale rank security, trust, compliance, and oversight among their top concerns. A survey [^16] TechRadar (2024) of IT professionals found that while 98% plan to expand AI agent usage, 96% view them as growing security threats, with only 54% having full visibility into their data access. In parallel, 67% of business leaders are increasing budgets for security oversight [^17] Cybersecurity Dive (2024), and 53% specifically cite data privacy and compliance as their primary challenge in scaling Agentic AI [^18] Cloudera (2024). Table LABEL:tab:agentic_governance operationalizes the governance controls into concrete implementations and evidence and next we detail the compliance to AI goverace and policy in Agentic AI systems.
7.1 Regulatory Standards
Regulatory frameworks provide baseline requirements that all AI systems, including agent-based architectures, must meet. Key frameworks include:
  • NIST AI Risk Management Framework (AI RMF) [tabassi2023ai]. A voluntary U.S. framework organized around the Govern, Map, Measure, Manage functions. It emphasizes organizational governance, risk identification and measurement, and operational risk treatments. Mapping to the RMF also benefits from NISTIR 8312 (Explainable AI principles) for transparency obligations.
  • EU AI Act [eu_ai_act]. A comprehensive EU regulation with phased application: bans on certain practices apply after six months (Feb 2025), obligations for general-purpose AI begin Aug 2025, and most high-risk system obligations (e.g., risk management, documentation, human oversight, post-market monitoring) start Aug 2026. High-risk systems require ongoing risk assessment, technical documentation (Annex IV), logging/traceability (Arts. 11–12), data governance (Art. 10), human oversight (Art. 14), and robustness/accuracy (Art. 15).
  • ISO/IEC management and assessment standards. ISO/IEC 42001:2023 (AI management systems) sets organization-level governance requirements; ISO/IEC 42005:2025 (AI impact assessment) provides system-level impact-assessment guidance; ISO/IEC 23894 addresses AI risk management; ISO/IEC 24029-1 supports robustness evaluation; ISO/IEC TR 24027 addresses bias. These align well with AI RMF and the EU AI Act’s lifecycle duties.
  • Domain-specific laws. Sectoral/privacy regimes still apply: GDPR (e.g., data protection by design, Art. 25), HIPAA (45 CFR §164), and regional data-residency rules. Agentic systems should enforce data minimization, purpose limitation, and consent bases; consider DPIA for high-risk processing and AI impact assessments per ISO/IEC 42005 in parallel.
7.2 Auditability
AMAS can produce emergent and opaque behaviors; auditors must reconstruct the chain of decisions.
  • Comprehensive logging. Record prompts/contexts, plans, actions, tool calls, inputs/outputs, timestamps, and agent role/identity; capture rationales where feasible. Preserve hashes and version IDs for models, tools, and data snapshots.
[…]
  • Role-granular trails. Tag entries by agent role (e.g., Researcher, Coder, Reviewer) to localize anomalies and bias. Require reviewer countersignatures for high-impact actions.
  • Immutability. Use append-only, tamper-evident storage (e.g., WORM buckets, cryptographic sealing) for audit logs; rotate keys and attest integrity regularly.
7.3 Policy Enforcement
While regulations and audits provide oversight and after-the-fact analysis, policy enforcement mechanisms work in real-time to keep an Agentic AI system’s behavior within allowed bounds. Policy in this context refers to the codified rules and constraints that the AI agents must obey. In practice, effective policy enforcement in agentic AI includes:
  • Orchestrator-level controls. A meta-controller enforces allow/deny policies for plans and tool calls; escalate to human approval for sensitive operations (e.g., data deletion, external writes).
  • Memory & retention. Implement TTLs, redaction, and scoped memory (task- and tenant-scoped) to satisfy minimization and purpose limits; support subject-access and deletion.
  • Tool & data access. Enforce RBAC/ABAC for agents/roles; use per-tool API tokens, least privilege, and environment scoping (dev/stage/prod). Validate/normalize tool arguments.
  • Isolation. Run untrusted code/tooling in sandboxes (e.g., containers/VMs) with syscall/network/file egress policies; prefer hardware isolation (TEEs) for sensitive workloads.
  • I/O mediation. Apply content filters and egress controls on model outputs; mediate RAG inputs to prevent injection/exfiltration; detect sensitive data (PII/PHI, secrets) before release.
  • Real-time monitoring. Track anomaly signals (policy violations, unusual tool chains, data exfiltration patterns) and enforce runtime kill-switches, rollbacks, and incident triggers.
7.4 Documentation & Record-Keeping (Evidence)
To meet audit and regulator expectations, maintain:
  • Technical documentation (EU AI Act [EU_AI_Act2025]): system purpose and risk class, training/eval data governance, performance, limitations, foreseeable misuse, human oversight measures.
  • Model/Data cards with lineage, licenses, and known hazards; evaluation reports with stress tests (robustness, red-teaming), calibration/fairness metrics, and residual risk.
  • Versioned artifacts: model versions, datasets, prompts, agents’ policies, and tool catalogs with change logs; signed releases and rollback points.
7.5 Post-Market Monitoring & Incident Response
Agentic systems must be monitored after deployment and respond to harm quickly.
  • Telemetry and drift. Continuously measure real-world performance, jailbreak/abuse rates, and data distribution shifts; trigger retraining or guard updates.
  • Serious-incident handling. Define thresholds, triage workflows, and reporting timelines; keep case files linking logs, affected users, and remediation.
  • Periodic reassessment. Re-run impact/risk assessments (ISO/IEC 42005 [ISO_IEC_42005_2025], ISO/IEC 23894 [ISOIEC23894_2023] ) when capabilities, data, or context change materially.
7.6 Supply-Chain & Secure Development (GenAI)
Strengthen provenance and development hygiene for agents and tools.
  • Secure SDLC for GenAI. Follow NIST SP 800-218 (SSDF) [^19] NIST SP 800-218 (SSDF) and the GenAI profile SP 800-218A  [^20] GenAI profile SP 800-218A for model/system development; integrate threat modeling for agent tool-chains and LLM-specific risks.
  • OWASP LLM risks. Mitigate prompt injection, insecure output handling, data poisoning, model DoS, and supply-chain risks using the OWASP [owasp2024llmtop10] Top 10 for LLM applications.
  • Provenance. Track model/tool/dataset origins, licenses, and checksums; require attestations for third-party components and run dependency/secret scans in CI.
Bringing Agentic systems into compliance is not a single checklist; it is a lifecycle program that couples standards (NIST AI RMF, ISO/IEC 42001/42005), regulation (EU AI Act), secure development (NIST SP 800-218/218A), and runtime enforcement (policies, isolation, monitoring). This combination creates evidenceable trust and reduces the blast radius of inevitable failures.
Key Insights
This section operationalized compliance through systematic mappings between TRiSM pillars and regulatory instruments (EU AI Act, NIST AI RMF, ISO/IEC 42001, GDPR, HIPAA). Comprehensive auditability via immutable logging and decision provenance graphs enables regulatory audit and incident investigation. Policy-as-code mechanisms encode runtime constraints, preventing violations through pre-execution validation and human sign-off requirements for high-impact operations. Documentation rigor (technical dossiers, model cards, evaluation reports) provides regulator-ready evidence packages, while post-market monitoring implements continuous telemetry, incident handling workflows, and periodic reassessment protocols. Supply chain security addresses provenance verification through Software Bill of Materials documentation and dependency scanning. This integrated lifecycle program transforms compliance from checkbox exercise to strategic capability, building evidenceable trust through systematic controls and adaptive risk management.
8 Discussion
In this section, we summarize key findings and outline implications and future directions for deploying TRiSM in AMAS.
8.1 Technical Implications of TRiSM for Agentic AI Design
Rather than treating LLM agents as black-box decision-makers, TRiSM encourages designing agents with continuous oversight guardrails [guardrails_ai_guardrails_2024]. Recent discussions also highlight the value of specialized “guardian agents” within AMAS [securiti2025aitrism], which can filter sensitive data, establish baselines of normal behavior, and enforce runtime policies (e.g., blocking disallowed actions such as exposing personally identifiable information). Prior work also warns about “excessive agency” [owasp2025agentic], where overly autonomous agents with broad tool access may trigger unintended harmful actions (e.g., hallucination-driven execution or misinterpreted goals). TRiSM mitigates these risks by limiting autonomy through role separation, scoped permissions, and safety constraints. Emerging AMAS-specific threats such as prompt injection, memory poisoning, and cascading hallucinations further motivate built-in risk controls.
8.2 Ethical and Societal Ramifications
Beyond technical risks, deploying networks of autonomous agents raises ethical and societal concerns around accountability, fairness, and human oversight [raza2025responsible]. TRiSM governance emphasizes that organizations must retain responsibility for AI actions and avoid obscuring accountability behind opaque decision-making. In practice, each autonomous agentic decisions should be transparent enough to be understood and challenged by human reviewers when necessary [vayani2025all]. TRiSM does not remove humans from the loop; instead, it structures human-agent collaboration through defined oversight roles and escalation pathways [raza2025humanibench]. In AMAS deployments, humans may monitor agent swarms in real time, pause or shut down anomalous behavior, and adjust policies on the fly. The risk of user complacency, over-trusting autonomous agents; has also been noted as a hazard [parasuraman2010complacency], and TRiSM counteracts this by formalizing oversight and supervision requirements.
8.3 Critical Analysis of Evaluation Approaches
The evaluation framework in Sections 5 and 6.3 supports targeted diagnosis by separating trustworthiness, explainability, user-centered performance, and coordination metrics, rather than relying on monolithic accuracy [raza2024exploring]. The proposed CSS and TUE further address gaps in capturing multi-agent phenomena that single-agent metrics cannot represent [ribeiro2016should, lundberg2017unified]. However, several limitations remain. Human-centered evaluation is resource-intensive and does not scale to continuous deployment, and AMAS evaluation often lacks ground truth for emergent coordination behaviors. Improvements in one dimension may degrade another (e.g., verbose explanations increasing latency), and composite weights vary by domain, requiring clearer guidance for practitioners. Current protocols also struggle to capture behavioral drift over time, and subjective measures often omit inter-rater reliability. Finally, evaluation metrics may be gamed (Goodhart’s Law [^21] Goodhart’s law.), especially when systems optimize for a single composite score rather than underlying capability.
8.4 Implementation Challenges and Real-World Constraints
While TRiSM provides a strong foundation, deploying it in production introduces practical constraints. Continuous monitoring, validation, and anomaly detection add computational overhead and latency, which can harm user experience in time-sensitive settings. Large-scale enterprise deployments may involve hundreds or thousands of agents, increasing compute, storage, networking, and staffing costs. Integration is also challenging because many organizations rely on legacy infrastructure or third-party frameworks not designed for agentic security; adding memory protection, access control, and behavioral monitoring often requires cross-component changes and ongoing maintenance. Operationally, monitoring can produce large volumes of alerts, including false positives, which may overwhelm security teams. A phased, risk-driven adoption strategy is therefore often more feasible: start with input validation, restricted tool access, and logging, and then add stronger controls as system maturity and risk exposure increase.
8.5 Alignment with Emerging AI Regulations and Standards
TRiSM principles align closely with emerging regulatory frameworks and standards. For example, the EU AI Act [eu_ai_act] emphasizes risk management, transparency, data governance, and human oversight for high-risk systems, for example capabilities directly supported by TRiSM-aligned governance and controls [yang2025agentic, farooq2025evaluating]. This convergence suggests that TRiSM-based practices can help organizations translate technical safeguards into compliance-ready processes and evidence.
Despite progress, some gaps remain. For example, there is still no consensus on standardized benchmarks and metrics for evaluating agentic systems under TRiSM principles, which limits cross-study comparison and objective progress tracking. Many controls also lack real-world validation beyond lab settings or narrowly scoped tasks, and adversarial learning remains an ongoing challenge as attackers adapt to defenses. Future work should prioritize standardized stress-testing protocols (e.g., red-teaming), longitudinal evaluation to detect drift, and deployment case studies that report both effectiveness and operational costs.
Extension to Multimodal and Embodied Agents. While this survey focuses on LLM-based AMAS, the Agentic AI landscape increasingly encompasses multimodal agents (e.g., vision-language models such as LLaVA, GPT-4V) and embodied agents [xia2018gibson] operating in physical environments (e.g., robotic systems, autonomous vehicles, industrial automation). Future work should systematically extend TRiSM frameworks to address the unique challenges these systems present. Multimodal agents introduce additional security considerations including cross-modal adversarial attacks (e.g., adversarial images triggering unsafe text outputs), visual hallucinations, and sensor spoofing vulnerabilities. Embodied agents face physical safety risks, real-time decision-making constraints under uncertainty, sim-to-real transfer gaps, and complex human-robot trust calibration requirements. Extending the proposed CSS and TUE metrics to capture sensor fusion quality, physical action safety scores, and human-robot collaboration efficacy represents a promising direction for comprehensive agentic AI evaluation.
8.6 Future Roadmap for Agentic AI TRiSM
Agentic AI Roadmap Scalable Agent Architectures Secure Multi-Agent Collaboration Explainable Multi-Agent Decisions Human-Centered Trust Design Autonomous Lifecycle Management Ethics & Governance Controllability Benchmarking & Evaluation Cognitive Capability Expansion
Figure 13: Agentic AI Roadmap: Mindmap Representation.
Drawing from our findings and best practices from multiple disciplines, we propose several actionable directions for future research and implementation, as shown in Figure 13. These recommendations span both technical system design improvements and governance-level policy initiatives:
The community should create open benchmarks and challenge environments to test AMAS governance. For instance, a suite of scenario-based tasks (with built-in threats and ethical dilemmas) could be used to evaluate how well a TRiSM-governed ASAM that performs relative to one without such controls. This will enable direct comparisons and drive progress on measurable metrics of trust (e.g. frequency of prevented failures or fairness outcomes).
Future system design must anticipate a continually evolving threat landscape. Techniques from cybersecurity (e.g. adversarial training, AI model “penetration testing” [bacudio2011overview], and formal verification) should be integrated into the LLM agent development pipeline. Cross-disciplinary collaboration with security experts can yield LLM-specific hardening methods, such as dynamic prompt anomaly detectors or robust tool APIs that constrain agent actions. Additionally, creating red-team/blue-team exercises for AMAS , akin to cyber wargames [schneider2020wargaming] , can help discover vulnerabilities in a controlled way before real adversaries do.
We encourage designing better interfaces and protocols for human oversight of Agentic AI systems. Borrowing from human-computer interaction [AmershiWVFNCSIB19] and cognitive engineering [hexmoor2025behaviour], researchers could devise dashboards that visualize an agent society’s state, flag important decisions, and allow intuitive human intervention (pausing agents, rolling back actions, etc.).
Policymakers and industry should collaborate to create regulatory sandboxes for multi-agent AI trials. There is much to learn from other high-stakes domains. For example, the safety engineering field (e.g. aerospace, automotive) has mature practices for redundant controls and failure mode analysis; these could inspire analogous practices in AI agent design. Likewise, ethics boards in biomedical research provide a template for AI ethics committees that review agent behaviors and approve high-risk deployments. We advocate establishing multidisciplinary governance boards that include ethicists, legal experts, domain specialists, and community representatives to oversee significant deployments of autonomous AI.
9 Conclusion
Agentic AI systems built as AMAS are rapidly changing how autonomous systems collaborate, plan, and make decisions in high-stakes domains such as healthcare, finance, and public services. At the same time, these systems expand the risk surface and raise new challenges in trust, risk, and security management (TRiSM). This review addressed those challenges and introduced a TRiSM-oriented framework for LLM-based AMAS organized around five pillars: Explainability, ModelOps, Security, Privacy, and lifecycle Governance. We contributed a risk taxonomy that captures AMAS-specific threats, including prompt injection, memory poisoning, and collusive or cascading failures. We also proposed practical evaluation metrics to measure inter-agent enablement and the correctness and efficiency of tool use. In addition, we summarized key technical mechanisms to support TRiSM in practice, including explainability methods and security/privacy protections. Looking ahead, progress will depend on stronger adversarial robustness, clearer governance and audit protocols, and standardized benchmarks that evaluate both trustworthiness and coordination under realistic deployment conditions. Aligning Agentic AI development with TRiSM principles can help ensure these systems remain not only capable and scalable, but also safe, transparent, and accountable for real-world use.
Acknowledgments
This research was funded by the European Union’s Horizon Europe research and innovation programme under the AIXPERT project (Grant Agreement No. 101214389), which aims to develop an agentic, multi-layered, GenAI-powered framework for creating explainable, accountable, and transparent AI systems. Additionally, this work was supported in part by the National Science Foundation (NSF); in part by United States Department of Agriculture (USDA); in part by the National Institute of Food and Agriculture (NIFA), through the “Artificial Intelligence (AI) Institute for Agriculture” Program, Accession Number 1029004 for the Project Titled “Robotic Blossom Thinning with Soft Manipulators” under Award AWD003473, Award AWD004595, and Award USDA-NIFA; and in part by United States Department of Agriculture National Science Foundation (USDANSF), Accession Number 1031712, under the Project “ExPanding University of Central Florida (UCF) AI Research To Novel Agricultural EngineeRing Applications (PARTNER)” under Grant 2024-67022-41788.
Declaration of Competing Interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Author Contributions
S.R. and R.S. contributed equally to this work. S.R.: Conceptualization, Methodology, Literature review, Bibliographic analysis, Writing, Re-writing, Editing, Revision R.S: Conceptualization, Methodology, Literature review, Bibliographic analysis, Writing, Re-writing, Editing M.K.: Supervision, Detailed review, Structuring, Writing. C.E.: Detailed review, Writing, Re-writing, Supervision, Revision. All authors have read and approved the final manuscript.
Appendix
Table A.1: List of abbreviations used in this paper.
Abbreviation Description
AI Artificial Intelligence
AMAS Agentic Multi-Agent Systems
CoT Chain-of-Thought
CSS Component Synergy Score
DP Differential Privacy
FHE Fully Homomorphic Encryption
HITL Human-in-the-Loop
HE Homomorphic Encryption
IAM Identity and Access Management
LLM Large Language Model
MLOps Machine Learning Operations
ModelOps Model Lifecycle Operations, Monitoring, and Governance
MPC Secure Multi-Party Computation
PII Personally Identifiable Information
RAG Retrieval-Augmented Generation
RBAC Role-Based Access Control
RMF Risk Management Framework
TEEs Trusted Execution Environments
TRiSM Trust, Risk, and Security Management
TUE Tool Utilization Efficacy
XAI Explainable Artificial Intelligence
Figure A.1: Traditional AI Agent vs. Agentic AI.
Table A.2: Key terminologies for LLM‐based agentic AI systems.
Term Definition
Agentic AI System A multi-agent artificial intelligence system powered by large language models (LLMs), in which multiple autonomous agents coordinate, communicate, and reason collectively to plan and execute tasks over extended horizons. Such systems typically incorporate persistent memory, tool use, and dynamic role assignment to support adaptive, goal-driven behavior.
Autonomy Model The internal decision-making mechanism that governs how an agent selects goals, plans actions, and executes tasks without continuous human intervention, often leveraging planning, reasoning traces, feedback signals, and environmental context.
Chain of Thought (CoT) A prompting and reasoning paradigm in which an LLM explicitly generates intermediate reasoning steps before producing a final output, enabling multi-step problem solving, improved transparency, and more controllable agent behavior.
Counterfactual Analysis An interpretability and diagnostic technique that evaluates how changes to inputs, agent roles, or internal states would alter system outcomes, helping reveal causal dependencies, failure modes, and contribution disparities across agents.
Explainability The ability of an AI system or individual agent to provide human-understandable rationales for its decisions, actions, or outputs, often supported by techniques such as feature attribution, decision provenance, or structured reasoning traces.
Foundation Model (LLM) A large pretrained language model (e.g., GPT-4, LLaMA) that serves as the core reasoning and generation component of an agent, providing linguistic competence, contextual understanding, tool invocation, and general-purpose reasoning capabilities.
Shared Memory (Persistent Memory) A centralized or distributed memory substrate, commonly implemented using vector databases or structured logs, that enables agents to store, retrieve, and share contextual information across time, supporting long-term coherence and coordination.
ModelOps A set of operational practices for managing models and agent prompts throughout their lifecycle, including versioning, deployment, monitoring, auditing, and retirement, with an emphasis on reliability, reproducibility, and governance.
Application Security Protective mechanisms and design practices that safeguard agentic systems against adversarial behaviors such as prompt injection, identity spoofing, privilege escalation, and cross-agent manipulation.
Model Privacy Privacy-preserving techniques that ensure sensitive data remains protected during model training, inference, and inter-agent communication, including differential privacy, secure computation, and trusted execution environments.
Prompt Injection A class of security attacks in which maliciously crafted inputs introduce hidden or indirect instructions that override or corrupt an agent’s intended behavior, potentially propagating harmful actions across agent interactions.
Retrieval-Augmented Generation (RAG) An architectural approach in which agents dynamically retrieve relevant external knowledge (e.g., documents or facts) from a datastore and condition LLM generation on this retrieved context to improve factuality and grounding.
Role-Specialized Coordination A coordination strategy where agents are assigned distinct functional roles (e.g., planner, verifier, executor) and interact via structured communication protocols to collectively solve complex tasks more efficiently and robustly.
Decision Provenance Graph A structured, graph-based representation that records data flows, reasoning steps, and inter-agent interactions, enabling post-hoc auditing, accountability analysis, and system-level interpretability.
Tool-Use Interface The mechanism that allows agents to issue structured commands to external tools, APIs, or execution environments and integrate returned results into subsequent reasoning and decision-making steps.
Trust Score A composite quantitative indicator of an agent’s reliability and alignment, typically combining task performance, safety compliance, behavioral consistency, and calibration of expressed confidence over time.
Composite Metric An aggregated evaluation measure that combines multiple dimensions—such as trustworthiness, explainability, coordination quality, and user-centered performance—to benchmark and compare agentic AI systems holistically.